### v1

# RingMo-Aerial: An Aerial Remote Sensing Foundation Model With A Affine Transformation Contrastive Learning

Abstract
Aerial Remote Sensing (ARS) vision tasks pose significant challenges due to the unique characteristics of their viewing angles. Existing research has primarily focused on algorithms for specific tasks, which have limited applicability in a broad range of ARS vision applications. This paper proposes the RingMo-Aerial model, aiming to fill the gap in foundation model research in the field of ARS vision. By introducing the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism and an affine transformation-based contrastive learning pre-training method, the model’s detection capability for small targets is enhanced and optimized for the tilted viewing angles characteristic of ARS. Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and effectiveness in various ARS vision tasks. Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks. This indicates the practicality and effectiveness of RingMo-Aerial in enhancing the performance of ARS vision tasks.
Index Terms:
Aerial Remote Sensing (ARS), foundation model, contrastive learning(CL), pretraining, self-supervised.
I Introduction
Figure 1: Number of Aerospace Remote Sensing and Aerial Remote Sensing papers in Web of Science in the last decade. The green bars represent aerospace remote sensing, while the orange bars represent aerial remote sensing. Although research on aerial remote sensing has grown, the number of papers remains significantly lower than that of aerospace remote sensing.
Figure 2: Illustration of multi-view, multi-resolution and occlusion of aerial remote sensing images:: (a) Multi-view imagery shows images captured from different angles, including tilt, top, and flat views; (b) Multi-resolution imagery demonstrates how varying low, medium, and high resolutions affect the level of detail; (c) Target occlusion compares cases where the target is partially occluded versus not occluded. These characteristics highlight the complexity of aerial remote sensing imagery in different scenarios.
Figure 3: Schematic Comparison of Aerospace Remote Sensing and Aerial Remote Sensing Image Features. Satellite images contain more low-frequency information with continuous features, while aerial images have more detailed variations and relatively less low-frequency information.
Aerial Remote Sensing (ARS) is a significant method of information acquisition that, unlike satellite remote sensing, is typically not constrained by revisit cycles, satellite sensor incidence angles, or other observation timing and viewing angle limitations. Aerial remote sensing sensors are commonly mounted on Unmanned Aerial Vehicles (UAVs), airplanes, or balloons. Compared to remote sensing satellites, these aerial platforms can swiftly reach the observation area and conduct comprehensive observations from multiple perspectives and at various resolutions, thereby capturing real-time and detailed information about the target area. Consequently, aerial remote sensing finds extensive applications in emergency response, environmental monitoring, and national defense security. Therefore, research on enhancing aerial remote-sensing images’ intelligent information extraction capabilities holds significant academic and practical value.
However, despite the advantages and value of aerial remote sensing imagery, our research reveals that the current focus in intelligent remote sensing interpretation remains predominantly on space-borne remote sensing. As shown in Figure 1, although research on aerial remote sensing has seen rapid growth in recent years, studies related to its intelligent interpretation remain relatively scarce. The number of published papers in this area accounts for less than 10% of space-borne remote sensing. Compared to space-borne remote sensing, aerial remote sensing possesses distinct physical characteristics, such as multi-view, multi-resolution, and occlusion effects, owing to its unique observational perspectives and sensor properties, as illustrated in Figure 2. In Figure 2(a), various perspectives are depicted, including oblique views (i and ii), a vertical view (iii), and a horizontal view (iv). Unlike the predominantly vertical perspective of satellite remote sensing, aerial remote sensing imagery captures surface information from oblique angles, enhancing the three-dimensional perception of the targets. However, this also results in distant objects appearing smaller and more densely arranged, increasing the complexity of image interpretation. Figure 2(b) compares images of different resolutions: low-resolution (i, iv), medium-resolution (ii, v), and high-resolution (iii, vi). These multi-resolution characteristics result from the varying data resolutions of aerial sensors operating at different flight altitudes and speeds. The variation in resolution affects the clarity and detail representation of the same target across different images, thereby placing higher demands on the adaptability of image interpretation models. Furthermore, aerial imagery is often characterized by significant occlusion, as shown in Figure 2(c). In this Figure, some targets are partially or fully occluded by other objects (i and iii illustrate occluded scenarios, while ii and iv show non-occluded scenarios). Such occlusions further complicate the interpretation of ARS images.
Given the challenges mentioned above, it is difficult to design specialized intelligent models for aerial remote sensing that are tailored to different applications. Existing foundational models in computer vision and remote sensing vision are no longer sufficient to meet the demands of aerial remote sensing applications. The recently developed foundational models, which have gained significant attention in remote sensing interpretation, are considered an effective solution to these challenges. These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
However, existing foundational model studies in the remote sensing domain primarily focus on the interpretation of spaceborne remote sensing imagery, as evidenced by studies like RingMo[1] and Skysense[2]. This underscores the pressing need to develop a pre-trained foundational model specifically tailored for aerial remote sensing to address its unique application requirements. When adapting current foundational models to aerial remote sensing, several critical challenges persist:
(1) Issues in the Pre-training Phase: The pre-training phase typically relies on large-scale datasets for self-supervised learning, primarily adopting two paradigms: Masked Image Modeling (MIM) and Contrastive Learning (CL). MIM, as exemplified by models like MAE [3] and SimMIM [4], involves applying masks to images and training the model to reconstruct the masked parts, thereby enhancing the model’s feature extraction capabilities. On the other hand, CL, as seen in models like MOCO [5], improves feature extraction by creating positive and negative sample pairs and training the model to learn the relationships between these pairs. Both paradigms have advantages and limitations, with MIM being more widely utilized in most visual scenarios. For instance, Swin Transformer and RingMo both utilize MIM-based pre-training methods. However, for aerial remote sensing imagery, the characteristics of images at different perspectives and resolutions differ significantly at the feature level. As shown in Figure 3, vertical images tend to exhibit more structured and continuous features, often characterized by large areas of color and texture consistency, with fewer variations in detail. Such images typically contain abundant low-frequency information due to their broad display range and relatively uniform visual presentation. In contrast, oblique images capture the sides of buildings and their height information, offering a more dynamic and three-dimensional perspective. These images are rich in detail variation and contain relatively less low-frequency information. The current self-supervised learning methods rely on reconstructing local information in images and struggle to capture these variations effectively.
(2) Issues in the Fine-tuning Phase: The traditional approach of fine-tuning the entire model presents significant challenges, primarily due to the time-consuming nature of the process. Additionally, since the foundational model is pre-trained on a diverse set of downstream tasks, fine-tuning the model globally for a specific downstream task can lead to overfitting, especially in aerial remote sensing. Unlike the single perspective and resolution typically encountered in spaceborne remote sensing, aerial remote sensing models are more prone to overfitting to specific perspectives and resolutions of a particular downstream task. This overfitting can reduce the ability of the model to generalize to different perspectives and resolutions in various downstream tasks.
To address these challenges, this study introduces RingMo-Aerial, the world’s first foundational large model designed for aerial remote sensing tasks, based on the existing spaceborne remote sensing foundational model, RingMo. In response to the unique characteristics and difficulties of aerial remote sensing tasks, we propose several novel theories and methods, as shown in Figure4.
First, we introduce Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) to address the multi-scale and occlusion issues caused by the oblique perspective in aerial remote sensing images. Unlike traditional multi-head self-attention mechanisms, FE-MSA incorporates an additional image patch expansion layer, ensuring that dense small objects are captured by more tokens when patches are divided into tokens for the Transformer model. Furthermore, FE-MSA leverages the Swin Transformer’s window attention and shifted window mechanisms. These improvements significantly enhance the model’s ability to detect dense small objects and occluded targets.
Second, to solve the problem of detecting and recognizing small distant targets in oblique perspectives, we propose a Contrastive Learning Affine Framework (CLAF) that scales and enlarges distant targets through affine transformations to construct positive samples. By inputting the transformed and original images as sample pairs into the model, this method guides the model to learn the unique visual characteristics of ARS images, improving its ability to recognize small targets at a distance.
In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic aerial remote sensing foundational models, called ARS-Adapter. Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting. ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter [6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling. This approach optimizes the performance of the aerial remote sensing foundational model, making it more adaptable and effective across various aerial remote sensing vision tasks.
In conclusion, the core contributions of this paper are as follows:
  • Addressing the Characteristics and Challenges of Aerial Remote Sensing Imagery: This paper introduces RingMo-Aerial, the world’s first foundational model designed for aerial remote sensing imagery. RingMo-Aerial supports various tasks across different perspectives, resolutions, and modalities, including scene classification, semantic segmentation, object detection, object tracking, change detection, and 3D reconstruction. It can be widely applied across multiple different scenarios.
  • FEMSA: By incorporating a patch expansion layer and integrating the window attention mechanism, FEMSA effectively enhances the detection capability of small, dense targets in aerial remote sensing images.
  • CLAF: Utilizing affine transformations, CLAF guides the model to learn the features of aerial remote sensing imagery captured from oblique perspectives, thereby improving the model’s understanding of distant regions in oblique images.
  • ARS-Adapter: The ARS-Adapter innovatively introduces a fine-tuning method tailored to foundational models in aerial remote sensing, reducing the number of parameters adjusted, improving fine-tuning efficiency, and significantly lowering the risk of overfitting.
In order to evaluate the effectiveness of the model, we performed six downstream tasks using 13 aerial remote sensing datasets and 5 aerospace remote sensing datasets with different viewpoints, resolutions and modes. The test results show that RingMo-Aerial achieves SOTA performance on all aerial remote sensing datasets and reaches the leading level on aerospace remote sensing datasets, which proves its strong generalization ability for intelligent interpretation of aerial and aerospace remote sensing images. Compared with the single-task model, RingMo-Aerial has obvious advantages in handling multi-tasks in different scenarios.
II Related Works
II-A Aerial remote sensing
Aerial remote sensing offers unique characteristics and has found widespread applications across various domains. Common tasks in aerial remote sensing include object detection, object tracking, semantic segmentation, and 3D reconstruction. The most frequently utilized modalities are optical imaging [7], infrared imaging [8], and synthetic aperture radar (SAR) [9].
Compared to satellite remote sensing, aerial remote sensing provides higher image resolution and more flexible spatiotemporal coverage. However, it also presents distinct characteristics that set it apart from satellite-based methods. One notable difference is that aerial imagery is typically captured at oblique angles with varying shooting distances. This results in significant variations in the size of similar targets and increases the likelihood of target occlusion. [10, 11, 12, 13] have specifically modeled occluded targets at the structural level, significantly improving the detection and segmentation performance of occluded instances. [14, 15, 16, 17] have focused on optimizing the oblique perspective through multi-scale transformations, enhancing the resolution of small targets. Another key distinction lies in the frequency content of aerial imagery, which contains both high- and low-frequency signals, unlike satellite imagery that predominantly captures low-frequency information. Effectively leveraging high-frequency signals can greatly aid in the interpretation of aerial images. [18, 19] introduced frequency-domain enhancement modules to balance high- and low-frequency signals, leading to improved performance across various aerial remote sensing tasks.
II-B Visual foundation model
Figure 4: Comparison of model frameworks between RingMo and RingMo-Aerial.
Visual foundation models have gained attention for their ability to serve as general-purpose solutions in various computer vision tasks. Convolutional neural networks (CNNs) were among the earliest architectures used for visual foundation models, with well-known examples such as VGG [20] and ResNet [21]. These models rely on convolutional layers for feature extraction and are typically pre-trained on classification tasks using the ImageNet [22] dataset. With the advent of Transformers in natural language processing, visual Transformer models like ViT [23] and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training. In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies. Additionally, a number of novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed. However, due to technical immaturity, these newer models have yet to achieve widespread adoption.
The key issue with these models lies in the pre-training phase, where they learn broad visual representations from large datasets. This chapter will also summarize three types of self-supervised visual training methods: CL, MIM, and their combination, with each method addressing both natural and RS.
II-B1 Contrastive Learning
Contrastive learning is an self-supervised learning approach that emphasizes learning the common features among similar instances while distinguishing differences between dissimilar instances. Major contrastive learning methods include the MoCo [5, 27, 28] series. MoCov2 [27], an optimization of MoCo [5] inspired by SimCLR [29], introduces enhancements such as the MLP projection head and increased data augmentation. MoCov3 [28] integrates MoCov2 and SimSiam [30], but training with ViT [23] as the backbone often leads to instability and reduced model accuracy. Stability and performance are improved by freezing the patch projection layer of ViT during training. DINO [31], another self-supervised learning method, employs knowledge distillation techniques to train ViT for semantic segmentation and classification feature learning without labels, leveraging momentum encoders and multi-crop training for enhanced model performance.
In the field of RS, contrastive learning also has many applications. GLCNet [32] presents a self-supervised contrastive learning method for semantic segmentation of high-resolution RS images. SeCo [33] introduces a self-supervised pretraining method that leverages temporal information in RS images to generate positive sample pairs, learning feature representations invariant and diverse to seasonal variations.
II-B2 Masked Image Modeling
MIM is an emerging self-supervised pre-training method. It involves masking a portion of an input image and then reconstructing it through a pretext task. Unlike traditional supervised pre-training, which relies on labeled data, MIM uses self-supervised learning to learn meaningful representations from images. MAE [3] pioneered the MIM by masking and reconstructing images using Vision Transformer (ViT). SimMIM [4] differs by only reconstructing masked areas and incorporates the Swin Transformer [24]. Combining MAE and SimMIM, MixMAE [34] uses visible tokens from different images for reconstruction. GreenMIM [35] optimizes MAE by introducing a layered architecture to lower computational costs. UM-MAE [36] applies uniform masking to address complexities in ViTs, while CAE [37] adds a regressor for distinct feature encoding separate from reconstruction.
In the field of RS, RingMo [1] pioneered the foundation model framework based on MIM. Unlike random masking techniques used in natural scenes, RingMo developed the PIMask strategy tailored to the complex backgrounds and small targets characteristic of RS data, focusing on preserving information about small targets. By integrating ViT and Swin Transformer, RingMo enhanced the accuracy of various downstream RS tasks, laying a foundation for developing large models in RS scenes.
II-B3 The combination of Contrastive Learning &\&& Masked Image Modeling
Combining CL and MIM has shown significant advantages in current research, particularly in enhancing feature extraction performance. For instance, CMAE [38] utilizes an asymmetric encoder-decoder structure along with a momentum encoder. This enables learning visual representations from masked images and enhances feature discriminability through CL. Additionally, MimCo [39] introduces patch-level and image-level contrastive reconstruction losses in MIM, effectively improving the model’s transferability.
In the field of RS, TMAC [40] and SegMind [41] demonstrate that combining MIM and CL significantly improves the utilization of unlabeled data. It also enhances the interaction between different image regions, boosting feature recognition capabilities.
Despite the success of models like ViT and RingMo in general and satellite remote sensing tasks, their performance in aerial remote sensing remains limited. Currently, there is no dedicated foundation model for aerial remote sensing. To address this gap, we propose RingMo-Aerial, a model designed specifically for aerial remote sensing tasks, and demonstrate its effectiveness in common applications.
III Problem Description
The unique imaging characteristics of ARS photography stem from the inherent properties of aerial perspective. Specifically, this perspective manifests as a size disparity within the same image due to the inclined angle of capture. Objects that are closer to the ARS (towards the bottom of the image) appear larger, while those further away (towards the top) appear smaller, as illustrated in Figure 5.
To elaborate, as depicted in Figure 6, when a ARS captures two objects of equal actual size at inclined angles of $\theta_{1}$ and $\theta_{2}$, the resulting pixel dimensions and their ratio can be mathematically expressed. For illustration, consider two identical cars on a highway, both with a height of L. The vertical component of the cars in the camera’s view can be expressed as $L\times sin(\theta_{1})$ and $L\times sin(\theta_{2})$. Given the drone’s flight height is $H$, the focal length of the drone’s camera is $f$, and the horizontal distances between the two cars and the drone are $d_{1}$ and $d_{2}$ respectively, the slant distances between the cars and the drone are calculated as $D_{1}=\sqrt{H^{2}+d_{1}^{2}}$ and $D_{1}=\sqrt{H^{2}+d_{2}^{2}}$. According to the principle of perspective projection, the height of the objects in pixels in the image is calculated as follows:
$h_{1}=\frac{f\cdot L\cdot\sin(\theta_{1})}{D_{1}}$ (1)
$h_{2}=\frac{f\cdot L\cdot\sin(\theta_{2})}{D_{2}}$ (2)
The ratio of the heights of the pixels occupied by the cars is:
$\frac{h_{1}}{h_{2}}=\frac{D_{2}}{D_{1}}\times\frac{\sin(\theta_{1})}{\sin(\theta_{2})}=\frac{\sqrt{H^{2}+d_{2}^{2}}}{\sqrt{H^{2}+d_{1}^{2}}}\times\frac{\sin(\theta_{1})}{\sin(\theta_{2})}$ (3)
Then the proportion of pixels occupied by car 1 and car 2 can be estimated as:
$\frac{p_{1}}{p_{2}}=\frac{h_{1}^{2}}{h_{2}^{2}}$ (4)
Figure 5: Similar targets with significant size differences in the same picture. In the Vision Transformer model, objects of different sizes occupy different numbers of tokens.The image is from VisDrone dataset.
Figure 6: The difference in shooting angles causes the final pixel sizes of similar targets to be different.
Considering that in the visual transformer model, the number of pixels occupied by an object is related to the number of tokens it occupies, as 5. Therefore, it is obviously useful to adopt a method to balance the number of tokens occupied by similar objects of different sizes in an image.
In addition to the oblique angles, ARS images typically contain many small objects, which can be very dense. CNNs have fixed convolution kernel sizes and limited receptive fields, which makes it challenging to capture large-scale information and detailed information in ARS images simultaneously. This can result in missed detections and false detections of objects. For instance, when Sensors capture congested urban streets, pedestrians and vehicles may randomly occlude other objects, making it challenging for detection algorithms to locate and classify them accurately. ViT employs the same downsampling rate and directly extracts features from image tokens, which fails to account for the spatial context of objects.
Furthermore, ARS images are typically high resolution, containing many intricate details and complex background information. This places greater demands on pre-trained models’ feature extraction and processing capabilities. Traditional pre-trained models are typically trained on large-scale natural scene datasets, which cannot adequately represent the characteristics of ARS images.
Consequently, a practical pre-training framework and method, specifically designed to address the distinctive visual characteristics of ARS images, is essential to achieve satisfactory performance in a range of downstream tasks.
Figure 7: The main structure of RingMo-Aerial.
IV Methodology
IV-A Method Overview
The proposed RingMo-Aerial model serves as a foundational framework with broad applicability. Its implementation involves three main stages: model design, pretraining, and fine-tuning for downstream tasks. In Section IV-B, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase. To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown as Figure 7. In Section IV-C, we combine the commonly used masked image modeling (MIM) pretraining with an additional contrastive learning (CL) method based on affine transformations during the pretraining stage. This hybrid pretraining approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown as Figure 7. Furthermore, in Section IV-D, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase. The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks, shown as Figure 10.
IV-B Model Architecture
RingMo-Aerial’s design is based on Swin Transformer. The architecture of Swin Transformer consists of four stages, each comprising either a linear encoder or Patch Merging, along with an even number of Swin Transformer Blocks. Each block includes Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA). These attention mechanisms effectively compute attention maps within the windows, extracting spatial feature information between pixels while reducing the computational complexity of the model.
Although the classical Swin Transformer introduces spatial interactions through two types of window attention mechanisms, emphasizing high-level semantic information within feature vectors, challenges arise in ARS imagery. The varying shooting distances, angles, and scale differences between distant and nearby objects present significant difficulties. Relying solely on semantic information may not be sufficient to achieve optimal performance in downstream tasks.
To address the aforementioned challenges, we propose the overall architecture of RingMo-Aerial. As shown in Figure 7(a), the overall architecture of RingMo-Aerial is based on the Swin Transformer. Considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
Figure 8: The effect of FE-MSA. Existing window attention mechanisms cannot balance the feature vector information of objects of different sizes in the image. At the same time, FE-MSA enhances the model’s receptive field for small objects (purple box).
Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module. To better understand the FE-MSA mechanism, it is essential to first recognize the limitations of the existing window attention mechanisms. These mechanisms often struggle to balance the semantic information in feature vectors of objects of different sizes. For instance, as illustrated in Figure 8, larger objects may occupy $4\times 4$ tokens, while smaller distant objects may occupy only $3\times 3$ tokens. The FE-MSA mechanism introduces frequency domain processing to the existing window attention, indirectly expanding the receptive field for smaller objects.
As shown in Figure 7(b), the RingMo-Aerial Block structure is illustrated, including two window attention mechanisms optimized by the FE-MSA mechanism. Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block. Taking the first stage as an example, the input feature vector $z^{l}$ adopts the same slicing parameters as the standard Swin-Transformer, resulting in four-channel slices $C=\{c_{1},c_{2},c_{3},c_{4}\}$. After linear normalization, it enters the FE-MSA module. For each patch with a dimension of $C$, a $7\times 7$ convolution kernel is used for depthwise convolution to obtain the feature map $Fz^{(l-1)}=\text{Concat}\{Fc_{1},Fc_{2},Fc_{3},Fc_{4}\}$. Then, W-MSA or SW-MSA is used to compute the attention map. After the FE-MSA, the edge and texture features of small objects are further enhanced, improving the model’s ability to perceive both local and global features. The computation process for each token patch through the RingMo-Aerial block is as follows:
$z^{l}=\text{FE-MSA}(z^{l-1})$ (5)
[…]
The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, enhancing the model’s performance in complex aerial scenes.
IV-C Pre-train
Figure 9: Positive samples for contrastive learning are obtained after affine transformation, cropping and resizing of the image.
For pretrain, we construct a large-scale pre-training dataset. Our comprehensive pre-training dataset covers a wide range of aerial scenes, including high-altitude, low-altitude, and near-ground perspectives. It spans diverse geographical regions such as urban, rural, and wilderness areas. The dataset includes various targets, such as roads, vehicles, and pedestrians, and captures different seasons and time periods. Notably, it exhibits multi-source, multi-temporal, and multi-instance characteristics. Prior to training, we crop the entire dataset into 640x640 or 448x448 resolution slices, all captured using optical three-channel imaging. In total, the sliced dataset comprises approximately 5.4 million images, as Table I.
TABLE I: The Number of Post-Cropping Images Per Dataset.
Datasets Images
AU-AIR [42] 328230
DAC-SDC [43] 73740
DarkTrack2021 [44] 223606
DroneCrowd [45] 42960
Okutama-Action [46] 530874
Seadronessee [47] 913986
StanfordDroneDataset [48] 580223
UAVDark135 [49] 124980
UZH-FPV [50] 184458
VTUAV [51] 1576800
WebUAV3M [52] 376903
DTB70 [53] 47331
ERA [54] 39720
ITCVD [55] 27680
UAV123 [56] 11161
UAVHuman [57] 22549
UAV-ROD [58] 42088
UCLA Aerial Event dataset [59] 27370
Internet collection 228316
Total 5402975
During the pre-training phase, we adopt a combination of two techniques: Masked Image Modeling (MIM) and Contrastive Learning (CL). For MIM training, we use the standard SimMIM pre-training framework. The primary goal of this approach is to enhance the model’s ability to capture pixel-level semantic associations, thus initially equipping the model with feature extraction capabilities. However, as analyzed previously, solely using the MIM pre-training method for ARS foundational model training is insufficient for the model to learn the intrinsic object characteristics specific to ARS vision. Therefore, building upon the MIM pre-training, we propose a contrastive learning framework based on affine transformations. This involves generating positive and negative image pairs through affine transformations, which serve as the basis for the contrastive learning pre-training of the model.
The process of affine transformation of an image is completed through perspective transformation. Perspective transformation first defines the source point and the destination point, which are defined as:
[…]
The transformation relationship between the source point and the destination point used in the proposed method is:
[…]
Among them, $\alpha$ is the transformation coefficient, and the value range is randomly between 0.05 and 0.35. On this basis, the perspective transformation matrix can be calculated. The perspective transformation matrix is defined as a $3\texttimes 3$ matrix in the form of:
$M=\begin{bmatrix}m_{11}&m_{12}&m_{13}\\ m_{21}&m_{22}&m_{23}\\ m_{31}&m_{32}&1\end{bmatrix}$ (9)
The solution process of this transformation matrix is:
$\begin{cases}u_{i}=\frac{m_{11}x_{i}+m_{12}y_{i}+m_{13}}{m_{31}x_{i}+m_{32}y_{i}+1}\\ v_{i}=\frac{m_{21}x_{i}+m_{22}y_{i}+m_{23}}{m_{31}x_{i}+m_{32}y_{i}+1}\end{cases}~{}\text{for}~{}i=0,1,2,3$ (10)
Based on formulas (7) and (8), the above transformation matrix can be solved by the least square method. After obtaining the transformation matrix, the point $(x,y)$ on the source image can be mapped to the point $(u,v)$ on the destination image:
$\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=M\cdot\begin{bmatrix}x\\ y\\ 1\end{bmatrix}$ (11)
Through the aforementioned affine transformation process, we obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$. The source image $I_{i}$ and its corresponding transformed image $I_{i}^{\prime}$ form a positive sample pair, while the remaining sample pairs $I_{i}$ and $I_{j}^{\prime}$ (where $i\neq j$) form negative sample pairs. By calculating the contrastive loss, we guide the model to learn the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.
IV-D Downstream Fine-tuning
Figure 10: The proposed ARS-Adapter.
Figure 11: Qualitative results for the UAVID dataset when using different methods.
In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks. Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model. The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, as 10. The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking. In a fashion akin to AdaptFormer [60], the introduced ARS-Adapter module operates in parallel with the model. The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer. Therefore, the forward propagation is:
[…]
Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling. The Adapter in parallel with FFN introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
[…]
V Experiments
TABLE II: Performance Comparison on AID and RESISC-45 Datasets
Model AID (OA) RESISC-45 (OA)
(TR=20%/50%) (TR=10%/20%)
GASSL [61] 93.55/95.92 90.86/93.06
SeCo [33] 93.47/95.99 89.64/92.91
SatMAE [62] 95.02/96.94 91.72/94.10
RingMo [1] 96.90/98.34 94.25/95.67
RVSA [63] 97.03/98.50 93.93/95.69
DINO-MC [64] 95.16/97.09 -
TOV [65] 95.16/97.09 90.97/93.79
SSL4EO [66] 91.06/94.74 87.60/91.27
CMID [67] 96.11/97.79 91.84/94.07
CACo [68] 90.88/95.05 88.28/91.94
CROMA [69] 96.44/97.58 92.63/95.04
SatLas [70] 94.96/97.38 92.16/94.70
GFM [71] 95.47/97.09 -
Scale-MAE [72] 96.44/97.58 92.63/95.04
SkySense [2] 97.68/98.60 94.85/96.32
RingMo-Aerial 95.81/96.46 92.28/95.65
TABLE III: Compare our model with other methods on LEVIR-CD and CDD dataset
Methods LEVIR-CD CDD
F1 IoU F1 IoU
FC-Siam-Di [73] 86.31 83.31 70.61 54.57
FC-Siam-Conc [73] 83.69 76.77 75.11 60.14
IFNet [74] 88.13 82.93 90.30 71.91
STANet [75] 87.30 77.40 84.12 72.22
BIT [76] 89.31 89.37 88.90 80.01
SNUNet [77] 88.16 87.17 83.89 72.11
ChangeFormer [78] 90.40 88.80 89.83 81.53
RingMo [1] 89.53 82.38 90.90 84.09
RingMo-Aerial (Ours) 92.71 87.10 94.62 90.09
TABLE IV: Comp. on UAVID
Methods Building Tree Clutter Road Low vegetation Static car Moving car Human mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
DeeplabV3+ [79] 82.5 80.6 78.1 73.9 71.8 45.3 77.9 65.2 58.3 45.3 45.8 24.2 70.2 53.3 8.2 1.8 61.6 48.7
DANet [80] 81.8 81.2 80.2 72.3 69.3 46.1 78.2 63.6 56.9 46.2 48.3 23.8 70.8 54.2 8.6 3.0 61.8 48.8
ACNet [81] 82.3 81.6 80.8 72.6 68.9 45.3 79.4 64.1 57.3 46.5 48.1 23.2 71.2 55.0 8.8 3.2 62.1 48.9
OCRNet [82] 81.9 81.4 80.6 72.4 69.2 46.0 78.6 63.7 57.0 46.4 48.5 23.7 71.0 54.8 8.5 2.9 61.9 48.9
SETR [83] 82.3 78.9 78.3 71.2 66.9 45.8 76.2 63.8 60.6 43.7 45.6 22.6 70.1 54.8 7.8 1.6 61.0 47.8
Segformer [84] 83.6 81.2 80.3 74.1 70.3 46.3 76.8 64.6 61.0 46.2 45.3 24.6 70.3 56.1 12.2 5.3 62.5 49.8
CSwin [85] 86.3 82.8 80.6 75.6 68.9 49.2 79.2 66.3 61.2 47.8 48.3 26.9 70.6 55.0 16.5 10.8 64.0 51.8
UAVFormer [86] 88.5 81.5 82.3 76.2 67.5 48.8 78.6 67.1 63.8 48.5 49.6 28.8 71.2 62.2 19.3 12.5 65.1 53.2
BiSeNetV2 [87] - 81.6 - 76.0 - 61.2 - 77.1 - 61.3 - 38.5 - 66.4 - 15.4 - 59.7
Segmenter [88] - 84.4 - 76.1 - 64.2 - 79.8 - 57.6 - 34.5 - 59.2 - 14.2 - 58.7
CoaT [89] - 88.5 - 79.3 - 69.0 - 80.0 - 62.0 - 59.1 - 70.0 - 18.9 - 65.8
BANet [90] - 85.4 - 78.9 - 66.6 - 80.7 - 62.1 - 52.8 - 69.3 - 21.0 - 64.6
UNetFormer [91] - 87.4 - 80.2 - 68.4 - 81.5 - 63.5 - 56.4 - 73.6 - 31.0 - 67.8
RingMo(U)*[1] 90.9 85.4 82.6 74.3 73.1 56.8 81.6 73.7 71.9 58.2 69.9 58.3 75.4 67.5 27.4 21.0 71.6 61.9
RingMo-Aerial(U)* 92.6 83.8 88.5 77.2 78.4 65.5 87.5 80.2 73.3 60.7 72.5 65.3 84.4 71.6 34.8 26.9 76.5 66.4
RingMo-Aerial(A)* 92.4 83.5 89.0 77.7 77.2 63.5 85.8 79.4 75.2 60.4 60.5 53.4 85.7 72.7 28.5 24.7 74.3 64.4
RingMo-Aerial(M)* 93.2 85.7 89.8 79.0 79.0 66.3 87.8 82.0 76.8 62.0 72.5 65.5 88.9 76.6 41.3 32.7 78.7 68.7
    Swin-Base (U) means using the Swin-Base backbone plus UperNet segmentation algorithm. RingMo-Aerial (U), (A), (M) represent the RingMo aerial backbone with UperNet segmentation algorithm, ARS-Adapter and UperNet algorithm, and Mask2Former algorithm. Same below.
TABLE V: Comp. on UDD6
Methods Other Facade Road Vegetation Vehicle Roof mAcc mIoU
Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU Acc IoU
DeeplabV3+ [79] 92.6 81.8 88.6 72.3 85.3 71.2 83.6 70.8 78.9 62.9 92.4 80.2 86.9 73.2
DANet [80] 92.8 82.1 89.3 72.6 84.5 71.3 81.8 71.6 82.6 63.1 93.8 81.3 87.5 73.7
ACNet [81] 93.1 82.8 90.1 73.2 84.8 71.5 81.6 71.2 83.2 64.2 94.2 81.8 87.8 74.1
OCRNet [82] 92.7 81.9 90.2 73.4 84.3 71 81.9 71.8 83 63.6 94 81.6 87.7 73.9
SETR [83] 91.8 79.6 88.3 70.9 82.6 69.8 81.6 69.6 77.8 61.8 90.8 79.9 85.5 71.9
Segformer [84] 93 82.5 88.9 73.1 86.2 72.3 89.3 71.8 80.1 68.6 91.6 80.8 88.1 74.9
CSwin [85] 94.2 84.8 89.8 74.9 88.5 73.8 86.3 73.6 83.2 64.9 95.6 86.9 89.6 76.5
UAVFormer [86] 95.2 86.9 90.3 78.3 89.6 75.6 88.1 74.8 82.3 65.6 96.3 84.6 90.3 77.6
RingMo(U) [1] 91 80.4 89.4 74.6 85.5 73.2 89.6 78.4 77.8 62.3 91.1 82.9 87.4 75.3
RingMo-Aerial(U) 90.4 81.1 89.7 77.1 88.7 77.4 90.4 80.2 81.3 68.9 95.9 86.3 89.4 78.5
RingMo-Aerial(M) 90.7 80.9 90.2 77.4 88.4 77.1 90.5 80.8 82.1 69.4 96.3 87.8 89.7 78.9
TABLE VI: Comp. on Floodnet
Methods Building Road Water Tree Vehicle Pool Grass OA mIoU
Flooded Non-Flooded Flooded Non-Flooded
SegNet [92] 44.72 69.85 39.70 76.44 71.17 76.88 32.74 35.87 85.84 87.98 59.25
FCN8s 48.93 73.84 43.11 80.22 71.71 77.33 38.19 45.25 86.85 88.78 62.83
FCN16s 44.97 67.54 34.14 72.80 65.39 76.07 34.28 32.27 85.66 87.23 57.01
PSPNet 45.24 70.30 43.05 76.06 74.32 74.59 24.41 40.74 85.88 88.01 59.40
DeepLabV3 50.32 78.02 48.46 82.06 76.94 79.90 47.14 48.45 88.10 90.13 66.60
DeepLabV3+ [79] 47.62 74.46 43.12 79.07 68.35 76.36 35.47 36.69 85.60 87.79 60.75
UNet 49.95 77.58 43.90 81.90 73.46 78.88 43.30 49.30 87.54 89.56 65.09
HRNet 50.04 79.23 48.62 81.59 75.05 80.16 49.80 50.89 88.38 90.17 67.08
TransUNet [93] 48.97 73.67 43.11 79.09 72.86 77.75 26.27 35.02 86.47 88.82 60.36
RingMo(U) [1] 71.95 66.80 44.22 74.24 61.97 77.97 23.11 33.01 85.22 88.46 59.83
RingMo-Aerial(U) 74.48 77.06 55.78 83.00 70.35 81.41 53.96 58.08 87.04 90.40 71.24
RingMo-Aerial(M) 76.17 77.90 48.57 82.62 62.75 81.70 63.16 65.79 87.16 90.43 71.76
V-A Pre-train
In the pretraining phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized. Prior to pretraining, we initialize the model parameters using Kaiming initialization. Subsequently, we perform 100 epochs of masked image modeling (MIM) pretraining. During MIM pretraining, the input image size is set to 224, and random cropping and scaling are applied to the input images. The mask size is set to 32 pixels. The batch size is set to 2048, with a base learning rate of 2e-4, cosine annealing learning rate scheduling, and a warm-up period of 2 epochs. Following MIM pretraining, we continue with contrastive learning (CL) pretraining, jointly training for 36 epochs. The input image size remains 224, and similar random cropping and scaling are applied. The batch size is adjusted to 1024, with a base learning rate of 2e-4, cosine annealing learning rate scheduling, and a warm-up period of 1 epoch.
V-B Scene Classification
V-B1 Datasets and Evaluation Metrics
For semantic segmentation, the AID [94] and RESISC-45 [95] datasets are employed.
  • AID. The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification. It contains images collected from Google Earth and covers 30 different types of scenes.
  • NWPU-RESISC45. The NWPU-RESISC45 dataset is a publicly available REmote Sensing Image Scene Classification (RESISC) benchmark. The dataset contains 31,500 images covering 45 scene classes, with 700 images in each class.
In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively. In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively. The experiments use overall accuracy (OA) as a performance metric, which is the most commonly used performance metric in scene classification.
V-B2 Experiment Settings
The experiment of the scene classification is based on mmpretrain , Cascade RCNN [96] is employed as the classification framework. The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs). The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05. In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
V-B3 Comparison With the SOTA
The comparison results are shown in Table II. RingMo-Aerial combined with Cascade RCNN can achieve classification results far exceeding other methods.
V-C Change Detection
V-C1 Datasets and Evaluation Metrics
For semantic segmentation, the LEVIR [75] and CDD [98] datasets are employed.
  • LEVIR. The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset. It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5 meter resolution. Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels. Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases, respectively.
  • CDD. The CDD dataset is a comprehensive collection of bitemporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels. These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches. This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs. To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
V-C2 Experiment Settings
Our model adopts BIT with Swin Transformer as the backbone. The experiments use overall accuracy (OA) and F1 and IoU scores with regard to the change detection as the primary evaluation metrics. The model is trained with AdamW optimizer with a weight decay 0.01 and beta values equal to (0.9, 0.999). The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size of 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
V-C3 Comparison With the SOTA
Figure 12: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset.
The comparison results are shown in Table III. RingMo-Aerial combined with BIT can achieve change detection results far exceeding other methods. In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection. When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated in our architecture.
The comparison of visual results12 on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks. The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images. In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
RingMo-Aerial’s commendable performance not only makes it a key player in the field, but also highlights the importance of the impact of foundation model innovation. These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
V-D 3D-Reconstruction
V-D1 Datasets and Evaluation Metrics
For 3D-Reconstruction, the and WHU [99] datasets are employed.
  • WHU. The WHU dataset is a synthetic aerial dataset designed for large-scale multi-view stereo reconstruction tasks. It was generated from a highly accurate 3D surface model and includes a complete set of aerial images as well as cropped sub-image sets specifically for deep learning applications. The images in the dataset cover an area of approximately 6.7 × 2.2 square kilometers in Meitan County, Guizhou Province, China, with a ground resolution of 0.1 meters. The WHU dataset provides crucial data support for the study of large-scale Earth surface and urban reconstruction and can be utilized for training and testing multi-view stereo matching algorithms.
  • LuoJia-MVS. The LuoJia-MVS dataset comprises 7,972 five-view images, each with a spatial resolution of 10 cm, accompanied by pixel-wise depth information and precise camera parameters. These images were generated from a highly accurate digital surface model (DSM) constructed using thousands of stereo aerial images.
During the experiment, we used the WHU dataset[99] and the LuoJia-MVS dataset[100] to verify the effectiveness of our proposed method. WHU dataset[99] is the first large-scale aerial multi-view stereo dataset for remote sensing. It was collected from Meitan, Guizhou, China, and was obtained by taking large-scale photos with a tilted five-eye camera on a drone platform. The area of the entire dataset is approximately $6.7\times 2.2km^{2}$, and the image resolution is about 0.1m. In addition, the tilt angle of each camera to the central camera is 40°. The WHU dataset gathers high-rise buildings, mountains, forests, and factory scenes, providing a good benchmark for remote sensing multi-view stereo. LuoJia-MVS[100] dataset is a large-scale multi-view stereo dataset collected from a variety of landform types. It was collected in a part of Baiyun, Guiyang, Guizhou, China. The image size of the entire dataset is $768\times 384$, and the image resolution is about 0.1m. The data consists of five views and pixel-by-pixel depth labels. Land types cover cultivated land, forest land, industrial areas, residential areas, etc.
V-D2 Experiment Settings
Our work is coded based on the PyTorch framework. The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1. During the training process, our method uses the Adam optimizer[101] for parameter optimization, and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001. In addition, our method is consistent with the existing multi-stage networks[102]. First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively. Then, deep optimization is performed through three stages of deep hypotheses, and the number of hypotheses is 48, 32, and 8, respectively. During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
V-D3 Comparison With the SOTA
TABLE VII: Performance comparisons on the LuoJia-MVS and WHU datasets.
Number of Views Model LuoJia-MVS WHU
MAE $<$0.6m $<$3-interval MAE $<$0.6m $<$3-interval
Three views * PatchmatchNet[103] 0.252 92.7 87.2 0.173 96.5 94.8
Fast-MVSNet[104] 0.194 95.7 92.0 0.184 95.5 94.1
MVSNet[105] 0.172 96.1 92.4 0.190 95.0 94.3
R-MVSNet[106] 0.177 96.0 93.5 0.183 95.3 93.5
RED-Net[99] 0.109 98.2 96.9 0.112 98.1 97.9
Cas-MVSNet[102] 0.103 98.4 97.1 0.111 97.7 97.6
RingMo-Aerial 0.095 98.7 97.8 0.121 97.8 97.5
Five views * PatchmatchNet[103] 0.283 90.4 84.1 0.160 96.9 95.0
Fast-MVSNet[104] 0.357 84.6 74.9 0.157 96.1 95.6
MVSNet[105] 0.270 91.2 81.8 0.160 95.8 95.5
R-MVSNet[106] 0.259 92.3 86.7 0.173 95.4 93.8
RED-Net[99] 0.156 94.9 90.5 0.104 98.1 97.9
Cas-MVSNet[102] 0.141 97.9 95.4 0.095 97.8 97.8
RingMo-Aerial 0.122 98.1 96.5 0.091 98.0 97.9
    ‘Three Views’ and ‘Five Views’ represent the number of views for the input of the network is three and five, respectively.
Table VII provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets. Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error. On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet, and improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs. 0.141), 0.2% (98.1% vs. 97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a large improvement. In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%. It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
V-E Semantic Segmentation
V-E1 Datasets and Evaluation Metrics
For semantic segmentation, the UAVID [107], UDD6 [108] and FloodNet[109] datasets are employed.
  • UAVID. The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid and 15 test), and each sequence contains 10 pictures. In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap. A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
  • UDD6. The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images. All of the images are 4K resolution. In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap. The training set is used for training and the validation set for testing.
  • FloodNet. The FloodNet dataset contains 500 high-resolution remote sensing images with a resolution of 4K. The dataset is divided into training, validation, and test sets, with the training set containing 300 images, the validation set containing 100 images, and the test set containing 100 images. Images are annotated as flood and non-flood areas, supporting multi-class flood impact assessment. All images are cropped into 1024×1024 pixel slices with a 256-pixel overlap in the experiment, facilitating the training and evaluation of machine learning models.
In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc) and average accuracy (aAcc).
V-E2 Experiment Settings
The experiment of the semantic segmentation is based on mmsegmentation [110], UperNet [111] and Mask2Former [112] are employed as the segmentation framework. The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used). The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05. In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
V-E3 Comparison With the SOTA
The comparison results on the UAVID dataset are shown in Table IV, results on the UDD6 dataset are shown in Table V, results on the FloodNet dataset are shown in Table VI. RingMo-Aerial combined with Mask2Former can achieve segmentation results far exceeding other methods. Experiments on UperNet (RingMo-Aerial vs. Swin-Base) also show that RingMo-Aerial is a more suitable basic model for ARS segmentation tasks.
Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks. As illustrated in the Figure11, RingMo-Aerial consistently produces superior segmentation maps. Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures. In particular, the RingMo-Aerial + UperNet + ARS-Adapter and RingMo-Aerial + Mask2Former methods demonstrate higher accuracy and clarity in the segmentation of small targets (e.g., stationary and moving vehicles), with enhanced boundary and detail capture. This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
V-F Object Detection
TABLE VIII: Comp. on VisDrone-DET
Methods Backbone mAP mAP50 mAP75
Light-RCNN[113] ResNet-50 16.5 32.8 15.1
CornerNet[114] ResNet-50 17.4 34.1 15.8
RetinaNet[115] ResNet-50 11.8 21.4 11.6
RetinaNet[115] Swin-Base 22.6 38.1 23.5
YOLOv4[116] CSPDarknet-53 - 38.7
Cascade R-CNN[96] ResNet-50 16.1 31.9 15.0
Cascade R-CNN[96] Swin-Base 28.4 45.2 30.1
CEASC[117] ResNet-18 28.7 50.7 28.4
SOD-YOLOv7[118] - - 53.2
MFFSODNet[119] - - 45.5
GLSAN[120] ResNet-50 25.8 51.5 22.9
QueryDet[17] ResNet-50 28.3 48.1 28.7
VistrongerDet[121] ResNet-50 33.9 57.3
DMNet[122] ResNet-101 29.4 49.3 30.6
HRDNet[123] ResNet-101 28.3 49.2 28.1
ClusDet[124] ResNet-101 26.7 50.4
ClusDet[124] ResNeXt-101 28.4 53.2
SDPDet[125] ResNet-101 34.2 57.8 34.9
AMRNet[126] ResNet-101 31.7 52.6 33.0
OGMN[11] ResNeXt-101 35.0 59.7 35.8
DDQ-DETR[127] Swin-Base 36.1 57.1 37.2
RingMo-Aerial (C) * RingMo-Aerial 31.1 51.8 32.5
RingMo-Aerial (A) * RingMo-Aerial 31.0 51.4 32.1
RingMo-Aerial (D) * RingMo-Aerial 38.6 63.3 39.4
    RingMo-Aerial (C), (A), (D) represent the RingMo aerial backbone with Cascade RCNN detection algorithm, ARS-Adapter and Cascade RCNN algorithm, and DDQ DETR algorithm. Same below.
TABLE IX: Comp. on UAVDT
Methods Backbone mAP mAP50 mAP75
GLSAN[120] ResNet-50 17.0 28.1 18.8
DMNet[122] ResNet-50 14.7 24.6 16.3
ClusDet[124] ResNet-50 13.7 26.5 12.5
AMRNet[126] ResNet-50 18.2 30.4 19.8
DSHNet[128] ResNet-50 17.8 30.4 19.7
EfficientDet[129] Efficient-B7 13.1 31.8 10.9
YOLOv4[116] CSPDarknet-53 13.4 30.6 11.3
SDPDet[125] ResNet-50 20.0 32.0 23.1
QueryDet[17] ResNet-50 14.3 27.2 16.6
Cascade R-CNN[96] Swin-Base 19.0 31.4 19.2
OGMN[11] ResNet-50 20.9 34.5 23.2
OGMN[11] ResNet-101 24.2 39.9 26.8
RingMo-Aerial (C) RingMo-Aerial 21.4 33.1 21.9
RingMo-Aerial (D) RingMo-Aerial 25.2 39.9 27.1
TABLE X: Comp. on ShipDataset
Methods Backbone mAP mAP50 mAPs
RetinaNet[115] RetinaNet 76.0 98.6 40.4
UAV-YOLO[130] DarkNet-53 86.3 99.5 -
TPH-YOLOv5[131] CSPDarknet53+SPP 89.6 99.2 -
YOLOv7[132] EfficientNet-B0 88.7 99.6 -
NanoDet-Plus[133] MobileNetV2 83.1 96.4 41.8
YOLOX-L[134] CSPDarkNet-53 68.5 93.5 27.6
Center-Net++[135] ResNet-101 88.8 99.0 70.8
YOLOX-Tiny[134] CSPDarkNet-53 71.1 96.4 30.1
PP-PicoDet-L[136] ESNet 89.7 98.9 67.3
Cascade R-CNN[96] Swin-Base 88.3 99.0 62.6
RingMo-Aerial(C) RingMo-Aerial 83.7 99.1 72.3
TABLE XI: Comp. on DroneVehicle
Methods Modality Backbone mAP50
Faster R-CNN(OBB)[137] RGB ResNet101 44.6
RoITransformer[138] RGB ResNet101 47.9
YOLOv7[132] RGB EfficientNet-B0 68.5
$S^{2}$et[139] RGB RetinaNet 57.3
Oriented R-CNN[140] RGB ResNet101 62.3
Cascade R-CNN[96] RGB Swin-Base 69.5
Faster R-CNN(OBB)[137] IR ResNet101 44.6
RoITransformer[138] IR ResNet101 59.2
YOLOv7[132] IR EfficientNet-B0 66.7
$S^{2}$et[139] IR RetinaNet 64.8
Oriented R-CNN[140] IR ResNet101 65.5
Cascade R-CNN[96] IR Swin-Base 69.8
UA-CMDet[141] RGB+IR EfficientNet-B0 64.0
Halfway Fusion(OBB)[142] RGB+IR Halfway Fusion 68.2
CIAN(OBB)[143] RGB+IR VGG16 70.2
Dual-YOLO[144] RGB+IR Dual-YOLO 71.5
Cascade R-CNN[96] RGB+IR Swin-Base 69.8
RingMo-Aerial(C) RGB+IR RingMo-Aerial 71.6
Figure 13: Qualitative results for the VisDrone-DET dataset when using different methods.
TABLE XII: Comp. on SARData
Methods Backbone mAP mAP50
RetinaNet[115] RetinaNet 22.9 64.3
YOLOX-S[134] CSPDarkNet-53 34.5 85.0
Cascade R-CNN[96] ResNet101 34.5 81.3
RingMo-Aerial(A) * RingMo-Aerial 30.9 77.6
RingMo-Aerial(Y) * RingMo-Aerial 34.0 81.5
RingMo-Aerial(C) * RingMo-Aerial 34.9 82.4
    RingMo-Aerial (A), (Y), (C) represent the RingMo aerial backbone with ARS-Adapter and YOLOX-S algorithm, YOLOX-S algorithm, and Cascade RCNN algorithm.
V-F1 Datasets and Evaluation Metrics
To conduct a comprehensive comparison of object detection methods, we selected four unimodal benchmark datasets, including the VisDrone-DET[145], UAVDT[146] and ShipDataset[147], as well as one multimodal benchmark dataset, DroneVehicle[141] and our collection of SAR airborne remote sensing data, to demonstrate the performance of the representative methods discussed in this paper.
  • VisDrone-DET. The Visdrone-DET dataset is the detection task part of the VisDrone dataset. The resolution of the images is about 2k. The RingMo-Aerial experiment crops the original image to 640×640 and uses multi-scale cropping and resizing for enhancement. The training set and validation set are used for training, and the labeled test-dev set is used for testing.
  • UAVDT. The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with the average resolution of 1080 × 540 pixels. The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
  • ShipDataset. The ShipDataset, captured at a resolution of $3840\times 2160$, encompasses five distinct scenarios in Shanghai, China, characterized by varying shooting angles and lighting conditions. This dataset has $96.97\%$ small ships. The images within this dataset predominantly feature congested ship traffic. The RingMo-Aerial experiment uses multi-scale resizing for enhancement.
  • DroneVehicle. The DroneVehicle dataset is a large-scale dataset specifically designed for drone-based RGB-Infrared cross-modality vehicle detection tasks. It includes 28,439 pairs of RGB-Infrared images, totaling 56,878 images, covering various urban scenarios such as roads, residential areas, and parking lots under different lighting conditions, including day, night, and dark night. The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
  • SARData. To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside. This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection. The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels. The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
In our object detection experiments, the evaluation metrics employed are mean average precision (mAP), mAP at an IoU threshold of 50 (mAP50), and mAP at an IoU threshold of 75 (mAP75). For datasets containing smaller objects, we also utilize mAPs, which represents the mean average precision for detecting small objects (with pixels smaller than 32×32).
V-F2 Experiment Settings
The experiment of the object detection is based on mmdetection [148], Cascade RCNN [96], YOLOX-S[134] and DDQ DETR [127] are employed as the detection framework. The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used). That for YOLOX-S spanned 36 epochs with a batch size of 4 total (4 per GPU and 1 NVIDIA A40 GPU are used). That for DDQ DETR spanned 36 epochs with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used). The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05. In addition, the learning rate is decreased by a factor of 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
TABLE XIII: Object Tracking: Comp. on Visdrone-MOT
Methods MOTA IDF1
SiamMOT[149] 31.9 48.3
ByteTrack[150] 35.7 37.0
UAVMOT[151] 36.1 51.0
OCSORT[152] 39.0 50.4
MOTR[153] 22.8 41.4
TrackFormer[154] 24.0 30.5
FOLT[155] 42.1 56.9
U2MOT[156] 52.3 69.0
RingMo[1] 48.4 63.1
RingMo-Aerial 52.6 67.8
TABLE XIV: Object Tracking: Comp. on AIR-MOT
Methods MOTA IDF1
CKDNet+SMTNet[157] 64.2 64.9
ByteTrack [150] 56.1 73.5
OCSORT[158] 35.6 53.8
StrongSORT[159] 36.1 54.1
CenterTrack[160] 66.3 79.3
TGram[161] 65.7 66.8
RSMOT[162] 67.0 80.4
CFTrack[163] 70.8 82.7
RingMo-Aerial 68.6 81.3
TABLE XV: Object Tracking: Comp. on AIR-HSAO
Methods MOTA IDF1
TraDeS[164] 59.6 75.1
CenterTrack[160] 60.3 74.5
DSFNet[165] 60.2 75.8
FairMOT [166] 54.8 71.3
CTracker[167] 51.1 65.1
ByteTrack[150] 60.3 74.5
MGTrack[168] 63.5 77.4
RingMo-Aerial 60.7 74.4
Figure 14: Qualitative results for DroneVehicle-IR when using different methods
Figure 15: Qualitative results for SARData when using different methods
V-F3 Comparison With the SOTA
An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance. On unimodal datasets such as VisDrone-DET (as shown in Table VIII) and UAVDT (as shown in Table IX), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods. The experimental results further demonstrate that RingMo-Aerial, when compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Table X), and PP-PicoDet-L follows, while YOLOX-L and YOLOX-Tiny show the worst results. Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
For the multimodal dataset, DroneVehicle (as shown in Table XI), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods. Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial. This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications. Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Table XII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as YOLOX-S, Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance. Particularly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and an mAP50 of 84.4%, highlighting its superiority in SAR modality. These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery. Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments. These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields. For optical images, as shown in Figure 13, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets. In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups. Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms. RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images. As shown in Figures 14 and 15, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence. The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images. RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for aerial remote sensing.
V-G Object Tracking
V-G1 Datasets and Evaluation Metrics
For Object Tracking, following aerial dataset is employed.
  • VisDrone-MOT. The VisDrone-MOT[169] dataset consists of 96 sequences, including 5 class such as pedestrian, car, van, bus and truck. The VisDrone-MOT train and val set are used for training and test set is used for test.
In addition to the aerial dataset, we conducted experiments on two satellite remote sensing datasets.
  • AIR-MOT. The AIR-MOT dataset[163], derived from the JiLin-1 satellite, is designed for multi-class multiple object tracking (MOT). It includes two categories: airplanes and ships. The dataset comprises 152 videos, with frame rates ranging from 5 to 10 frames per second, a spatial resolution of 1–1.2 meters, and a pixel resolution of 1920 × 1080. The target categories in this dataset move relatively slowly.
  • AIR-HSAO. The AIR-HSAO dataset[168] is designed for tracking high-speed aerial objects, derived from the JiLin-1 satellite. It contains 199 videos, though fewer instances are present due to the rapid movement of these objects. The dataset spans 20 diverse regions, including urban areas, harbors, and airports, offering complex backgrounds that challenge object detection and tracking.
In our experiment of object tracking, Multiple Object Tracking Accuracy (MOTA) and Identification Flatness (IDF1) are used as evaluation metrics.
V-G2 Experiment Settings
In the proposed experiment, object tracking is predicated on the detection model of RingMo-Aerial on VisDrone-DET. Specifically, the detection model trained by RingMo-Aerial will undergo an additional 5 epochs of training on the VisDrone-MOT dataset, during which the backbone network parameters will be frozen to ensure consistency. The training will be conducted with a batch size of 8, distributed as 2 per GPU across 4 NVIDIA A40 GPUs. Throughout the training phase, original images from the VisDrone-MOT dataset will be processed, accompanied by multi-scale augmentation to enhance robustness. Upon completion of training on the VisDrone-MOT dataset, the detection model will be subjected to inference on the VisDrone-MOT test set using the MOT tools provided by mmdetection [148]. Subsequently, the model’s multi-object tracking performance will be quantified and analyzed.
V-G3 Comparison With the SOTA
The comparison results on the VisDrone-MOT dataset are shown in Table XIII. RingMo-Aerial can achieve the best tracking results far exceeding other methods. In the AIR-MOT dataset (as shown in TableXIV), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack. Meanwhile, in the AIR-HSAO dataset (as shown in TableXV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods. Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
Figure 16: Visualization Results on AIR-MOT: Four video sequences are selected, where each target in the sequences is indicated with different colors. The category and target ID are displayed above the bounding boxes.
Futhermore, the visualization results of the AIR-MOT sequences (as shown in Figure16) clearly demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios. The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions. This visualization provides strong qualitative evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in aerial remote sensing videos.
V-H ARS-Adapter
TABLE XVI: Comp. between fine-tuning methods
Methods Parameters UAVID VisDrone-DET
mAcc mIoU mAP mAP50
Froezn 0% 61.2 49.3 21.5 36.2
Full 100% 76.5 66.4 31.1 51.8
Adapter 0.39% 68.9 57.7 28.9 49.1
AdaptFormer 0.39% 71.3 62.1 30.3 50.8
ARS-Adapter 0.41% 74.3 64.4 31.0 51.4
As previously discussed, this paper proposes the ARS-Adapter, a parameter-efficient fine-tuning method tailored for ARS vision. This section focuses on comparing the performance of the ARS-Adapter with other fine-tuning methods in ARS vision applications. The comparisons are conducted using the UAVID and VisDrone-DET datasets, employing UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively. The evaluation aims to highlight the efficacy and advantages of the ARS-Adapter in enhancing ARS vision tasks compared to conventional fine-tuning techniques.
[…]
V-I Ablation Experiments
Figure 17: The ablation experiment on UAVID dataset and UperNet is used as the segmentation head. RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
TABLE XVII: Ablation Experiments
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo 71.6 61.9 28.4 45.2
RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
TABLE XVIII: Ablation Experiments with the proposed ARS-Adapter
UAVID Visdrone-DET
mAcc mIoU mAP mAP50
RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
The ablation experiments in this study are conducted utilizing the UperNet segmentation model on the UAVID dataset and the Cascade RCNN detection model on the Visdrone-DET dataset. The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL. This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
The comparison is shown in Figure 17 and Table XVII. The experimental findings reveal a progressive enhancement in convergence speed and performance across the three foundation models. Notably, the RingMo-Aerial model trained with both MIM and CL exhibits the swiftest convergence and attains the optimal performance. This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pretraining strategy, affirming their mutually reinforcing nature and consistent effectiveness.
In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter. Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL. Comparative results on the UAVID and VisDrone-DET datasets are presented in Table XVIII. The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
VI Conclusion
The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks. Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein.
[…]
Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks. This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks.
In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research. Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment. Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.

### v3

@@
-RingMo-Aerial: An Aerial Remote Sensing Foundation Model With A Affine Transformation Contrastive Learning
+RingMo-Aerial: An Aerial Remote Sensing Foundation Model With Affine Transformation Contrastive Learning
 Abstract
@@
 Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks.
-This indicates the practicality and effectiveness of RingMo-Aerial in enhancing the performance of ARS vision tasks.
+This indicates the practicality and efficacy of RingMo-Aerial in enhancing the performance of ARS vision tasks.
 Index Terms:
-Aerial Remote Sensing (ARS), foundation model, contrastive learning(CL), pretraining, self-supervised.
-I Introduction
+Aerial Remote Sensing (ARS), Foundation Model, Contrastive Learning(CL), Affine Transformation
 Figure 1: Number of Aerospace Remote Sensing and Aerial Remote Sensing papers in Web of Science in the last decade.
 The green bars represent aerospace remote sensing, while the orange bars represent aerial remote sensing.
-Although research on aerial remote sensing has grown, the number of papers remains significantly lower than that of aerospace remote sensing.
-Figure 2: Illustration of multi-view, multi-resolution and occlusion of aerial remote sensing images:: (a) Multi-view imagery shows images captured from different angles, including tilt, top, and flat views; (b) Multi-resolution imagery demonstrates how varying low, medium, and high resolutions affect the level of detail; (c) Target occlusion compares cases where the target is partially occluded versus not occluded.
-These characteristics highlight the complexity of aerial remote sensing imagery in different scenarios.
-Figure 3: Schematic Comparison of Aerospace Remote Sensing and Aerial Remote Sensing Image Features.
+Although research on ARS has grown, the number of papers remains significantly lower than that of aerospace remote sensing.
+Figure 2: Illustration of multi-view, multi-resolution, and occlusion of ARS images: (a) Multi-view imagery shows images captured from different angles, including tilt, top, and flat views; (b) Multi-resolution imagery demonstrates how varying low, medium, and high resolutions affect the level of detail; (c) Target occlusion compares cases where the target is partially occluded versus not occluded.
+These characteristics highlight the complexity of ARS imagery in different scenarios.
+1 Introduction
+Aerial Remote Sensing (ARS) is an essential method of data acquisition, distinct from aerospace remote sensing (ApRS) in its flexibility, as it is typically free from constraints such as revisit cycles, satellite sensor incidence angles, and timing limitations.
+ARS sensors, usually mounted on Unmanned Aerial Vehicles (UAVs), airplanes, or balloons, offer advantages over satellite-based remote sensing by quickly reaching target areas and capturing real-time, multi-perspective observations at various resolutions.
+This capability enables ARS to provide detailed information in applications ranging from emergency response to environmental monitoring and national defense.
+Thus, enhancing intelligent information extraction from ARS images holds significant academic and practical value.
+Despite ARS’s distinct advantages and applications, research efforts in intelligent remote sensing interpretation are still largely focused on ApRS imagery.
+As shown in Fig.1, although interest in ARS has grown rapidly in recent years, studies in its intelligent interpretation remain relatively limited, with publications in this area comprising less than 10% of ApRS-related work.
+Unlike ApRS, ARS imagery is characterized by unique physical properties, including multi-view perspectives, varied resolutions, and occlusion effects, due to its flexible observational angles and sensor characteristics, as shown in Fig.2.
+In Fig.2(a), various perspectives are depicted, including oblique views (i and ii), a vertical view (iii), and a horizontal view (iv).
+Unlike ApRS, which predominantly captures vertical views, ARS imagery often incorporates oblique angles, providing more dimensional perspectives of targets.
+However, this also results in distant objects appearing smaller and more densely arranged, increasing the complexity of image interpretation.
+Fig.2(b) shows images at low, medium, and high resolutions (i, ii, iii, respectively), reflecting the diverse resolutions produced by ARS sensors at varying altitudes and speeds.
+Such differences in resolution affect target clarity and detail across images, placing higher demands on model adaptability.
+Additionally, significant occlusion is common in aerial imagery, as demonstrated in Fig.2(c), where some targets are partially or fully obscured by other objects, further complicating image interpretation.
+These challenges highlight the need for specialized intelligent models tailored to ARS applications.
+While recent foundational models in computer vision and RS have shown potential, they are often inadequate for ARS due to its unique requirements.
+These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
+However, current foundational model research primarily focuses on ApRS imagery, as evidenced by studies like RingMo [1] and Skysense [2].
+This underscores the need for a dedicated pre-trained foundational model specifically designed for ARS.
+Adapting existing foundational models for ARS involves addressing several critical challenges:
+(1) Challenges in Pre-training: Pre-training generally relies on large-scale datasets for self-supervised learning, primarily using Masked Image Modeling (MIM) and Contrastive Learning (CL).
+MIM, employed by models like MAE [3] and SimMIM [4], masks parts of images for reconstruction to enhance feature extraction, while CL, seen in models like MOCO [5], leverages positive and negative sample pairs to learn relational features.
+Both methods have merits but also limitations; MIM is widely used across visual tasks, as RingMo[1].
+However, ARS images vary significantly in features due to differences in perspective and resolution.
+As shown in Fig.3, the vertical view displays structured, continuous features with broad areas of color consistency, containing abundant low-frequency information.
+In contrast, the oblique view, capturing building sides and heights, offers dynamic perspectives with high detail variability and lower low-frequency content.
+Current methods, which often focus on reconstructing local image information, struggle to capture these unique variations in ARS imagery.
+(2) Challenges in Fine-tuning: The traditional approach of global fine-tuning of models faces significant challenges, mainly due to the fact that fine-tuning may lead to overfitting, especially in ARS.
+Unlike single viewpoints and resolutions that are common in spatial remote sensing, ARS models are more likely to be overfitted with specific viewpoints and resolutions for specific downstream tasks.
+This overfitting reduces the ability of the model to generalize to different viewpoints and resolutions in various downstream tasks.
+Figure 3: Schematic Comparison of ApRS and ARS Image Features.
 Satellite images contain more low-frequency information with continuous features, while aerial images have more detailed variations and relatively less low-frequency information.
-Aerial Remote Sensing (ARS) is a significant method of information acquisition that, unlike satellite remote sensing, is typically not constrained by revisit cycles, satellite sensor incidence angles, or other observation timing and viewing angle limitations.
-Aerial remote sensing sensors are commonly mounted on Unmanned Aerial Vehicles (UAVs), airplanes, or balloons.
-Compared to remote sensing satellites, these aerial platforms can swiftly reach the observation area and conduct comprehensive observations from multiple perspectives and at various resolutions, thereby capturing real-time and detailed information about the target area.
-Consequently, aerial remote sensing finds extensive applications in emergency response, environmental monitoring, and national defense security.
-Therefore, research on enhancing aerial remote-sensing images’ intelligent information extraction capabilities holds significant academic and practical value.
-However, despite the advantages and value of aerial remote sensing imagery, our research reveals that the current focus in intelligent remote sensing interpretation remains predominantly on space-borne remote sensing.
-As shown in Figure 1, although research on aerial remote sensing has seen rapid growth in recent years, studies related to its intelligent interpretation remain relatively scarce.
-The number of published papers in this area accounts for less than 10% of space-borne remote sensing.
-Compared to space-borne remote sensing, aerial remote sensing possesses distinct physical characteristics, such as multi-view, multi-resolution, and occlusion effects, owing to its unique observational perspectives and sensor properties, as illustrated in Figure 2.
-In Figure 2(a), various perspectives are depicted, including oblique views (i and ii), a vertical view (iii), and a horizontal view (iv).
-Unlike the predominantly vertical perspective of satellite remote sensing, aerial remote sensing imagery captures surface information from oblique angles, enhancing the three-dimensional perception of the targets.
-However, this also results in distant objects appearing smaller and more densely arranged, increasing the complexity of image interpretation.
-Figure 2(b) compares images of different resolutions: low-resolution (i, iv), medium-resolution (ii, v), and high-resolution (iii, vi).
-These multi-resolution characteristics result from the varying data resolutions of aerial sensors operating at different flight altitudes and speeds.
-The variation in resolution affects the clarity and detail representation of the same target across different images, thereby placing higher demands on the adaptability of image interpretation models.
-Furthermore, aerial imagery is often characterized by significant occlusion, as shown in Figure 2(c).
-In this Figure, some targets are partially or fully occluded by other objects (i and iii illustrate occluded scenarios, while ii and iv show non-occluded scenarios).
-Such occlusions further complicate the interpretation of ARS images.
-Given the challenges mentioned above, it is difficult to design specialized intelligent models for aerial remote sensing that are tailored to different applications.
-Existing foundational models in computer vision and remote sensing vision are no longer sufficient to meet the demands of aerial remote sensing applications.
-The recently developed foundational models, which have gained significant attention in remote sensing interpretation, are considered an effective solution to these challenges.
-These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
-However, existing foundational model studies in the remote sensing domain primarily focus on the interpretation of spaceborne remote sensing imagery, as evidenced by studies like RingMo[1] and Skysense[2].
-This underscores the pressing need to develop a pre-trained foundational model specifically tailored for aerial remote sensing to address its unique application requirements.
-When adapting current foundational models to aerial remote sensing, several critical challenges persist:
-(1) Issues in the Pre-training Phase: The pre-training phase typically relies on large-scale datasets for self-supervised learning, primarily adopting two paradigms: Masked Image Modeling (MIM) and Contrastive Learning (CL).
-MIM, as exemplified by models like MAE [3] and SimMIM [4], involves applying masks to images and training the model to reconstruct the masked parts, thereby enhancing the model’s feature extraction capabilities.
-On the other hand, CL, as seen in models like MOCO [5], improves feature extraction by creating positive and negative sample pairs and training the model to learn the relationships between these pairs.
-Both paradigms have advantages and limitations, with MIM being more widely utilized in most visual scenarios.
-For instance, Swin Transformer and RingMo both utilize MIM-based pre-training methods.
-However, for aerial remote sensing imagery, the characteristics of images at different perspectives and resolutions differ significantly at the feature level.
-As shown in Figure 3, vertical images tend to exhibit more structured and continuous features, often characterized by large areas of color and texture consistency, with fewer variations in detail.
-Such images typically contain abundant low-frequency information due to their broad display range and relatively uniform visual presentation.
-In contrast, oblique images capture the sides of buildings and their height information, offering a more dynamic and three-dimensional perspective.
-These images are rich in detail variation and contain relatively less low-frequency information.
-The current self-supervised learning methods rely on reconstructing local information in images and struggle to capture these variations effectively.
-(2) Issues in the Fine-tuning Phase: The traditional approach of fine-tuning the entire model presents significant challenges, primarily due to the time-consuming nature of the process.
-Additionally, since the foundational model is pre-trained on a diverse set of downstream tasks, fine-tuning the model globally for a specific downstream task can lead to overfitting, especially in aerial remote sensing.
-Unlike the single perspective and resolution typically encountered in spaceborne remote sensing, aerial remote sensing models are more prone to overfitting to specific perspectives and resolutions of a particular downstream task.
-This overfitting can reduce the ability of the model to generalize to different perspectives and resolutions in various downstream tasks.
-To address these challenges, this study introduces RingMo-Aerial, the world’s first foundational large model designed for aerial remote sensing tasks, based on the existing spaceborne remote sensing foundational model, RingMo.
-In response to the unique characteristics and difficulties of aerial remote sensing tasks, we propose several novel theories and methods, as shown in Figure4.
-First, we introduce Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) to address the multi-scale and occlusion issues caused by the oblique perspective in aerial remote sensing images.
-Unlike traditional multi-head self-attention mechanisms, FE-MSA incorporates an additional image patch expansion layer, ensuring that dense small objects are captured by more tokens when patches are divided into tokens for the Transformer model.
-Furthermore, FE-MSA leverages the Swin Transformer’s window attention and shifted window mechanisms.
-These improvements significantly enhance the model’s ability to detect dense small objects and occluded targets.
-Second, to solve the problem of detecting and recognizing small distant targets in oblique perspectives, we propose a Contrastive Learning Affine Framework (CLAF) that scales and enlarges distant targets through affine transformations to construct positive samples.
-By inputting the transformed and original images as sample pairs into the model, this method guides the model to learn the unique visual characteristics of ARS images, improving its ability to recognize small targets at a distance.
-In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic aerial remote sensing foundational models, called ARS-Adapter.
+Figure 4: Comparison of model frameworks between RingMo and RingMo-Aerial.
+(a) RingMo uses SimMIM for pre-training to perform image reconstruction.
+(b) RingMo-Aerial builds upon this by introducing affine transformation and contrastive learning, enhancing its ability to process multi-view images.
+In the fine-tuning stage, the ARS adapter is introduced to adjust a small number of parameters, further enhancing the model’s generalization ability for downstream tasks.
+To address the challenges above, we introduces RingMo-Aerial, the first foundational large model explicitly designed for ARS tasks.
+Building on the existing ApRS foundational model RingMo, RingMo-Aerial addresses the unique characteristics of ARS tasks through several novel approaches, as illustrated in Fig.4.
+First, we propose the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module, designed to address the multi-scale and occlusion challenges inherent to oblique perspectives in ARS imagery.
+Unlike traditional multi-head attention mechanisms, FE-MSA includes an image patch expansion layer to capture dense, small objects as tokens for the Transformer model, leveraging Swin Transformer’s window and shifted window attention.
+This enhances the model’s ability to detect densely packed and occluded objects.
+Second, to improve detection and recognition of distant small targets in oblique views, we propose the Contrastive Learning Affine Framework (CLAF).
+CLAF uses affine transformations to scale and enlarge distant targets, constructing positive samples to guide the model in learning ARS-specific visual features.
+In addition, during the fine-tuning phase, we propose an efficient fine-tuning method for basic ARS foundational models called ARS-Adapter.
 Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting.
 ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter[6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling.
-This approach optimizes the performance of the aerial remote sensing foundational model, making it more adaptable and effective across various aerial remote sensing vision tasks.
-In conclusion, the core contributions of this paper are as follows:
-  • Addressing the Characteristics and Challenges of Aerial Remote Sensing Imagery: This paper introduces RingMo-Aerial, the world’s first foundational model designed for aerial remote sensing imagery.
-RingMo-Aerial supports various tasks across different perspectives, resolutions, and modalities, including scene classification, semantic segmentation, object detection, object tracking, change detection, and 3D reconstruction.
-It can be widely applied across multiple different scenarios.
-  • FEMSA: By incorporating a patch expansion layer and integrating the window attention mechanism, FEMSA effectively enhances the detection capability of small, dense targets in aerial remote sensing images.
-  • CLAF: Utilizing affine transformations, CLAF guides the model to learn the features of aerial remote sensing imagery captured from oblique perspectives, thereby improving the model’s understanding of distant regions in oblique images.
-  • ARS-Adapter: The ARS-Adapter innovatively introduces a fine-tuning method tailored to foundational models in aerial remote sensing, reducing the number of parameters adjusted, improving fine-tuning efficiency, and significantly lowering the risk of overfitting.
-In order to evaluate the effectiveness of the model, we performed six downstream tasks using 13 aerial remote sensing datasets and 5 aerospace remote sensing datasets with different viewpoints, resolutions and modes.
-The test results show that RingMo-Aerial achieves SOTA performance on all aerial remote sensing datasets and reaches the leading level on aerospace remote sensing datasets, which proves its strong generalization ability for intelligent interpretation of aerial and aerospace remote sensing images.
-Compared with the single-task model, RingMo-Aerial has obvious advantages in handling multi-tasks in different scenarios.
-II Related Works
-II-A Aerial remote sensing
-Aerial remote sensing offers unique characteristics and has found widespread applications across various domains.
-Common tasks in aerial remote sensing include object detection, object tracking, semantic segmentation, and 3D reconstruction.
+This approach optimizes the performance of the ARS foundational model, making it more adaptable and effective across various ARS vision tasks.
+In summary, the contributions of this paper can be summarised as follows:
+  • Addressing ARS-specific Characteristics: We introduce RingMo-Aerial, the first foundational model for ARS, supporting multi-perspective, multi-resolution, and multi-modality tasks across various applications.
+  • FE-MSA Module: By incorporating a patch expansion layer and integrating the window attention mechanism, FE-MSA effectively enhances the detection capability of small, dense targets in ARS images.
+  • CLAF Module: CLAF uses affine transformations to capture ARS-specific visual features, particularly in oblique perspectives, improving the model’s recognition of distant objects.
+  • ARS-Adapter: The ARS-Adapter provides a fine-tuning approach tailored to ARS models, enhancing efficiency and reducing overfitting risks.
+To evaluate the effectiveness of the model, we performed six downstream tasks using 13 ARS datasets and 5 ApRS datasets with different viewpoints, resolutions, and modes.
+The test results show that RingMo-Aerial achieves SOTA performance on all ARS datasets and reaches the leading level on ApRS datasets, which proves its strong generalization ability for intelligent interpretation of aerial and ApRS images.
+Compared with the single-task model, RingMo-Aerial has apparent advantages in handling multi-tasks in different scenarios.
+2 Related Work
+2.1 Aerial remote sensing(ARS)
+ARS offers unique characteristics and applications across various domains, supporting tasks such as object detection, object tracking, semantic segmentation, and 3D-reconstruction.
 The most frequently utilized modalities are optical imaging [7], infrared imaging [8], and synthetic aperture radar (SAR) [9].
-Compared to satellite remote sensing, aerial remote sensing provides higher image resolution and more flexible spatiotemporal coverage.
-However, it also presents distinct characteristics that set it apart from satellite-based methods.
-One notable difference is that aerial imagery is typically captured at oblique angles with varying shooting distances.
-This results in significant variations in the size of similar targets and increases the likelihood of target occlusion.
-[10, 11, 12, 13] have specifically modeled occluded targets at the structural level, significantly improving the detection and segmentation performance of occluded instances.
-[14, 15, 16, 17] have focused on optimizing the oblique perspective through multi-scale transformations, enhancing the resolution of small targets.
-Another key distinction lies in the frequency content of aerial imagery, which contains both high- and low-frequency signals, unlike satellite imagery that predominantly captures low-frequency information.
-Effectively leveraging high-frequency signals can greatly aid in the interpretation of aerial images.
-[18, 19] introduced frequency-domain enhancement modules to balance high- and low-frequency signals, leading to improved performance across various aerial remote sensing tasks.
-II-B Visual foundation model
-Figure 4: Comparison of model frameworks between RingMo and RingMo-Aerial.
-Visual foundation models have gained attention for their ability to serve as general-purpose solutions in various computer vision tasks.
-Convolutional neural networks (CNNs) were among the earliest architectures used for visual foundation models, with well-known examples such as VGG [20] and ResNet [21].
-These models rely on convolutional layers for feature extraction and are typically pre-trained on classification tasks using the ImageNet [22] dataset.
-With the advent of Transformers in natural language processing, visual Transformer models like ViT [23] and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training.
+Compared to ApRS, ARS enables higher image resolution and flexible spatiotemporal coverage.
+However, it also presents distinct characteristics.
+Aerial imagery is typically captured at oblique angles with varying distances, leading to differences in target sizes and an increased likelihood of occlusion.
+For example, [10, 11, 12, 13] addressed target occlusion by modeling occluded objects at the structural level, significantly enhancing detection and segmentation performance.
+Studies such as [14, 15, 16, 17] have improved small target resolution by optimizing for oblique perspectives through multi-scale transformations.
+Another critical difference lies in the frequency content of ARS imagery.
+Unlike ApRS images, which primarily capture low-frequency information, aerial images contain a mix of high-frequency and low-frequency signals, enabling detailed image interpretation.
+Recent work [18, 19] introduced frequency-domain enhancement modules to balance these signals, thereby improving performance across various ARS tasks.
+2.2 Visual foundation model
+Visual foundation models have gained attention as general-purpose solutions for computer vision tasks.
+Early models, based on convolutional neural networks (CNNs) like VGG [20] and ResNet [21], used convolutional layers for feature extraction and were typically pre-trained on classification tasks using ImageNet [22].
+With the advent of Transformers in natural language processing, visual Transformer models like ViT [23], and Swin Transformer [24] have gained prominence, utilizing masked image modeling (MIM) or contrastive learning (CL) for self-supervised pre-training.
 In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies.
-Additionally, a number of novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed.
-However, due to technical immaturity, these newer models have yet to achieve widespread adoption.
-The key issue with these models lies in the pre-training phase, where they learn broad visual representations from large datasets.
-This chapter will also summarize three types of self-supervised visual training methods: CL, MIM, and their combination, with each method addressing both natural and RS.
-II-B1 Contrastive Learning
-Contrastive learning is an self-supervised learning approach that emphasizes learning the common features among similar instances while distinguishing differences between dissimilar instances.
-Major contrastive learning methods include the MoCo [5, 27, 28] series.
+Several novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed.
+However, these newer models have yet to achieve widespread adoption due to technical immaturity.
+A major challenge for these models lies in the pre-training phase, where they learn broad visual representations from large datasets.
+This section provides an overview of three self-supervised training methods—contrastive learning, masked image modeling, and their combination—each addressing both natural and RS images.
+2.2.1 Contrastive Learning
+Contrastive learning is a self-supervised learning approach that emphasizes learning the standard features among similar instances while distinguishing differences between dissimilar cases.
+Primary contrastive learning methods include the MoCo [5, 27, 28] series.
 MoCov2 [27], an optimization of MoCo [5] inspired by SimCLR [29], introduces enhancements such as the MLP projection head and increased data augmentation.
@@
 DINO [31], another self-supervised learning method, employs knowledge distillation techniques to train ViT for semantic segmentation and classification feature learning without labels, leveraging momentum encoders and multi-crop training for enhanced model performance.
-In the field of RS, contrastive learning also has many applications.
+In RS, contrastive learning also has many applications.
 GLCNet [32] presents a self-supervised contrastive learning method for semantic segmentation of high-resolution RS images.
-SeCo [33] introduces a self-supervised pretraining method that leverages temporal information in RS images to generate positive sample pairs, learning feature representations invariant and diverse to seasonal variations.
-II-B2 Masked Image Modeling
+SeCo [33] introduces a self-supervised pre-training method that leverages temporal information in RS images to generate positive sample pairs, learning feature representations invariant and diverse to seasonal variations.
+2.2.2 Masked Image Modeling
 MIM is an emerging self-supervised pre-training method.
@@
 UM-MAE [36] applies uniform masking to address complexities in ViTs, while CAE [37] adds a regressor for distinct feature encoding separate from reconstruction.
-In the field of RS, RingMo [1] pioneered the foundation model framework based on MIM.
+In RS, RingMo [1] pioneered the foundation model framework based on MIM.
 Unlike random masking techniques used in natural scenes, RingMo developed the PIMask strategy tailored to the complex backgrounds and small targets characteristic of RS data, focusing on preserving information about small targets.
 By integrating ViT and Swin Transformer, RingMo enhanced the accuracy of various downstream RS tasks, laying a foundation for developing large models in RS scenes.
-II-B3 The combination of Contrastive Learning &\&& Masked Image Modeling
+2.2.3 The combination of Contrastive Learning &\&& Masked Image Modeling
 Combining CL and MIM has shown significant advantages in current research, particularly in enhancing feature extraction performance.
@@
 It also enhances the interaction between different image regions, boosting feature recognition capabilities.
-Despite the success of models like ViT and RingMo in general and satellite remote sensing tasks, their performance in aerial remote sensing remains limited.
-Currently, there is no dedicated foundation model for aerial remote sensing.
-To address this gap, we propose RingMo-Aerial, a model designed specifically for aerial remote sensing tasks, and demonstrate its effectiveness in common applications.
-III Problem Description
-The unique imaging characteristics of ARS photography stem from the inherent properties of aerial perspective.
-Specifically, this perspective manifests as a size disparity within the same image due to the inclined angle of capture.
-Objects that are closer to the ARS (towards the bottom of the image) appear larger, while those further away (towards the top) appear smaller, as illustrated in Figure 5.
-To elaborate, as depicted in Figure 6, when a ARS captures two objects of equal actual size at inclined angles of $\theta_{1}$ and $\theta_{2}$, the resulting pixel dimensions and their ratio can be mathematically expressed.
-For illustration, consider two identical cars on a highway, both with a height of L.
-The vertical component of the cars in the camera’s view can be expressed as $L\times sin(\theta_{1})$ and $L\times sin(\theta_{2})$.
-Given the drone’s flight height is $H$, the focal length of the drone’s camera is $f$, and the horizontal distances between the two cars and the drone are $d_{1}$ and $d_{2}$ respectively, the slant distances between the cars and the drone are calculated as $D_{1}=\sqrt{H^{2}+d_{1}^{2}}$ and $D_{1}=\sqrt{H^{2}+d_{2}^{2}}$.
-According to the principle of perspective projection, the height of the objects in pixels in the image is calculated as follows:
-$h_{1}=\frac{f\cdot L\cdot\sin(\theta_{1})}{D_{1}}$ (1)
-$h_{2}=\frac{f\cdot L\cdot\sin(\theta_{2})}{D_{2}}$ (2)
-The ratio of the heights of the pixels occupied by the cars is:
-$\frac{h_{1}}{h_{2}}=\frac{D_{2}}{D_{1}}\times\frac{\sin(\theta_{1})}{\sin(\theta_{2})}=\frac{\sqrt{H^{2}+d_{2}^{2}}}{\sqrt{H^{2}+d_{1}^{2}}}\times\frac{\sin(\theta_{1})}{\sin(\theta_{2})}$ (3)
-Then the proportion of pixels occupied by car 1 and car 2 can be estimated as:
-$\frac{p_{1}}{p_{2}}=\frac{h_{1}^{2}}{h_{2}^{2}}$ (4)
+Despite the success of models like ViT and RingMo in general and satellite remote sensing tasks, their performance in ARS remains limited.
+Currently, there is no dedicated foundation model for ARS.
+To address this gap, we propose RingMo-Aerial, a model designed specifically for ARS tasks, and demonstrate its effectiveness in typical applications.
+3 Problem Description
 Figure 5: Similar targets with significant size differences in the same picture.
-In the Vision Transformer model, objects of different sizes occupy different numbers of tokens.The image is from VisDrone dataset.
-Figure 6: The difference in shooting angles causes the final pixel sizes of similar targets to be different.
-Considering that in the visual transformer model, the number of pixels occupied by an object is related to the number of tokens it occupies, as 5.
-Therefore, it is obviously useful to adopt a method to balance the number of tokens occupied by similar objects of different sizes in an image.
+In the Vision Transformer model, objects of different sizes occupy different numbers of tokens.
+The image is from the VisDrone dataset.
+The unique imaging properties of ARS photography stem from the inherent properties of aerial perspective.
+Specifically, this perspective manifests itself in size differences within the same image due to the tilted angle of the shot.
+As shown in Fig.5, similar objects at different distances from the ARS are rendered at different sizes in the image.
+Figure 6: The same target will be represented differently in the image if its distance from the lens and imaging angle are different.
+As the horizontal distance increases, the number of pixels occupied by the height content increases and then decreases, and the height content occupies the most pixels at the horizontal distance $l=\sqrt{H(H-h)}$.
+As shown in Fig.6, when the camera is positioned at a height $H$ and the target object has a height $h$, the angular size $\alpha$ of the target’s height in the camera’s field of view varies as a function of the horizontal distance $l$ between the target and the camera.
+At different horizontal distances, the imaging size corresponding to the same target height differs.
+Specifically, when the target is directly beneath the camera, its height is not captured in the image.
+As the horizontal distance increases, the number of pixels occupied by the target’s height initially increases and then decreases.
+By considering the distances from the target’s lowest and highest points to the camera, given by $x=\sqrt{H^{2}+l^{2}}$ and $y=\sqrt{(H-h)^{2}+l^{2}}$ respectively, we can derive the relationship between the viewing angle $\alpha$ and the horizontal distance $l$ between the target and the camera:
+Using the cosine rule:
+$x^{2}+y^{2}-2xy\cos\alpha=h^{2}$ (1)
+Substituting the values of $x$ and $y$, The final cosine formula becomes:
+$\cos\alpha=\frac{H^{2}-Hh+l^{2}}{\sqrt{H^{2}+l^{2}}\cdot\sqrt{(H-h)^{2}+l^{2}}}$ (2)
+By setting $a=H$ and $b=H-h$, we have the target function is:
+$f(l)=\cos\alpha=\frac{ab+l^{2}}{\sqrt{a^{2}+l^{2}}\cdot\sqrt{b^{2}+l^{2}}}$ (3)
+Taking the derivative of $f(l)$:
+$f^{\prime}(l)=\frac{2l(a^{2}+l^{2})(b^{2}+l^{2})-l(ab+l^{2})(a^{2}+b^{2}+2l^{2})}{(a^{2}+l^{2})^{3/2}(b^{2}+l^{2})^{3/2}}$ (4)
+Solving $f^{\prime}(l)=0$ results in the equation:
+$2(a^{2}+l^{2})(b^{2}+l^{2})=(ab+l^{2})(a^{2}+b^{2}+2l^{2})$ (5)
+Simplifying further:
+$2a^{2}b^{2}+a^{2}l^{2}+b^{2}l^{2}=a^{3}b+ab^{3}+2abl^{2}$ (6)
+Combining terms:
+$l^{2}(a-b)^{2}=ab(a-b)^{2}$ (7)
+Since $a=H$ and $b=H-h$, and $a>b$, it follows that $a-b=h>0$.
+When $f^{\prime}(l)=0$, we obtain:
+$l^{2}=ab$ (8)
+Therefore:
+$l=\sqrt{ab}=\sqrt{H(H-h)}$ (9)
+This completes the derivation of the maximum value $\alpha$ and the corresponding minimum value $\cos\alpha$.
+From Eq.
+(2), when $l=0$, $\alpha$ is equal to 0.
+From Eqs.(4) and (9), $\alpha$ increases with $l$ is less than $\sqrt{H(H-h)}$, $\alpha$ is maximum when $l=\sqrt{H(H-h)}$, and decreases with $l$ when $l$ is greater than $\sqrt{H(H-h)}$, and decreases with $l$ in the case of $l\rightarrow\infty$ when $\alpha$ is again equal to 0.
+Assuming that the optical axis of the camera is always at the center of the target viewpoint $\alpha$, the pixel change of the target height information in the picture is the same as the change of the corresponding viewpoint $\alpha$ of the target.
+This indicates that if the same target is at different distances from the lens and imaging angles, it will have different representations in the image.
+Figure 7: RingMo-Aerial’s two-stage pre-training principles Capturing Pixel-level Semantic Associations and Learning Aerial Visual Physical Characteristics, and the architecture of the FE-MSA module.
+In the visual transformer model, the number of pixels occupied by an object is related to the number of tokens it occupies, as Fig.5.
+Therefore, it is helpful to adopt a method to balance the number of tokens occupied by similar objects of different sizes in an image.
 In addition to the oblique angles, ARS images typically contain many small objects, which can be very dense.
@@
 Consequently, a practical pre-training framework and method, specifically designed to address the distinctive visual characteristics of ARS images, is essential to achieve satisfactory performance in a range of downstream tasks.
-Figure 7: The main structure of RingMo-Aerial.
-IV Methodology
-IV-A Method Overview
+4 Methodology
+4.1 Method Overview
 The proposed RingMo-Aerial model serves as a foundational framework with broad applicability.
-Its implementation involves three main stages: model design, pretraining, and fine-tuning for downstream tasks.
-In Section IV-B, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase.
-To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown as Figure 7.
-In Section IV-C, we combine the commonly used masked image modeling (MIM) pretraining with an additional contrastive learning (CL) method based on affine transformations during the pretraining stage.
-This hybrid pretraining approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown as Figure 7.
-Furthermore, in Section IV-D, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase.
-The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks, shown as Figure 10.
-IV-B Model Architecture
-RingMo-Aerial’s design is based on Swin Transformer.
-The architecture of Swin Transformer consists of four stages, each comprising either a linear encoder or Patch Merging, along with an even number of Swin Transformer Blocks.
+Its implementation involves three main stages: model design, pre-training, and fine-tuning for downstream tasks.
+In Section 4.2, we address the challenge of detecting small objects in ARS imagery caused by oblique viewing angles during the model design phase.
+To enhance small object detection, we introduce the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) layer by incorporating a patch expansion layer, shown in Fig.7.
+In Section 4.3, we combine the commonly used masked image modeling (MIM) pre-training with additional contrastive learning (CL) methods based on affine transformations during the pre-training stage.
+This hybrid pre-training approach significantly improves both training speed and the performance of the converged model on downstream tasks, shown in Fig.7.
+Furthermore, in Section 4.4, we introduce a parallel Adapter fine-tuning method during the fine-tuning phase.
+The proposed ARS-Adapter module includes downsampling, channel-wise convolution, and upsampling components, enhancing its adaptability and learning capacity for specific tasks.
+4.2 Model Architecture
+RingMo-Aerial’s design is based on the RingMo.
+The architecture of the Swin Transformer consists of four stages, each comprising either a linear encoder or Patch Merging, along with an even number of Swin Transformer Blocks.
 Each block includes Window Multi-Head Self-Attention (W-MSA) and Shifted Window Multi-Head Self-Attention (SW-MSA).
@@
 Relying solely on semantic information may not be sufficient to achieve optimal performance in downstream tasks.
-To address the aforementioned challenges, we propose the overall architecture of RingMo-Aerial.
-As shown in Figure 7(a), the overall architecture of RingMo-Aerial is based on the Swin Transformer.
-Considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
+To address the challenges above, we propose the overall architecture of RingMo-Aerial.
+As shown in Fig.7 Pre-training Stage 1, considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
 Figure 8: The effect of FE-MSA.
@@
 At the same time, FE-MSA enhances the model’s receptive field for small objects (purple box).
+Figure 9: Positive samples for contrastive learning are obtained after the image’s affine transformation, cropping, and resizing.
 Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module.
-To better understand the FE-MSA mechanism, it is essential to first recognize the limitations of the existing window attention mechanisms.
+To better understand the FE-MSA mechanism, it is essential first to recognize the limitations of the existing window attention mechanisms.
 These mechanisms often struggle to balance the semantic information in feature vectors of objects of different sizes.
-For instance, as illustrated in Figure 8, larger objects may occupy $4\times 4$ tokens, while smaller distant objects may occupy only $3\times 3$ tokens.
+For instance, as illustrated in Fig.8, larger objects may occupy $4\times 4$ tokens, while smaller distant objects may occupy only $3\times 3$ tokens.
 The FE-MSA mechanism introduces frequency domain processing to the existing window attention, indirectly expanding the receptive field for smaller objects.
-As shown in Figure 7(b), the RingMo-Aerial Block structure is illustrated, including two window attention mechanisms optimized by the FE-MSA mechanism.
+Fig.7(b) illustrates the RingMo-Aerial Block structure, including two window attention mechanisms optimized by the FE-MSA mechanism.
 Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block.
@@
 Then, W-MSA or SW-MSA is used to compute the attention map.
-After the FE-MSA, the edge and texture features of small objects are further enhanced, improving the model’s ability to perceive both local and global features.
+After the FE-MSA, small objects’ edge and texture features are further enhanced, improving the model’s ability to perceive local and global features.
 The computation process for each token patch through the RingMo-Aerial block is as follows:
@@
 […]
-The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, enhancing the model’s performance in complex aerial scenes.
-IV-C Pre-train
-Figure 9: Positive samples for contrastive learning are obtained after affine transformation, cropping and resizing of the image.
-For pretrain, we construct a large-scale pre-training dataset.
+The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, improving the model’s performance in complex aerial scenes.
+Fig.10 presents a comparative analysis of the feature extraction capabilities of the RingMo-Aerial and RingMo models in processing multiview ARS images.
+The heatmaps generated by RingMo-Aerial show more focused attention on critical areas, such as vehicles and buildings, as indicated by the bright regions.
+In contrast, the RingMo in the lower half has a more diffuse focus, with less concentration on specific features.
+This comparison demonstrates the superior ability of RingMo-Aerial to handle complex aerial scenes, enabling better feature extraction and object recognition in multiview ARS images.
+These results visually confirm the improvements introduced by RingMo-Aerial, particularly in addressing challenges such as varying scales and perspectives in aerial imagery.
+4.3 Pre-training
+Figure 10: Comparative analysis of feature heatmaps extracted from multi-view aerial imagery using RingMo-Aerial and RingMo.
+The feature heatmaps extracted by the RingMo-Aerial model are located in the upper half, while those from the RingMo model are in the lower half of the figure.
+In these images, the brighter the region, the more the model is focusing on that particular feature, indicating that these areas have higher feature importance for the model.
+For pre-training, we construct a large-scale pre-training dataset.
 Our comprehensive pre-training dataset covers a wide range of aerial scenes, including high-altitude, low-altitude, and near-ground perspectives.
 It spans diverse geographical regions such as urban, rural, and wilderness areas.
-The dataset includes various targets, such as roads, vehicles, and pedestrians, and captures different seasons and time periods.
+The dataset includes various targets, such as roads, vehicles, and pedestrians, and captures different seasons and periods.
 Notably, it exhibits multi-source, multi-temporal, and multi-instance characteristics.
-Prior to training, we crop the entire dataset into 640x640 or 448x448 resolution slices, all captured using optical three-channel imaging.
-In total, the sliced dataset comprises approximately 5.4 million images, as Table I.
-TABLE I: The Number of Post-Cropping Images Per Dataset.
-Datasets Images
-AU-AIR [42] 328230
-DAC-SDC [43] 73740
-DarkTrack2021 [44] 223606
-DroneCrowd [45] 42960
-Okutama-Action [46] 530874
-Seadronessee [47] 913986
-StanfordDroneDataset [48] 580223
-UAVDark135 [49] 124980
-UZH-FPV [50] 184458
-VTUAV [51] 1576800
-WebUAV3M [52] 376903
-DTB70 [53] 47331
-ERA [54] 39720
-ITCVD [55] 27680
-UAV123 [56] 11161
-UAVHuman [57] 22549
-UAV-ROD [58] 42088
-UCLA Aerial Event dataset [59] 27370
-Internet collection 228316
-Total 5402975
-During the pre-training phase, we adopt a combination of two techniques: Masked Image Modeling (MIM) and Contrastive Learning (CL).
+Before training, we crop the entire dataset into $640\times 640$ or $448\times 448$ resolution slices, all captured using optical three-channel imaging.
+The sliced dataset comprises approximately 5.4 million images (Fig.11).
+Figure 11: The figure shows the number of slices corresponding to the slicing process performed on multiple collected open source airborne remote sensing datasets, which were used as the dataset for pre-training of the base model.
+During the pre-training phase, we combine Masked Image Modeling (MIM) and Contrastive Learning (CL) techniques.
 For MIM training, we use the standard SimMIM pre-training framework.
@@
 On this basis, the perspective transformation matrix can be calculated.
-The perspective transformation matrix is defined as a $3\texttimes 3$ matrix in the form of:
+The perspective transformation matrix is defined as a $3\times 3$ matrix in the form of:
 $M=\begin{bmatrix}m_{11}&m_{12}&m_{13}\\ m_{21}&m_{22}&m_{23}\\ m_{31}&m_{32}&1\end{bmatrix}$ (14)
-The solution process of this transformation matrix is:
+The solution process of this transformation matrix is as follows:
 $\begin{cases}u_{i}=\frac{m_{11}x_{i}+m_{12}y_{i}+m_{13}}{m_{31}x_{i}+m_{32}y_{i}+1}\\ v_{i}=\frac{m_{21}x_{i}+m_{22}y_{i}+m_{23}}{m_{31}x_{i}+m_{32}y_{i}+1}\end{cases}~{}\text{for}~{}i=0,1,2,3$ (15)
@@
 $\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=M\cdot\begin{bmatrix}x\\ y\\ 1\end{bmatrix}$ (16)
-Through the aforementioned affine transformation process, we obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$.
+We obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$ through the aforementioned affine transformation process.
 The source image $I_{i}$ and its corresponding transformed image $I_{i}^{\prime}$ form a positive sample pair, while the remaining sample pairs $I_{i}$ and $I_{j}^{\prime}$ (where $i\neq j$) form negative sample pairs.
-By calculating the contrastive loss, we guide the model to learn the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.
-IV-D Downstream Fine-tuning
-Figure 10: The proposed ARS-Adapter.
-Figure 11: Qualitative results for the UAVID dataset when using different methods.
-In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks.
-Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model.
-The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, as 10.
-The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking.
-In a fashion akin to AdaptFormer [60], the introduced ARS-Adapter module operates in parallel with the model.
-The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer.
-Therefore, the forward propagation is:
-[…]
-[…]
-Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling.
-The Adapter in parallel with FFN introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
-[…]
-V Experiments
-TABLE II: Performance Comparison on AID and RESISC-45 Datasets
-Model AID (OA) RESISC-45 (OA)
-(TR=20%/50%) (TR=10%/20%)
-GASSL [61] 93.55/95.92 90.86/93.06
-SeCo [33] 93.47/95.99 89.64/92.91
-SatMAE [62] 95.02/96.94 91.72/94.10
-RingMo [1] 96.90/98.34 94.25/95.67
-RVSA [63] 97.03/98.50 93.93/95.69
-DINO-MC [64] 95.16/97.09 -
-TOV [65] 95.16/97.09 90.97/93.79
-SSL4EO [66] 91.06/94.74 87.60/91.27
-CMID [67] 96.11/97.79 91.84/94.07
-CACo [68] 90.88/95.05 88.28/91.94
-CROMA [69] 96.44/97.58 92.63/95.04
-SatLas [70] 94.96/97.38 92.16/94.70
-GFM [71] 95.47/97.09 -
-Scale-MAE [72] 96.44/97.58 92.63/95.04
-SkySense [2] 97.68/98.60 94.85/96.32
-RingMo-Aerial 95.81/96.46 92.28/95.65
-TABLE III: Compare our model with other methods on LEVIR-CD and CDD dataset
-Methods LEVIR-CD CDD
-F1 IoU F1 IoU
-FC-Siam-Di [73] 86.31 83.31 70.61 54.57
-FC-Siam-Conc [73] 83.69 76.77 75.11 60.14
-IFNet [74] 88.13 82.93 90.30 71.91
-STANet [75] 87.30 77.40 84.12 72.22
-BIT [76] 89.31 89.37 88.90 80.01
-SNUNet [77] 88.16 87.17 83.89 72.11
-ChangeFormer [78] 90.40 88.80 89.83 81.53
-RingMo [1] 89.53 82.38 90.90 84.09
-RingMo-Aerial (Ours) 92.71 87.10 94.62 90.09
-TABLE IV: Comp. on UAVID
+By calculating the contrastive loss, we guide the model in learning the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.
+4.4 Downstream Fine-tuning
+Figure 12: ARS-Adapter is an enhancement module designed to improve transformer models by adding efficient, low-cost adaptation layers.
+It integrates down-sampling, activation (e.g., GELU), and up-sampling, optionally including convolution, to enrich feature representations while reducing computational overhead.
+TABLE I: Semantic Segmentation: Comp. on UAVID
 Methods Building Tree Clutter Road Low vegetation Static car Moving car Human mAcc mIoU
@@
 RingMo-Aerial(M)* 93.2 85.7 89.8 79.0 79.0 66.3 87.8 82.0 76.8 62.0 72.5 65.5 88.9 76.6 41.3 32.7 78.7 68.7
-    Swin-Base (U) means using the Swin-Base backbone plus UperNet segmentation algorithm.
+    Swin-Base (U) uses the Swin-Base backbone plus the UperNet segmentation algorithm.
 RingMo-Aerial (U), (A), (M) represent the RingMo aerial backbone with UperNet segmentation algorithm, ARS-Adapter and UperNet algorithm, and Mask2Former algorithm.
 Same below.
-TABLE V: Comp. on UDD6
+In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks.
+Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model.
+The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, shown as Fig.12.
+The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking.
+Like AdaptFormer [55], the introduced ARS-Adapter module operates in parallel with the model.
+The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer.
+Therefore, the forward propagation is:
+[…]
+[…]
+Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling.
+In parallel with FFN, the adapter introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
+[…]
+5 Experiments
+TABLE II: Semantic Segmentation: Comp. on UDD6
 Methods Other Facade Road Vegetation Vehicle Roof mAcc mIoU
@@
 RingMo-Aerial(M) 90.7 80.9 90.2 77.4 88.4 77.1 90.5 80.8 82.1 69.4 96.3 87.8 89.7 78.9
-TABLE VI: Comp. on Floodnet
+TABLE III: Semantic Segmentation: Comp. on Floodnet
 Methods Building Road Water Tree Vehicle Pool Grass OA mIoU
 Flooded Non-Flooded Flooded Non-Flooded
+IoU IoU IoU IoU IoU IoU IoU IoU IoU
 SegNet [56] 44.72 69.85 39.70 76.44 71.17 76.88 32.74 35.87 85.84 87.98 59.25
-FCN8s 48.93 73.84 43.11 80.22 71.71 77.33 38.19 45.25 86.85 88.78 62.83
-FCN16s 44.97 67.54 34.14 72.80 65.39 76.07 34.28 32.27 85.66 87.23 57.01
-PSPNet 45.24 70.30 43.05 76.06 74.32 74.59 24.41 40.74 85.88 88.01 59.40
-DeepLabV3 50.32 78.02 48.46 82.06 76.94 79.90 47.14 48.45 88.10 90.13 66.60
+FCN8s [57] 48.93 73.84 43.11 80.22 71.71 77.33 38.19 45.25 86.85 88.78 62.83
+FCN16s [57] 44.97 67.54 34.14 72.80 65.39 76.07 34.28 32.27 85.66 87.23 57.01
+PSPNet [58] 45.24 70.30 43.05 76.06 74.32 74.59 24.41 40.74 85.88 88.01 59.40
+DeepLabV3 [59] 50.32 78.02 48.46 82.06 76.94 79.90 47.14 48.45 88.10 90.13 66.60
 DeepLabV3+ [42] 47.62 74.46 43.12 79.07 68.35 76.36 35.47 36.69 85.60 87.79 60.75
-UNet 49.95 77.58 43.90 81.90 73.46 78.88 43.30 49.30 87.54 89.56 65.09
-HRNet 50.04 79.23 48.62 81.59 75.05 80.16 49.80 50.89 88.38 90.17 67.08
+UNet [60] 49.95 77.58 43.90 81.90 73.46 78.88 43.30 49.30 87.54 89.56 65.09
+HRNet [61] 50.04 79.23 48.62 81.59 75.05 80.16 49.80 50.89 88.38 90.17 67.08
 TransUNet [62] 48.97 73.67 43.11 79.09 72.86 77.75 26.27 35.02 86.47 88.82 60.36
@@
 RingMo-Aerial(M) 76.17 77.90 48.57 82.62 62.75 81.70 63.16 65.79 87.16 90.43 71.76
-V-A Pre-train
-In the pretraining phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized.
-Prior to pretraining, we initialize the model parameters using Kaiming initialization.
-Subsequently, we perform 100 epochs of masked image modeling (MIM) pretraining.
-During MIM pretraining, the input image size is set to 224, and random cropping and scaling are applied to the input images.
+5.1 Pre-training Implementation
+In the pre-training phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized.
+Before pre-training, we initialize the model parameters using Kaiming initialization.
+Subsequently, we perform 100 epochs of masked image modeling (MIM) pre-training.
+During MIM pre-training, the input image size is 224, and random cropping and scaling are applied to the input images.
 The mask size is set to 32 pixels.
-The batch size is set to 2048, with a base learning rate of 2e-4, cosine annealing learning rate scheduling, and a warm-up period of 2 epochs.
-Following MIM pretraining, we continue with contrastive learning (CL) pretraining, jointly training for 36 epochs.
+The batch size is 2048, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 2 epochs.
+Following MIM pre-training, we continue with contrastive learning (CL) pre-training, jointly training for 36 epochs.
 The input image size remains 224, and similar random cropping and scaling are applied.
-The batch size is adjusted to 1024, with a base learning rate of 2e-4, cosine annealing learning rate scheduling, and a warm-up period of 1 epoch.
-V-B Scene Classification
-V-B1 Datasets and Evaluation Metrics
-For semantic segmentation, the AID [94] and RESISC-45 [95] datasets are employed.
-  • AID.
-The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification.
-It contains images collected from Google Earth and covers 30 different types of scenes.
-  • NWPU-RESISC45.
-The NWPU-RESISC45 dataset is a publicly available REmote Sensing Image Scene Classification (RESISC) benchmark.
-The dataset contains 31,500 images covering 45 scene classes, with 700 images in each class.
-In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively.
-In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively.
-The experiments use overall accuracy (OA) as a performance metric, which is the most commonly used performance metric in scene classification.
-V-B2 Experiment Settings
-The experiment of the scene classification is based on mmpretrain , Cascade RCNN [96] is employed as the classification framework.
-The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs).
-The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05.
-In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
-V-B3 Comparison With the SOTA
-The comparison results are shown in Table II. RingMo-Aerial combined with Cascade RCNN can achieve classification results far exceeding other methods.
-V-C Change Detection
-V-C1 Datasets and Evaluation Metrics
-For semantic segmentation, the LEVIR [75] and CDD [98] datasets are employed.
-  • LEVIR.
-The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset.
-It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5 meter resolution.
-Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels.
-Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases, respectively.
-  • CDD.
-The CDD dataset is a comprehensive collection of bitemporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels.
-These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches.
-This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs.
-To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
-V-C2 Experiment Settings
-Our model adopts BIT with Swin Transformer as the backbone.
-The experiments use overall accuracy (OA) and F1 and IoU scores with regard to the change detection as the primary evaluation metrics.
-The model is trained with AdamW optimizer with a weight decay 0.01 and beta values equal to (0.9, 0.999).
-The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size of 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
-V-C3 Comparison With the SOTA
-Figure 12: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset.
-The comparison results are shown in Table III.
-RingMo-Aerial combined with BIT can achieve change detection results far exceeding other methods.
-In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection.
-When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated in our architecture.
-The comparison of visual results12 on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks.
-The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images.
-In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
-RingMo-Aerial’s commendable performance not only makes it a key player in the field, but also highlights the importance of the impact of foundation model innovation.
-These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
-V-D 3D-Reconstruction
-V-D1 Datasets and Evaluation Metrics
-For 3D-Reconstruction, the and WHU [99] datasets are employed.
-  • WHU.
-The WHU dataset is a synthetic aerial dataset designed for large-scale multi-view stereo reconstruction tasks.
-It was generated from a highly accurate 3D surface model and includes a complete set of aerial images as well as cropped sub-image sets specifically for deep learning applications.
-The images in the dataset cover an area of approximately 6.7 × 2.2 square kilometers in Meitan County, Guizhou Province, China, with a ground resolution of 0.1 meters.
-The WHU dataset provides crucial data support for the study of large-scale Earth surface and urban reconstruction and can be utilized for training and testing multi-view stereo matching algorithms.
-  • LuoJia-MVS.
-The LuoJia-MVS dataset comprises 7,972 five-view images, each with a spatial resolution of 10 cm, accompanied by pixel-wise depth information and precise camera parameters.
-These images were generated from a highly accurate digital surface model (DSM) constructed using thousands of stereo aerial images.
-During the experiment, we used the WHU dataset[99] and the LuoJia-MVS dataset[100] to verify the effectiveness of our proposed method.
-WHU dataset[99] is the first large-scale aerial multi-view stereo dataset for remote sensing.
-It was collected from Meitan, Guizhou, China, and was obtained by taking large-scale photos with a tilted five-eye camera on a drone platform.
-The area of the entire dataset is approximately $6.7\times 2.2km^{2}$, and the image resolution is about 0.1m.
-In addition, the tilt angle of each camera to the central camera is 40°.
-The WHU dataset gathers high-rise buildings, mountains, forests, and factory scenes, providing a good benchmark for remote sensing multi-view stereo.
-LuoJia-MVS[100] dataset is a large-scale multi-view stereo dataset collected from a variety of landform types.
-It was collected in a part of Baiyun, Guiyang, Guizhou, China.
-The image size of the entire dataset is $768\times 384$, and the image resolution is about 0.1m.
-The data consists of five views and pixel-by-pixel depth labels.
-Land types cover cultivated land, forest land, industrial areas, residential areas, etc.
-V-D2 Experiment Settings
-Our work is coded based on the PyTorch framework.
-The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1.
-During the training process, our method uses the Adam optimizer[101] for parameter optimization, and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001.
-In addition, our method is consistent with the existing multi-stage networks[102].
-First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively.
-Then, deep optimization is performed through three stages of deep hypotheses, and the number of hypotheses is 48, 32, and 8, respectively.
-During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
-V-D3 Comparison With the SOTA
-TABLE VII: Performance comparisons on the LuoJia-MVS and WHU datasets.
-Number of Views Model LuoJia-MVS WHU
-MAE $<$0.6m $<$3-interval MAE $<$0.6m $<$3-interval
-Three views * PatchmatchNet[103] 0.252 92.7 87.2 0.173 96.5 94.8
-Fast-MVSNet[104] 0.194 95.7 92.0 0.184 95.5 94.1
-MVSNet[105] 0.172 96.1 92.4 0.190 95.0 94.3
-R-MVSNet[106] 0.177 96.0 93.5 0.183 95.3 93.5
-RED-Net[99] 0.109 98.2 96.9 0.112 98.1 97.9
-Cas-MVSNet[102] 0.103 98.4 97.1 0.111 97.7 97.6
-RingMo-Aerial 0.095 98.7 97.8 0.121 97.8 97.5
-Five views * PatchmatchNet[103] 0.283 90.4 84.1 0.160 96.9 95.0
-Fast-MVSNet[104] 0.357 84.6 74.9 0.157 96.1 95.6
-MVSNet[105] 0.270 91.2 81.8 0.160 95.8 95.5
-R-MVSNet[106] 0.259 92.3 86.7 0.173 95.4 93.8
-RED-Net[99] 0.156 94.9 90.5 0.104 98.1 97.9
-Cas-MVSNet[102] 0.141 97.9 95.4 0.095 97.8 97.8
-RingMo-Aerial 0.122 98.1 96.5 0.091 98.0 97.9
-    ‘Three Views’ and ‘Five Views’ represent the number of views for the input of the network is three and five, respectively.
-Table VII provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets.
-Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error.
-On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet, and improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs.
-0.141), 0.2% (98.1% vs.
-97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a large improvement.
-In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%.
-It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
-V-E Semantic Segmentation
-V-E1 Datasets and Evaluation Metrics
-For semantic segmentation, the UAVID [107], UDD6 [108] and FloodNet[109] datasets are employed.
-  • UAVID.
-The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid and 15 test), and each sequence contains 10 pictures.
-In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
-A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
-  • UDD6.
-The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images.
-All of the images are 4K resolution.
-In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
-The training set is used for training and the validation set for testing.
-  • FloodNet.
-The FloodNet dataset contains 500 high-resolution remote sensing images with a resolution of 4K.
-The dataset is divided into training, validation, and test sets, with the training set containing 300 images, the validation set containing 100 images, and the test set containing 100 images.
-Images are annotated as flood and non-flood areas, supporting multi-class flood impact assessment.
-All images are cropped into 1024×1024 pixel slices with a 256-pixel overlap in the experiment, facilitating the training and evaluation of machine learning models.
-In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc) and average accuracy (aAcc).
-V-E2 Experiment Settings
-The experiment of the semantic segmentation is based on mmsegmentation [110], UperNet [111] and Mask2Former [112] are employed as the segmentation framework.
-The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used).
-The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
-In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
-V-E3 Comparison With the SOTA
-The comparison results on the UAVID dataset are shown in Table IV, results on the UDD6 dataset are shown in Table V, results on the FloodNet dataset are shown in Table VI.
-RingMo-Aerial combined with Mask2Former can achieve segmentation results far exceeding other methods.
-Experiments on UperNet (RingMo-Aerial vs.
-Swin-Base) also show that RingMo-Aerial is a more suitable basic model for ARS segmentation tasks.
+The batch size is adjusted to 1024, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 1 epoch.
+5.2 Performance on Single/Multi-modal Downstream Tasks
+We validated the performance of RingMo-Aerial across six downstream tasks, covering both single-modal and multi-modal tasks, and involving a variety of data modalities and tasks commonly encountered in remote sensing scenarios.
+Detailed descriptions of the datasets, experimental setups, and additional visualizations for each task are provided in the Appendix.
+5.2.1 Semantic Segmantation
+Comparative results on the UAVID, UDD6, and FloodNet datasets are presented in Tabs.
+I, II, and III, respectively.
+The combination of RingMo-Aerial with Mask2Former achieves segmentation outcomes that significantly outperform those of other methods.
 Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks.
-As illustrated in the Figure11, RingMo-Aerial consistently produces superior segmentation maps.
+RingMo-Aerial consistently produces superior segmentation maps.
 Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures.
@@
 This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
-V-F Object Detection
-TABLE VIII: Comp. on VisDrone-DET
+5.2.2 Object Detection
+TABLE IV: Object Detection: Comp. on VisDrone-DET
 Methods Backbone mAP mAP50 mAP75
@@
     RingMo-Aerial (C), (A), (D) represent the RingMo aerial backbone with Cascade RCNN detection algorithm, ARS-Adapter and Cascade RCNN algorithm, and DDQ DETR algorithm. Same below.
-TABLE IX: Comp. on UAVDT
+TABLE V: Object Detection: Comp. on UAVDT
 Methods Backbone mAP mAP50 mAP75
@@
 RingMo-Aerial (D) RingMo-Aerial 25.2 39.9 27.1
-TABLE X: Comp. on ShipDataset
+TABLE VI: Object Detection: Comp. on ShipDataset
 Methods Backbone mAP mAP50 mAPs
@@
 RingMo-Aerial(C) RingMo-Aerial 83.7 99.1 72.3
-TABLE XI: Comp. on DroneVehicle
+TABLE VII: Object Detection: Comp. on DroneVehicle
 Methods Modality Backbone mAP50
@@
 RingMo-Aerial(C) RGB+IR RingMo-Aerial 71.6
-Figure 13: Qualitative results for the VisDrone-DET dataset when using different methods.
-TABLE XII: Comp. on SARData
+TABLE VIII: Object Detection: Comp. on AerialSARData
 Methods Backbone mAP mAP50
 RetinaNet[65] RetinaNet 22.9 64.3
-YOLOX-S[134] CSPDarkNet-53 34.5 85.0
 Cascade R-CNN[67] ResNet101 34.5 81.3
 RingMo-Aerial(A) * RingMo-Aerial 30.9 77.6
-RingMo-Aerial(Y) * RingMo-Aerial 34.0 81.5
-RingMo-Aerial(C) * RingMo-Aerial 34.9 82.4
-    RingMo-Aerial (A), (Y), (C) represent the RingMo aerial backbone with ARS-Adapter and YOLOX-S algorithm, YOLOX-S algorithm, and Cascade RCNN algorithm.
-V-F1 Datasets and Evaluation Metrics
-To conduct a comprehensive comparison of object detection methods, we selected four unimodal benchmark datasets, including the VisDrone-DET[145], UAVDT[146] and ShipDataset[147], as well as one multimodal benchmark dataset, DroneVehicle[141] and our collection of SAR airborne remote sensing data, to demonstrate the performance of the representative methods discussed in this paper.
-  • VisDrone-DET.
-The Visdrone-DET dataset is the detection task part of the VisDrone dataset.
-The resolution of the images is about 2k.
-The RingMo-Aerial experiment crops the original image to 640×640 and uses multi-scale cropping and resizing for enhancement.
-The training set and validation set are used for training, and the labeled test-dev set is used for testing.
-  • UAVDT.
-The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with the average resolution of 1080 × 540 pixels.
-The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
-  • ShipDataset.
-The ShipDataset, captured at a resolution of $3840\times 2160$, encompasses five distinct scenarios in Shanghai, China, characterized by varying shooting angles and lighting conditions.
-This dataset has $96.97\%$ small ships.
-The images within this dataset predominantly feature congested ship traffic.
-The RingMo-Aerial experiment uses multi-scale resizing for enhancement.
-  • DroneVehicle.
-The DroneVehicle dataset is a large-scale dataset specifically designed for drone-based RGB-Infrared cross-modality vehicle detection tasks.
-It includes 28,439 pairs of RGB-Infrared images, totaling 56,878 images, covering various urban scenarios such as roads, residential areas, and parking lots under different lighting conditions, including day, night, and dark night.
-The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
-  • SARData.
-To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside.
-This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection.
-The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels.
-The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
-In our object detection experiments, the evaluation metrics employed are mean average precision (mAP), mAP at an IoU threshold of 50 (mAP50), and mAP at an IoU threshold of 75 (mAP75).
-For datasets containing smaller objects, we also utilize mAPs, which represents the mean average precision for detecting small objects (with pixels smaller than 32×32).
-V-F2 Experiment Settings
-The experiment of the object detection is based on mmdetection [148], Cascade RCNN [96], YOLOX-S[134] and DDQ DETR [127] are employed as the detection framework.
-The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used).
-That for YOLOX-S spanned 36 epochs with a batch size of 4 total (4 per GPU and 1 NVIDIA A40 GPU are used).
-That for DDQ DETR spanned 36 epochs with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used).
-The AdamW [97] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
-In addition, the learning rate is decreased by a factor of 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
-TABLE XIII: Object Tracking: Comp. on Visdrone-MOT
+RingMo-Aerial(C) * RingMo-Aerial 34.9 82.6
+    RingMo-Aerial (A), (C) represent the RingMo-Aerial backbone with ARS-Adapter and Cascade RCNN algorithm.
+An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance.
+On unimodal datasets such as VisDrone-DET (as shown in Tab.IV) and UAVDT (as shown in Tab.V), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods.
+The experimental results further demonstrate that RingMo-Aerial, compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
+What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Tab.VI), and PP-PicoDet-L follows.
+In contrast, YOLOX-L and YOLOX-Tiny show the worst results.
+Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
+For the multimodal dataset, DroneVehicle (as shown in Tab.VII), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods.
+Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial.
+This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications.
+Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
+In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Tab.VIII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance.
+Mainly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and a mAP50 of 84.4%, highlighting its superiority in SAR modality.
+These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery.
+Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments.
+These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
+Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields.
+For optical images, as shown in the appendix, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets.
+In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups.
+Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms.
+RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images.
+As shown in the appendix, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence.
+The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images.
+RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for ARS.
+5.2.3 Scene Classification
+TABLE IX: Scene Classification: Comp. on AID and RESISC-45
+Model AID (OA) RESISC-45 (OA)
+(TR=20%/50%) (TR=10%/20%)
+GASSL [96] 93.55/95.92 90.86/93.06
+SeCo [33] 93.47/95.99 89.64/92.91
+SatMAE [97] 95.02/96.94 91.72/94.10
+RingMo [1] 96.90/98.34 94.25/95.67
+RVSA [98] 97.03/98.50 93.93/95.69
+DINO-MC [99] 95.16/97.09 -
+TOV [100] 95.16/97.09 90.97/93.79
+SSL4EO [101] 91.06/94.74 87.60/91.27
+CMID [102] 96.11/97.79 91.84/94.07
+CACo [103] 90.88/95.05 88.28/91.94
+CROMA [104] 96.44/97.58 92.63/95.04
+SatLas [105] 94.96/97.38 92.16/94.70
+GFM [106] 95.47/97.09 -
+Scale-MAE [107] 96.44/97.58 92.63/95.04
+SkySense [2] 97.68/98.60 94.85/96.32
+RingMo [1] 96.90/98.34 94.25/95.67
+RingMo-Aerial 95.81/96.46 92.28/95.65
+The comparison results are shown in Tab.IX.
+Although RingMo-Aerial does not perform as well as specialized satellite models like SkySense in nadir-view classification, it still demonstrates strong performance, reflecting its robust generalization capability.
+5.2.4 Change Detection
+TABLE X: Change Detection: Comp. on LEVIR-CD and CDD
+Methods LEVIR-CD CDD
+F1 IoU F1 IoU
+FC-Siam-Di [108] 86.31 83.31 70.61 54.57
+FC-Siam-Conc [108] 83.69 76.77 75.11 60.14
+IFNet [109] 88.13 82.93 90.30 71.91
+STANet [110] 87.30 77.40 84.12 72.22
+BIT [111] 89.31 89.37 88.90 80.01
+SNUNet [112] 88.16 87.17 83.89 72.11
+ChangeFormer [113] 90.40 88.80 89.83 81.53
+RingMo [1] 89.53 82.38 90.90 84.09
+RingMo-Aerial 92.71 87.10 94.62 90.09
+The comparison results are shown in Tab.X.
+RingMo-Aerial combined with BIT can achieve change detection results that far exceed other methods.
+In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection.
+When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated into our architecture.
+The comparison of visual results on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks, as shown in the appendix.
+The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images.
+In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
+RingMo-Aerial’s commendable performance not only makes it a key player in the field but also highlights the importance of the impact of foundation model innovation.
+These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
+5.2.5 3D-Reconstruction
+TABLE XI: 3D-Reconstruction: Comp. on LuoJia-MVS and WHU
+Number of Views Model LuoJia-MVS WHU
+MAE $<$0.6m $<$3-interval MAE $<$0.6m $<$3-interval
+Three views * PatchmatchNet[114] 0.252 92.7 87.2 0.173 96.5 94.8
+Fast-MVSNet[115] 0.194 95.7 92.0 0.184 95.5 94.1
+MVSNet[116] 0.172 96.1 92.4 0.190 95.0 94.3
+R-MVSNet[117] 0.177 96.0 93.5 0.183 95.3 93.5
+RED-Net[118] 0.109 98.2 96.9 0.112 98.1 97.9
+Cas-MVSNet[119] 0.103 98.4 97.1 0.111 97.7 97.6
+RingMo-Aerial 0.095 98.7 97.8 0.121 97.8 97.5
+Five views * PatchmatchNet[114] 0.283 90.4 84.1 0.160 96.9 95.0
+Fast-MVSNet[115] 0.357 84.6 74.9 0.157 96.1 95.6
+MVSNet[116] 0.270 91.2 81.8 0.160 95.8 95.5
+R-MVSNet[117] 0.259 92.3 86.7 0.173 95.4 93.8
+RED-Net[118] 0.156 94.9 90.5 0.104 98.1 97.9
+Cas-MVSNet[119] 0.141 97.9 95.4 0.095 97.8 97.8
+RingMo-Aerial 0.122 98.1 96.5 0.091 98.0 97.9
+    ‘Three Views’ and ‘Five Views’ represent the number of views for the network’s input is three and five, respectively.
+Tab.XI provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets.
+Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error.
+On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet.
+It improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs.
+0.141), 0.2% (98.1% vs.
+97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a significant improvement.
+In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%.
+It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
+5.2.6 Object Tracking
+TABLE XII: Object Tracking: Comp. on Visdrone-MOT
 Methods MOTA IDF1
@@
 RingMo-Aerial 52.6 67.8
-TABLE XIV: Object Tracking: Comp. on AIR-MOT
+TABLE XIII: Object Tracking: Comp. on AIR-MOT
 Methods MOTA IDF1
@@
 RingMo-Aerial 68.6 81.3
-TABLE XV: Object Tracking: Comp. on AIR-HSAO
+TABLE XIV: Object Tracking: Comp. on AIR-HSAO
 Methods MOTA IDF1
@@
 RingMo-Aerial 60.7 74.4
-Figure 14: Qualitative results for DroneVehicle-IR when using different methods
-Figure 15: Qualitative results for SARData when using different methods
-V-F3 Comparison With the SOTA
-An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance.
-On unimodal datasets such as VisDrone-DET (as shown in Table VIII) and UAVDT (as shown in Table IX), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods.
-The experimental results further demonstrate that RingMo-Aerial, when compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
-What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Table X), and PP-PicoDet-L follows, while YOLOX-L and YOLOX-Tiny show the worst results.
-Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
-For the multimodal dataset, DroneVehicle (as shown in Table XI), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods.
-Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial.
-This indicates that RingMo-Aerial holds significant advantages in handling cross-modality object detection tasks, particularly in UAV image applications.
-Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
-In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Table XII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as YOLOX-S, Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance.
-Particularly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and an mAP50 of 84.4%, highlighting its superiority in SAR modality.
-These results substantiate the efficacy of the RingMo-Aerial backbone in processing SAR imagery.
-Furthermore, the performance of RingMo-Aerial on multimodal datasets reinforces its viability as a universal pre-trained model, especially in enhancing precision and robustness for object detection tasks across diverse and intricate environments.
-These findings underscore the potential application value of RingMo-Aerial in the field of remote sensing, specifically for aerial object detection tasks utilizing SAR data.
-Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields.
-For optical images, as shown in Figure 13, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets.
-In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups.
-Similarly, RingMo-Aerial accurately detects distant vehicles in road scenarios, whereas RingMo underperforms.
-RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images.
-As shown in Figures 14 and 15, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence.
-The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images.
-RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for aerial remote sensing.
-V-G Object Tracking
-V-G1 Datasets and Evaluation Metrics
-For Object Tracking, following aerial dataset is employed.
+The comparison results on the VisDrone-MOT dataset are shown in Tab.XII.
+RingMo-Aerial can achieve the best tracking results, far exceeding other methods.
+In the AIR-MOT dataset (as shown in Tab.XIII), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack.
+Meanwhile, in the AIR-HSAO dataset (as shown in Tab.XIV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods.
+Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
+Furthermore, the visualization results of the AIR-MOT sequences demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios.
+The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions.
+This visualization provides qualitative solid evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in ARS videos.
+5.3 Ablation Experiments
+5.3.1 Ablation Experiments on pre-training
+Figure 13: The ablation experiment on the UAVID dataset and UperNet is used as the segmentation head.
+RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
+TABLE XV: Ablation Experiments
+UAVID Visdrone-DET
+mAcc mIoU mAP mAP50
+RingMo 71.6 61.9 28.4 45.2
+RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
+RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
+TABLE XVI: Ablation Experiments with the proposed ARS-Adapter
+UAVID Visdrone-DET
+mAcc mIoU mAP mAP50
+RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
+RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
+We utilize UperNet model on UAVID and Cascade RCNN model on Visdrone-DET.
+The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL.
+This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
+The comparison is shown in Fig.13 and Tab.XV.
+The experimental findings reveal a progressive convergence speed and performance enhancement across the three foundation models.
+Notably, the RingMo-Aerial model trained with MIM and CL exhibits the swiftest convergence and attains the optimal performance.
+This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pre-training strategy, affirming their mutually reinforcing nature and consistent effectiveness.
+In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter.
+Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL.
+Comparative results on the UAVID and VisDrone-DET datasets are presented in Tab.XVI.
+The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
+5.3.2 Performance Comparison Between ARS-Adapter and Other Fine-Tuning Methods
+TABLE XVII: Comp. between fine-tuning methods
+Methods Parameters UAVID VisDrone-DET
+mAcc mIoU mAP mAP50
+Froezn 0% 61.2 49.3 21.5 36.2
+Full 100% 76.5 66.4 31.1 51.8
+Adapter 0.39% 68.9 57.7 28.9 49.1
+AdaptFormer 0.39% 71.3 62.1 30.3 50.8
+ARS-Adapter * 0.41% 74.3 64.4 31.0 51.4
+    Although ARS-Adapter is less accurate than global fine-tuning, it significantly outperforms other fine-tuning methods when the proportion of fine-tuning parameters is almost constant.
+In the context of ARS, traditional model fine-tuning presents specific challenges.
+These include the propensity for overfitting due to the extensive pre-training on diverse tasks, which can lead to a model that is overly tailored to the nuances of the training data, thereby compromising its applicability to new, unseen data.
+Overfitting is particularly problematic in ARS, where models are not only expected to generalize across various tasks but also to adapt to the unique perspectives and resolutions inherent in aerial imagery.
+This contrasts with ApRS, which typically operates within a more constrained perspective and resolution, reducing the risk of overfitting to specific views.
+The ARS-Adapter, therefore, addresses these challenges by introducing a parameter-efficient fine-tuning approach that minimizes overfitting risks while maintaining high performance, as evidenced by the experimental results.
+This section compares ARS-Adapter against other fine-tuning methods, using UAVID and VisDrone-DET datasets with UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively.
+The aim is to assess the effectiveness of ARS-Adapter compared to conventional fine-tuning techniques.
+The experimental results are presented in Tab.XVII.
+For clarity, we compare ARS-Adapter with full fine-tuning, zero fine-tuning, and two other fine-tuning methods (Adapter and AdaptFormer).
+Full fine-tuning is a performance upper bound, while zero fine-tuning provides a reference for minimal parameter adjustment.
+ARS-Adapter requires a similar proportion of parameters(0.41%) to be adjusted as Adapter(0.39%) and AdaptFormer(0.39%), but achieves superior fine-tuning efficiency.
+Additionally, the performance loss compared to full fine-tuning(100%) is limited to approximately 2%, demonstrating that ARS-Adapter can achieve high performance with minimal parameter adjustments.
+These results indicate that ARS-Adapter maintains high performance while significantly reducing the number of parameters requiring adjustment, demonstrating its efficiency in fine-tuning.
+6 Conclusion
+The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks.
+Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein.
+Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks.
+This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks.
+In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research.
+Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment.
+Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.
+Appendix A Semantic Segmentation
+A.1 Datasets and Evaluation Metrics
+For semantic segmentation, the UAVID [140], UDD6 [141] and FloodNet[142] datasets are employed.
+  • UAVID.
+The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid, and 15 test), and each sequence contains ten pictures.
+In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
+A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
+  • UDD6.
+The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images.
+All of the images are 4K resolution.
+In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
+The training set is used for training, and the validation set is used for testing.
+  • FloodNet.
+The FloodNet dataset contains 500 high-resolution remote sensing images with a resolution of 4K.
+The dataset is divided into training, validation, and test sets, with the training set containing 300 images, the validation set containing 100 images, and the test set containing 100 images.
+Images are annotated as flood and non-flood areas, supporting multi-class flood impact assessment.
+All images are cropped into 1024×1024 pixel slices with a 256-pixel overlap in the experiment, facilitating the training and evaluation of machine learning models.
+In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc), and average accuracy (aAcc).
+A.2 Experiment Settings
+The experiment of the semantic segmentation is based on segmentation [143], UperNet [144], and Mask2Former [145] are employed as the segmentation framework.
+The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used).
+The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
+In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
+Figure 14: Qualitative comparison of semantic segmentation results on the UAVID dataset using different methods.
+The figure illustrates differences in capturing details such as vegetation, moving vehicles, and buildings, with highlighted regions emphasizing variations in detail and performance.
+Appendix B Object Detection
+B.1 Datasets and Evaluation Metrics
+To conduct a comprehensive comparison of object detection methods, we selected four unimodal benchmark datasets, including the VisDrone-DET[147], UAVDT[148] and ShipDataset[149], as well as one multimodal benchmark dataset, DroneVehicle[92] and our collection of SAR airborne remote sensing data, to demonstrate the performance of the representative methods discussed in this paper.
+  • VisDrone-DET.
+The Visdrone-DET dataset is the detection task part of the VisDrone dataset.
+The resolution of the images is about 2k.
+The RingMo-Aerial experiment crops the original image to 640×640 and uses multi-scale cropping and resizing for enhancement.
+The training set and validation set are used for training, and the labeled test-dev set is used for testing.
+  • UAVDT.
+The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with an average resolution of 1080 × 540 pixels.
+The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
+  • ShipDataset.
+The ShipDataset, captured at a resolution of $3840\times 2160$, encompasses five distinct scenarios in Shanghai, China, characterized by varying shooting angles and lighting conditions.
+This dataset has $96.97\%$ small ships.
+The images within this dataset predominantly feature congested ship traffic.
+The RingMo-Aerial experiment uses multi-scale resizing for enhancement.
+  • DroneVehicle.
+The DroneVehicle dataset is a large-scale dataset specifically designed for drone-based RGB-Infrared cross-modality vehicle detection tasks.
+It includes 28,439 pairs of RGB-Infrared images, totaling 56,878 images, covering various urban scenarios such as roads, residential areas, and parking lots under different lighting conditions, including day, night, and dark night.
+The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
+  • AerialSARData.
+To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside.
+This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection.
+The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels.
+The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
+In our object detection experiments, the evaluation metrics employed are mean average precision (mAP), mAP at an IoU threshold of 50 (mAP50), and mAP at an IoU threshold of 75 (mAP75).
+For datasets containing smaller objects, we also utilize mAPs, which represents the mean average precision for detecting small objects (with pixels smaller than 32×32).
+B.2 Experiment Settings
+The experiment of the object detection is based on mmdetection [150], Cascade RCNN [67] and DDQ DETR [78] are employed as the detection framework.
+The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used).
+For DDQ DETR, 36 epochs were spanned with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used).
+The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
+In addition, the learning rate is decreased by 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
+Figure 15: RingMo-Aerial combined with Cascade RCNN achieves superior detection accuracy compared to the original RingMo model, particularly in complex environments and small object scenarios.
+It demonstrates reduced false positives and improved bounding box alignment with Ground Truth, highlighting its effectiveness for aerial imagery detection.
+Figure 16: Qualitative comparison of object detection results on infrared images from the DroneVehicle-IR dataset using different methods.
+The figure highlights differences in the detection quality of vehicles, road elements, and pedestrian areas, with emphasized regions showcasing variations in accuracy and detail among the compared methods.
+Figure 17: Qualitative comparison of object detection results on the AerialSARData dataset using different methods.
+The figure highlights differences in detection performance for SAR images, particularly focusing on identifying objects of interest with highlighted regions showing variations in accuracy and detection quality among the methods compared.
+Appendix C Change Detection
+C.1 Datasets and Evaluation Metrics
+For semantic segmentation, the LEVIR [110] and CDD [151] datasets are employed.
+  • LEVIR.
+The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset.
+It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5-meter resolution.
+Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels.
+Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases.
+  • CDD.
+The CDD dataset is a comprehensive collection of bi-temporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels.
+These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches.
+This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs.
+To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
+C.2 Experiment Settings
+Our model adopts BIT with Swin Transformer as the backbone.
+The experiments use overall accuracy (OA) and F1 and IoU scores about change detection as the primary evaluation metrics.
+The model is trained with an AdamW optimizer with a weight decay of 0.01 and beta values equal to (0.9, 0.999).
+The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
+Figure 18: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset.
+That shows the results of aerial images, where RingMo-Aerial outperforms ResNet18 and RingMo, with fewer false positives and false negatives.
+RingMo-Aerial closely matches the Ground Truth, demonstrating the best accuracy in change Detection.
+The black region represents true negative, the white region represents true positive, the green region represents false negative, the red region represents false positive.
+Appendix D Scene Classification
+D.1 Datasets and Evaluation Metrics
+For scene classification, the AID [152] and RESISC-45 [153] datasets are employed.
+  • AID. The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification. It contains images collected from Google Earth and covers 30 scenes.
+  • NWPU-RESISC45.
+The NWPU-RESISC45 dataset is a publicly available Remote Sensing Image Scene Classification (RESISC) benchmark.
+The dataset contains 31,500 images covering 45 scene classes, with 700 images in each class.
+In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively.
+In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively.
+The experiments use overall accuracy (OA) as a performance metric, the most commonly used performance metric in scene classification.
+D.2 Experiment Settings
+The experiment of the scene classification is based on mmpretrain [154], Cascade RCNN [67] is employed as the classification framework.
+The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs).
+The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05.
+In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
+Appendix E 3D-Reconstruction
+E.1 Datasets and Evaluation Metrics
+For 3D-Reconstruction, the and WHU [118] datasets are employed.
+  • WHU.
+The WHU dataset is a synthetic aerial dataset designed for large-scale multi-view stereo reconstruction tasks.
+It was generated from a highly accurate 3D surface model and includes a complete set of aerial images as well as cropped sub-image sets specifically for deep learning applications.
+The images in the dataset cover an area of approximately 6.7 × 2.2 square kilometers in Meitan County, Guizhou Province, China, with a ground resolution of 0.1 meters.
+The WHU dataset provides crucial data support for the study of large-scale Earth surface and urban reconstruction and can be utilized for training and testing multi-view stereo matching algorithms.
+  • LuoJia-MVS.
+The LuoJia-MVS dataset comprises 7,972 five-view images, each with a spatial resolution of 10 cm, accompanied by pixel-wise depth information and precise camera parameters.
+These images were generated from a highly accurate digital surface model (DSM) constructed using thousands of stereo aerial images.
+During the experiment, we used the WHU dataset[118] and the LuoJia-MVS dataset[155] to verify the effectiveness of our proposed method.
+WHU dataset[118] is the first large-scale aerial multi-view stereo dataset for remote sensing.
+It was collected from Meitan, Guizhou, China, and was obtained by taking large-scale photos with a tilted five-eye camera on a drone platform.
+The area of the entire dataset is approximately $6.7\times 2.2km^{2}$, and the image resolution is about 0.1m.
+In addition, the tilt angle of each camera to the central camera is 40°.
+The WHU dataset gathers high-rise buildings, mountains, forests, and factory scenes, providing a good benchmark for remote sensing multi-view stereo.
+LuoJia-MVS[155] dataset is a large-scale multi-view stereo dataset collected from a variety of landform types.
+It was collected in a part of Baiyun, Guiyang, Guizhou, China.
+The image size of the entire dataset is $768\times 384$, and the image resolution is about 0.1m.
+The data consists of five views and pixel-by-pixel depth labels.
+Land types cover cultivated land, forest land, industrial areas, residential areas, etc.
+Figure 19: Qualitative results on the target tracking dataset AIR-MOT using RingMo-Aerial.
+Four video sequences are selected, where each target is indicated with different colors.
+The category and target ID are displayed above the bounding boxes, demonstrating RingMo-Aerial’s ability to effectively distinguish and track individual targets.
+E.2 Experiment Settings
+Our work is coded based on the PyTorch framework.
+The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1.
+During the training process, our method uses the Adam optimizer[156] for parameter optimization and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001.
+In addition, our method is consistent with the existing multi-stage networks[119].
+First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively.
+Then, deep optimization is performed through three stages of profound hypotheses, and the number of hypotheses is 48, 32, and 8, respectively.
+During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
+Appendix F Object Tracking
+F.1 Datasets and Evaluation Metrics
+For Object Tracking, the following aerial dataset is employed.
   • VisDrone-MOT.
-The VisDrone-MOT[169] dataset consists of 96 sequences, including 5 class such as pedestrian, car, van, bus and truck.
-The VisDrone-MOT train and val set are used for training and test set is used for test.
+The VisDrone-MOT[157] dataset consists of 96 sequences, including five classes such as pedestrian, car, van, bus, and truck.
+The VisDrone-MOT train and val set are used for training, and the test set is used for the test.
 In addition to the aerial dataset, we conducted experiments on two satellite remote sensing datasets.
@@
   • AIR-HSAO.
-The AIR-HSAO dataset[168] is designed for tracking high-speed aerial objects, derived from the JiLin-1 satellite.
+The AIR-HSAO dataset[139] is designed for tracking high-speed aerial objects derived from the JiLin-1 satellite.
 It contains 199 videos, though fewer instances are present due to the rapid movement of these objects.
@@
 In our experiment of object tracking, Multiple Object Tracking Accuracy (MOTA) and Identification Flatness (IDF1) are used as evaluation metrics.
-V-G2 Experiment Settings
+F.2 Experiment Settings
 In the proposed experiment, object tracking is predicated on the detection model of RingMo-Aerial on VisDrone-DET.
-Specifically, the detection model trained by RingMo-Aerial will undergo an additional 5 epochs of training on the VisDrone-MOT dataset, during which the backbone network parameters will be frozen to ensure consistency.
+Specifically, the detection model trained by RingMo-Aerial will undergo an additional five epochs of training on the VisDrone-MOT dataset, during which the backbone network parameters will be frozen to ensure consistency.
 The training will be conducted with a batch size of 8, distributed as 2 per GPU across 4 NVIDIA A40 GPUs.
-Throughout the training phase, original images from the VisDrone-MOT dataset will be processed, accompanied by multi-scale augmentation to enhance robustness.
+Throughout the training phase, original images from the VisDrone-MOT dataset will be processed, and multi-scale augmentation will be accompanied to enhance robustness.
 Upon completion of training on the VisDrone-MOT dataset, the detection model will be subjected to inference on the VisDrone-MOT test set using the MOT tools provided by mmdetection [150].
 Subsequently, the model’s multi-object tracking performance will be quantified and analyzed.
-V-G3 Comparison With the SOTA
-The comparison results on the VisDrone-MOT dataset are shown in Table XIII.
-RingMo-Aerial can achieve the best tracking results far exceeding other methods.
-In the AIR-MOT dataset (as shown in TableXIV), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack.
-Meanwhile, in the AIR-HSAO dataset (as shown in TableXV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods.
-Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
-Figure 16: Visualization Results on AIR-MOT: Four video sequences are selected, where each target in the sequences is indicated with different colors.
-The category and target ID are displayed above the bounding boxes.
-Futhermore, the visualization results of the AIR-MOT sequences (as shown in Figure16) clearly demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios.
-The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions.
-This visualization provides strong qualitative evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in aerial remote sensing videos.
-V-H ARS-Adapter
-TABLE XVI: Comp. between fine-tuning methods
-Methods Parameters UAVID VisDrone-DET
-mAcc mIoU mAP mAP50
-Froezn 0% 61.2 49.3 21.5 36.2
-Full 100% 76.5 66.4 31.1 51.8
-Adapter 0.39% 68.9 57.7 28.9 49.1
-AdaptFormer 0.39% 71.3 62.1 30.3 50.8
-ARS-Adapter 0.41% 74.3 64.4 31.0 51.4
-As previously discussed, this paper proposes the ARS-Adapter, a parameter-efficient fine-tuning method tailored for ARS vision.
-This section focuses on comparing the performance of the ARS-Adapter with other fine-tuning methods in ARS vision applications.
-The comparisons are conducted using the UAVID and VisDrone-DET datasets, employing UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively.
-The evaluation aims to highlight the efficacy and advantages of the ARS-Adapter in enhancing ARS vision tasks compared to conventional fine-tuning techniques.
-[…]
-V-I Ablation Experiments
-Figure 17: The ablation experiment on UAVID dataset and UperNet is used as the segmentation head.
-RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
-TABLE XVII: Ablation Experiments
-UAVID Visdrone-DET
-mAcc mIoU mAP mAP50
-RingMo 71.6 61.9 28.4 45.2
-RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
-RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
-TABLE XVIII: Ablation Experiments with the proposed ARS-Adapter
-UAVID Visdrone-DET
-mAcc mIoU mAP mAP50
-RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
-RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
-The ablation experiments in this study are conducted utilizing the UperNet segmentation model on the UAVID dataset and the Cascade RCNN detection model on the Visdrone-DET dataset.
-The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL.
-This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
-The comparison is shown in Figure 17 and Table XVII.
-The experimental findings reveal a progressive enhancement in convergence speed and performance across the three foundation models.
-Notably, the RingMo-Aerial model trained with both MIM and CL exhibits the swiftest convergence and attains the optimal performance.
-This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pretraining strategy, affirming their mutually reinforcing nature and consistent effectiveness.
-In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter.
-Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL.
-Comparative results on the UAVID and VisDrone-DET datasets are presented in Table XVIII.
-The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
-VI Conclusion
-The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks.
-Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein.
-[…]
-Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks.
-This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks.
-In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research.
-Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment.
-Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.
+Acknowledgments
+This work was supported by the National Nature Science Foundation of China under Grant 62331027 and 62301538.

### v4

@@
 Abstract
-Aerial Remote Sensing (ARS) vision tasks pose significant challenges due to the unique characteristics of their viewing angles.
+Aerial Remote Sensing (ARS) vision tasks present significant challenges due to the unique viewing angle characteristics.
 Existing research has primarily focused on algorithms for specific tasks, which have limited applicability in a broad range of ARS vision applications.
-This paper proposes the RingMo-Aerial model, aiming to fill the gap in foundation model research in the field of ARS vision.
-By introducing the Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism and an affine transformation-based contrastive learning pre-training method, the model’s detection capability for small targets is enhanced and optimized for the tilted viewing angles characteristic of ARS.
-Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and effectiveness in various ARS vision tasks.
+This paper proposes RingMo-Aerial, aiming to fill the gap in foundation model research in the field of ARS vision.
+A Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) mechanism is introduced to strengthen the model’s capacity for small-object representation.
+Complementarily, an affine transformation-based contrastive learning method improves its adaptability to the tilted viewing angles inherent in ARS tasks.
+Furthermore, the ARS-Adapter, an efficient parameter fine-tuning method, is proposed to improve the model’s adaptability and performance in various ARS vision tasks.
 Experimental results demonstrate that RingMo-Aerial achieves SOTA performance on multiple downstream tasks.
@@
 These challenges highlight the need for specialized intelligent models tailored to ARS applications.
-While recent foundational models in computer vision and RS have shown potential, they are often inadequate for ARS due to its unique requirements.
+While recent foundational models in computer vision and RS have shown potential capabilities, they are often inadequate for ARS due to its unique requirements.
 These models benefit from pre-training on vast amounts of data and large-scale parameters, enabling them to adapt more rapidly and effectively to various changes.
@@
 Unlike traditional comprehensive fine-tuning methods, ARS-Adapter achieves efficient model fine-tuning by adjusting fewer than 5% of the parameters, significantly improving the training efficiency of downstream tasks while reducing the risk of overfitting.
+Traditional efficient fine-tuning methods, such as Adapter and AdaptFormer, are insufficient in addressing the domain gap between self-supervised pretraining and downstream tasks.
+This domain gap arises from the inconsistency in feature extraction preferences between pretraining and task-specific objectives.
+ARS-Adapter mitigates this issue to a certain extent.
 ARS-Adapter applies to all layers of the model: in the attention layers, it employs a traditional Adapter[6] structure that operates parallel to the attention layers; in the Multi-Layer Perceptron (MLP) layers, it utilizes a bottleneck structure that includes MLP downsampling, deep convolutional feature extraction, and MLP upsampling.
@@
   • ARS-Adapter: The ARS-Adapter provides a fine-tuning approach tailored to ARS models, enhancing efficiency and reducing overfitting risks.
-To evaluate the effectiveness of the model, we performed six downstream tasks using 13 ARS datasets and 5 ApRS datasets with different viewpoints, resolutions, and modes.
+To evaluate the effectiveness of the model, we performed six downstream tasks using 12 ARS datasets and 5 ApRS datasets with different viewpoints, resolutions, and modes.
 The test results show that RingMo-Aerial achieves SOTA performance on all ARS datasets and reaches the leading level on ApRS datasets, which proves its strong generalization ability for intelligent interpretation of aerial and ApRS images.
@@
 In the field of remote sensing (RS), models such as RingMo [1] and SkySense [2] have adopted the visual Transformer architecture and self-supervised pre-training strategies.
-Several novel architectures, such as ConvNeXt [25] and Mamba[26], have been proposed.
+At this stage, methods integrating convolution into ViT architectures also began to emerge.
+Representative works such as CvT [25], Conformer [26], and CoaT [27] enhance local feature modeling by embedding convolution into token projections, using parallel convolution-attention branches, or inserting convolution in early encoder stages.
+However, most of these approaches rely on stacked 3×3 convolutions, which suffer from a mismatch between the theoretical and effective receptive fields, limiting their ability to capture broader contexts or object-level structures [28].
+Besides, several novel architectures, such as ConvNeXt [29] and Mamba[30], have been proposed.
 However, these newer models have yet to achieve widespread adoption due to technical immaturity.
@@
 By integrating ViT and Swin Transformer, RingMo enhanced the accuracy of various downstream RS tasks, laying a foundation for developing large models in RS scenes.
-2.2.3 The combination of Contrastive Learning &\&& Masked Image Modeling
+2.2.3 The combination of Contrastive Learning &\& Masked Image Modeling
 Combining CL and MIM has shown significant advantages in current research, particularly in enhancing feature extraction performance.
@@
 The image is from the VisDrone dataset.
+The yellow box represents the tokens occupied by the target.
 The unique imaging properties of ARS photography stem from the inherent properties of aerial perspective.
@@
 As shown in Fig.5, similar objects at different distances from the ARS are rendered at different sizes in the image.
+This phenomenon results in distant objects occupying fewer tokens in the Vision Transformer’s input sequence, while closer objects occupy more.
 Figure 6: The same target will be represented differently in the image if its distance from the lens and imaging angle are different.
 As the horizontal distance increases, the number of pixels occupied by the height content increases and then decreases, and the height content occupies the most pixels at the horizontal distance $l=\sqrt{H(H-h)}$.
-As shown in Fig.6, when the camera is positioned at a height $H$ and the target object has a height $h$, the angular size $\alpha$ of the target’s height in the camera’s field of view varies as a function of the horizontal distance $l$ between the target and the camera.
-At different horizontal distances, the imaging size corresponding to the same target height differs.
+As shown in Fig.6, when the camera is positioned at a height $H$ and the target object has a height $h$, the angular size $\theta$ of the target’s height in the camera’s field of view varies as a function of the horizontal distance $l$ between the target and the camera.At different horizontal distances, the imaging size corresponding to the same target height differs.
 Specifically, when the target is directly beneath the camera, its height is not captured in the image.
@@
 By considering the distances from the target’s lowest and highest points to the camera, given by $x=\sqrt{H^{2}+l^{2}}$ and $y=\sqrt{(H-h)^{2}+l^{2}}$ respectively, we can derive the relationship between the viewing angle $\alpha$ and the horizontal distance $l$ between the target and the camera:
-Using the cosine rule:
-$x^{2}+y^{2}-2xy\cos\alpha=h^{2}$ (1)
+$x^{2}+y^{2}-2xy\cos\theta=h^{2}$ (1)
 Substituting the values of $x$ and $y$, The final cosine formula becomes:
-$\cos\alpha=\frac{H^{2}-Hh+l^{2}}{\sqrt{H^{2}+l^{2}}\cdot\sqrt{(H-h)^{2}+l^{2}}}$ (2)
-By setting $a=H$ and $b=H-h$, we have the target function is:
-$f(l)=\cos\alpha=\frac{ab+l^{2}}{\sqrt{a^{2}+l^{2}}\cdot\sqrt{b^{2}+l^{2}}}$ (3)
+$\cos\theta=\frac{H^{2}-Hh+l^{2}}{\sqrt{H^{2}+l^{2}}\cdot\sqrt{(H-h)^{2}+l^{2}}}$ (2)
+By setting $a=H$ and b=Hhb=H-h, we have the target function is:
+$f(l)=\cos\theta=\frac{ab+l^{2}}{\sqrt{a^{2}+l^{2}}\cdot\sqrt{b^{2}+l^{2}}}$ (3)
 Taking the derivative of $f(l)$:
@@
 $l^{2}(a-b)^{2}=ab(a-b)^{2}$ (7)
-Since $a=H$ and $b=H-h$, and $a>b$, it follows that $a-b=h>0$.
+Since $a=H$ and b=Hhb=H-h, and $a>b$, it follows that ab=h>0a-b=h>0.
 When $f^{\prime}(l)=0$, we obtain:
@@
 $l=\sqrt{ab}=\sqrt{H(H-h)}$ (9)
-This completes the derivation of the maximum value $\alpha$ and the corresponding minimum value $\cos\alpha$.
+This completes the derivation of the maximum value $\theta$ and the corresponding minimum value $\cos\theta$.
 From Eq.
-(2), when $l=0$, $\alpha$ is equal to 0.
-From Eqs.(4) and (9), $\alpha$ increases with $l$ is less than $\sqrt{H(H-h)}$, $\alpha$ is maximum when $l=\sqrt{H(H-h)}$, and decreases with $l$ when $l$ is greater than $\sqrt{H(H-h)}$, and decreases with $l$ in the case of $l\rightarrow\infty$ when $\alpha$ is again equal to 0.
-Assuming that the optical axis of the camera is always at the center of the target viewpoint $\alpha$, the pixel change of the target height information in the picture is the same as the change of the corresponding viewpoint $\alpha$ of the target.
+(2), when $l=0$, $\theta$ is equal to 0.
+From Eqs.(4) and (9), $\theta$ increases with $l$ is less than $\sqrt{H(H-h)}$, $\theta$ is maximum when $l=\sqrt{H(H-h)}$, and decreases with $l$ when $l$ is greater than $\sqrt{H(H-h)}$, and decreases with $l$ in the case of $l\rightarrow\infty$ when $\theta$ is again equal to 0.
+Assuming that the optical axis of the camera is always at the center of the target viewpoint $\theta$, the pixel change of the target height information in the picture is the same as the change of the corresponding viewpoint $\theta$ of the target.
 This indicates that if the same target is at different distances from the lens and imaging angles, it will have different representations in the image.
@@
 4.1 Method Overview
-The proposed RingMo-Aerial model serves as a foundational framework with broad applicability.
+The proposed RingMo-Aerial serves as a foundational framework with broad applicability.
 Its implementation involves three main stages: model design, pre-training, and fine-tuning for downstream tasks.
@@
 As shown in Fig.7 Pre-training Stage 1, considering the unique characteristics of the ARS perspective and based on the window attention mechanisms (W-MSA and SW-MSA), a new attention mechanism called Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) is introduced to enhance the model’s capability in handling diverse aerial scenarios.
+It is an extension to standard Multi-Head Self-Attention (MSA), defined as:
+$\mathrm{FE\text{-}MSA}(X)$ $=\mathrm{MSA}(Y_{\mathrm{conv}})$
+$=\mathrm{MSA}\bigl{(}X+\mathrm{DWConv}_{7\times 7}(X)\bigr{)}.$ (10)
+The input $X$ is first processed through a residual depth-wise $7\times 7$ convolution branch, which expands the local receptive field and compensates for high-frequency components.
+The enhanced feature map $Y_{\mathrm{conv}}$ is then fed into the MSA layer to enable global information interaction.
 Figure 8: The effect of FE-MSA.
@@
 At the same time, FE-MSA enhances the model’s receptive field for small objects (purple box).
-Figure 9: Positive samples for contrastive learning are obtained after the image’s affine transformation, cropping, and resizing.
 Frequency-Enhanced Multi-Head Self-Attention (FE-MSA) module.
@@
 The FE-MSA mechanism introduces frequency domain processing to the existing window attention, indirectly expanding the receptive field for smaller objects.
+The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, improving the model’s performance in complex aerial scenes.
+Fig.9 presents a comparative analysis of the feature extraction capabilities of RingMo-Aerial and RingMo models in processing multiview ARS images.
+The heatmaps generated by RingMo-Aerial show more focused attention on critical areas, such as vehicles and buildings, as indicated by the bright regions.
+In contrast, the RingMo in the lower half has a more diffuse focus, with less concentration on specific features.
+This comparison demonstrates the superior ability of RingMo-Aerial to handle complex aerial scenes, enabling better feature extraction and object recognition in multiview ARS images.
+These results visually confirm the improvements introduced by RingMo-Aerial, particularly in addressing challenges such as varying scales and perspectives in aerial imagery.
+In ARS tasks such as object detection and semantic segmentation, small targets are far more prevalent and exhibit more extreme scale variation than in natural or satellite imagery.
+Standard Vision Transformer MSA or Swin Transformer Window Attention limits each token’s initial receptive field to a single patch, severely hindering edge and detail capture for small targets.
+Introducing an additional convolutional branch serves to (1) spatially expand the receptive field and (2) bolster the model’s ability to focus attention on fine-scale structures.
+From a frequency-domain perspective, small targets manifest stronger high-frequency energy, yet attention mechanisms tend to attenuate such components.
+By inserting a residual $7\times 7$ depth-wise separable convolution prior to attention, FE-MSA enhances band-pass filtering in the high-frequency range, improving detail modeling for small targets.
 Fig.7(b) illustrates the RingMo-Aerial Block structure, including two window attention mechanisms optimized by the FE-MSA mechanism.
 Given $z^{(l-1)}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$, where $l$ is the id of each RingMo-Aerial Block.
-Taking the first stage as an example, the input feature vector $z^{l}$ adopts the same slicing parameters as the standard Swin-Transformer, resulting in four-channel slices $C=\{c_{1},c_{2},c_{3},c_{4}\}$.
-After linear normalization, it enters the FE-MSA module.
-For each patch with a dimension of $C$, a $7\times 7$ convolution kernel is used for depthwise convolution to obtain the feature map $Fz^{(l-1)}=\text{Concat}\{Fc_{1},Fc_{2},Fc_{3},Fc_{4}\}$.
+Taking the first stage as an example, the input feature vector $z^{(l-1)}$ adopts the same partitioning strategy as the standard Swin Transformer.
+After linear normalization, the feature is passed into the FE-MSA module.
+In this module, all channels are independently processed using a shared $7\times 7$ depthwise convolution kernel, and the resulting outputs are concatenated along the channel dimension.
 Then, W-MSA or SW-MSA is used to compute the attention map.
@@
 $z^{l}=\text{FE-MSA}(z^{l-1})$ (11)
-[…]
-The FE-MSA mechanism enables RingMo-Aerial to effectively capture edge and texture features of different targets in ARS images, improving the model’s performance in complex aerial scenes.
-Fig.10 presents a comparative analysis of the feature extraction capabilities of the RingMo-Aerial and RingMo models in processing multiview ARS images.
-The heatmaps generated by RingMo-Aerial show more focused attention on critical areas, such as vehicles and buildings, as indicated by the bright regions.
-In contrast, the RingMo in the lower half has a more diffuse focus, with less concentration on specific features.
-This comparison demonstrates the superior ability of RingMo-Aerial to handle complex aerial scenes, enabling better feature extraction and object recognition in multiview ARS images.
-These results visually confirm the improvements introduced by RingMo-Aerial, particularly in addressing challenges such as varying scales and perspectives in aerial imagery.
+$\text{FE-MSA}(z^{l-1})=\text{SoftMax}\left(\frac{qk^{T}}{\sqrt{d}}+B\right)v$ (12)
+In the $l$-th RingMo-Aerial Block, the input feature map $z^{l-1}\in\mathbb{R}^{\frac{W}{4}\times\frac{H}{4}\times C}$ is first enhanced via a depthwise convolution in the frequency domain, after which multi-head self-attention is performed within each window: queries $q$, keys $k$, and values $v$ are obtained by linear projections of the enhanced features, their dot-products are scaled by $\sqrt{d}$, augmented with a relative position bias $B$, and normalized by SoftMax, then subsequently weighted-summed with $v$.
 4.3 Pre-training
 Figure 9: Comparative analysis of feature heatmaps extracted from multi-view aerial imagery using RingMo-Aerial and RingMo.
-The feature heatmaps extracted by the RingMo-Aerial model are located in the upper half, while those from the RingMo model are in the lower half of the figure.
+The feature heatmaps extracted by RingMo-Aerial are located in the upper half, while those from the RingMo model are in the lower half of the figure.
 In these images, the brighter the region, the more the model is focusing on that particular feature, indicating that these areas have higher feature importance for the model.
@@
 Perspective transformation first defines the source point and the destination point, which are defined as:
-[…]
+$\text{src\_points}=\begin{bmatrix}x_{0}&y_{0}\\ x_{1}&y_{1}\\ x_{2}&y_{2}\\ x_{3}&y_{3}\end{bmatrix},\quad\text{dst\_points}=\begin{bmatrix}u_{0}&v_{0}\\ u_{1}&v_{1}\\ u_{2}&v_{2}\\ u_{3}&v_{3}\end{bmatrix}$ (13)
 The transformation relationship between the source point and the destination point used in the proposed method is:
-[…]
+$\text{dst\_points}=\begin{bmatrix}u_{0}&v_{0}\\ u_{1}&v_{1}\\ u_{2}&v_{2}\\ u_{3}&v_{3}\end{bmatrix}=\begin{bmatrix}x_{0}&y_{0}\\ x_{1}&y_{1}\\ \alpha x_{2}+(1-\alpha)x_{3}&y_{2}\\ \alpha x_{3}+(1-\alpha)x_{2}&y_{3}\end{bmatrix}$ (14)
 Among them, $\alpha$ is the transformation coefficient, and the value range is randomly between 0.05 and 0.35.
@@
 $\begin{bmatrix}u\\ v\\ 1\end{bmatrix}=M\cdot\begin{bmatrix}x\\ y\\ 1\end{bmatrix}$ (17)
-We obtain the source image $I_{i}$ and the transformed image $I_{i}^{\prime}$ through the aforementioned affine transformation process.
-The source image $I_{i}$ and its corresponding transformed image $I_{i}^{\prime}$ form a positive sample pair, while the remaining sample pairs $I_{i}$ and $I_{j}^{\prime}$ (where $i\neq j$) form negative sample pairs.
-By calculating the contrastive loss, we guide the model in learning the invariant properties of the transformation process, thereby training the model’s ability to extract visual physical characteristics.
+Figure 11: Positive samples for contrastive learning are obtained after the image’s affine transformation, cropping, and resizing.
+When generating positive sample pairs via affine transformation, we first apply a random coefficient $\alpha\in[0.05,0.35]$ to compress the bottom width of the image to a specified proportion of its original size while preserving the top width, thereby creating a pixel-level “magnification” effect for distant small objects.
+The transformed image is then center-cropped to a rectangular region and rescaled to match the model’s input dimensions, effectively removing irrelevant background and further enhancing the visibility of small objects.
+By sampling $\alpha$ from a diversified range, we achieve multi-scale coverage—from subtle magnification to global micro-distortion—which enables the model to learn semantically consistent embeddings across varying perspectives and spatial resolutions, thereby improving its ability to perceive distant small targets in aerial imagery.
+For negative samples, we select all full, transformed images from other instances within the same mini-batch.
+Specifically, as illustrated in Figure 11, positive samples are generated by first applying an affine transformation that warps the image into an isosceles trapezoid while keeping the top width unchanged.
+This operation compresses the lower part of the image and stretches the upper part.
+The transformed image is then cropped to retain only the central rectangular region, which is subsequently resized to form the final positive sample.
+Although this procedure discards certain areas, it effectively amplifies the upper region of the image—i.e., distant small targets in aerial scenes.
+Incorporating such positive samples into contrastive pre-training helps the model to better capture scale-invariant features and geometric invariance, thereby improving its ability to detect small objects under complex aerial viewpoints.
+We obtain the source image $I_{i}$ and its transformed counterpart $I_{i}^{\prime}$ via the above affine process.
+The pair $(I_{i},I_{i}^{\prime})$ constitutes a positive sample, while the pairs $(I_{i},I_{j}^{\prime})$ where $i\neq j$, constitute negative samples.
+By optimizing a contrastive loss, the model is encouraged to learn view- and scale-invariant visual representations, thus enhancing its ability to extract stable physical characteristics from aerial scenes under diverse imaging conditions.
+Specifically, the contrastive loss used in our method is defined as follows:
+$\mathcal{L}_{\text{contrast}}=-\log\frac{\exp(\text{sim}(\mathbf{h}_{i},\mathbf{h}_{i}^{\prime})/\tau)}{\sum\limits_{j=1,j\neq i}^{N}\exp(\text{sim}(\mathbf{h}_{i},\mathbf{h}_{j}^{\prime})/\tau)}$ (18)
+where $\mathbf{h}_{i}$ and $\mathbf{h}_{i}^{\prime}$ are the normalized representations of a positive pair, $\text{sim}(\mathbf{h}_{i},\mathbf{h}_{j})=\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\,\|\mathbf{h}_{j}\|}$ is the cosine similarity between two vectors, $\tau$ is a temperature hyper parameter, and NNis the batch size.
 4.4 Downstream Fine-tuning
@@
 It integrates down-sampling, activation (e.g., GELU), and up-sampling, optionally including convolution, to enrich feature representations while reducing computational overhead.
-TABLE I: Semantic Segmentation: Comp. on UAVID
+In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks.
+Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model.
+The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, shown as Fig.12.
+The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking.
+Like AdaptFormer [46], the introduced ARS-Adapter module operates in parallel with the model.
+The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer.
+Therefore, the forward propagation is:
+$\text{Adapter-MSA}=\text{Up}(\text{GELU}(\text{Down}(x)))$ (19a)
+$\text{Adapter-FFN}=\text{Up}(\text{GELU}(\text{Conv}(\text{GELU}(\text{Down}(x)))))$ (19b)
+Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling.
+In parallel with FFN, the adapter introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
+[…]
+5 Experiments
+5.1 Pre-training Implementation
+In the pre-training phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized.
+Before pre-training, we initialize the model parameters using Kaiming initialization.
+Subsequently, we perform 100 epochs of masked image modeling (MIM) pre-training.
+During MIM pre-training, the input image size is 224, and random cropping and scaling are applied to the input images.
+The mask size is set to 32 pixels.
+The batch size is 2048, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 2 epochs.
+We adopt a masking ratio of 0.6 as the default configuration.
+This value is selected based on preliminary experiments, balancing semantic abstraction and structural preservation in ARS images characterized by multi-scale sparsity.
+Following MIM pre-training, we continue with contrastive learning (CL) pre-training, jointly training for 36 epochs.
+The input image size remains 224, and similar random cropping and scaling are applied.
+The batch size is adjusted to 1024, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 1 epoch.
+Figure 13: Performance improvement of RingMo-Aerial over previous SOTA methods across various remote sensing datasets. The horizontal axis denotes the percentage change in performance.
+5.2 Performance on Single/Multi-modal Downstream Tasks
+We validated the performance of RingMo-Aerial across six downstream tasks, covering both single-modal and multi-modal tasks, and involving a variety of data modalities and tasks commonly encountered in remote sensing scenarios.
+Fig.
+13 presents a comprehensive comparison between RingMo-Aerial and previous state-of-the-art (SOTA) methods on 17 datasets.
+It clearly demonstrates that RingMo-Aerial achieves superior performance, particularly in ARS imagery characterized by diverse viewing angles and object scales.
+Notably, RingMo-Aerial yields significant gains in semantic segmentation (e.g., +4.66 mIoU on FloodNet), object detection (e.g., +3.6 mAP on VisDrone-DET), and change detection tasks (e.g., +4.02 F1 on CDD), highlighting its strong generalization and representation capabilities across complex airborne environments.
+Detailed descriptions of the datasets, experimental setups, table of comparison with state-of-the-art methods and additional visualizations for each task are provided in the Appendix (Semantic Segmantation C, Object Detection D, Change Detection E, Scene Classification F, 3D-Reconstruction G and Object Tracking H).
+5.3 Ablation Experiments
+5.3.1 Ablation Experiments on pre-training
+Figure 14: The ablation experiment on the UAVID dataset and UperNet is used as the segmentation head.
+RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
+TABLE I: Ablation Experiments of pretrain methods
+UAVID Visdrone-DET
+mAcc mIoU mAP mAP50
+RingMo 71.6 61.9 28.4 45.2
+RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
+RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
+TABLE II: Ablation Experiments with the proposed ARS-Adapter
+UAVID Visdrone-DET
+mAcc mIoU mAP mAP50
+RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
+RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
+We utilize UperNet model on UAVID and Cascade RCNN model on Visdrone-DET.
+The comparative analysis involves three foundation models: RingMo pre-trained with MIM [1], RingMo-Aerial solely pre-trained with MIM, and RingMo-Aerial pre-trained with both MIM and CL.
+This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
+The comparison is shown in Fig.14 and Tab.I.
+The experimental findings reveal a progressive convergence speed and performance enhancement across the three foundation models.
+Notably, RingMo-Aerial trained with MIM and CL exhibits the swiftest convergence and attains the optimal performance.
+This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pre-training strategy, affirming their mutually reinforcing nature and consistent effectiveness.
+In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter.
+Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL.
+Comparative results on the UAVID and VisDrone-DET datasets are presented in Tab.II.
+The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
+5.3.2 Performance Comparison Between ARS-Adapter and Other Fine-Tuning Methods
+TABLE III: Comp. between fine-tuning methods
+Methods Parameters UAVID VisDrone-DET
+mAcc mIoU mAP mAP50
+Froezn 0% 61.2 49.3 21.5 36.2
+Full 100% 76.5 66.4 31.1 51.8
+Adapter 0.39% 68.9 57.7 28.9 49.1
+AdaptFormer 0.39% 71.3 62.1 30.3 50.8
+ARS-Adapter * 0.41% 74.3 64.4 31.0 51.4
+    Although ARS-Adapter is less accurate than global fine-tuning, it significantly outperforms other fine-tuning methods when the proportion of fine-tuning parameters is almost constant.
+In the context of ARS, traditional model fine-tuning presents specific challenges.
+These include the propensity for overfitting due to the extensive pre-training on diverse tasks, which can lead to a model that is overly tailored to the nuances of the training data, thereby compromising its applicability to new, unseen data.
+Overfitting is particularly problematic in ARS, where models are not only expected to generalize across various tasks but also to adapt to the unique perspectives and resolutions inherent in aerial imagery.
+This contrasts with ApRS, which typically operates within a more constrained perspective and resolution, reducing the risk of overfitting to specific views.
+The ARS-Adapter, therefore, addresses these challenges by introducing a parameter-efficient fine-tuning approach that minimizes overfitting risks while maintaining high performance, as evidenced by the experimental results.
+This section compares ARS-Adapter against other fine-tuning methods, using UAVID and VisDrone-DET datasets with UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively.
+The aim is to assess the effectiveness of ARS-Adapter compared to conventional fine-tuning techniques.
+The experimental results are presented in Tab.III.
+For clarity, we compare ARS-Adapter with full fine-tuning, zero fine-tuning, and two other fine-tuning methods (Adapter and AdaptFormer).
+Full fine-tuning is a performance upper bound, while zero fine-tuning provides a reference for minimal parameter adjustment.
+ARS-Adapter requires a similar proportion of parameters(0.41%) to be adjusted as Adapter(0.39%) and AdaptFormer(0.39%), but achieves superior fine-tuning efficiency.
+Additionally, the performance loss compared to full fine-tuning(100%) is limited to approximately 2%, demonstrating that ARS-Adapter can achieve high performance with minimal parameter adjustments.
+These results indicate that ARS-Adapter maintains high performance while significantly reducing the number of parameters requiring adjustment, demonstrating its efficiency in fine-tuning.
+5.3.3 Ablation of Mask ratio
+TABLE IV: Object detection performance of RingMo-Aerial under different masking ratios on the DroneVehicle dataset (modality: RGB+IR). The result at 0.6 is from the original configuration.
+Mask Ratio mAP mAP50 mAP75 mAPs mAPm mAPl
+0.5 0.489 0.728 0.566 0.196 0.493 0.664
+0.6 0.492 0.735 0.570 0.228 0.497 0.684
+0.7 0.487 0.726 0.567 0.195 0.490 0.668
+0.8 0.490 0.731 0.567 0.235 0.494 0.681
+TABLE V: Semantic segmentation performance of RingMo-Aerial under different masking ratios on the UAVid dataset. The result at 0.6 is from the original configuration.
+Mask Ratio mIoU (%) mAcc (%)
+0.5 68.29 78.18
+0.6 68.70 78.70
+0.7 67.72 77.60
+0.8 68.56 78.28
+To investigate the influence of masking ratio in pretraining, we conduct additional experiments under different ratios (0.5, 0.6, 0.7, 0.8).
+Results on both object detection (DroneVehicle, RGB+IR) and semantic segmentation (UAVid) are presented in Tab.
+IV and Tab.
+V The masking ratio of 0.6 achieves the highest mAP and mIoU, confirming its effectiveness across tasks.
+Interestingly, higher ratios (e.g., 0.8) slightly improve small object detection, while lower ratios (e.g., 0.5) provide marginal gains on medium-sized targets.
+These findings suggest that ARS imagery, due to its multi-scale structure and sparse layout, requires more carefully tuned masking strategies than natural images.
+5.3.4 Ablation of depthwise convolution placement
+TABLE VI: Ablation study on the placement of depthwise convolution in the FE-MSA module. Results are reported on the DroneVehicle dataset with RGB+IR modality.
+Variant mAP mAP50 mAP75 mAPs mAPm mAPl
+Original* 0.492 0.735 0.570 0.228e 0.497 0.684
+Conv1* 0.485 0.720 0.561 0.183 0.485 0.664
+Conv2* 0.483 0.719 0.556 0.193 0.485 0.670
+    Original: Depthwise convolution is inserted after norm1 and before the attention module (used in our final model).
+Conv1: Depthwise convolution is placed after the FFN block.
+Conv2: Depthwise convolution is placed between the multi-head attention and FFN.
+To analyze the effect of depthwise convolution placement in the FE-MSA module, we compare three configurations using object detection performance on the DroneVehicle dataset (RGB+IR).
+As shown in Tab.
+VI, the results demonstrate that the Original configuration achieves the best overall performance, with 0.492 mAP and consistent advantages across all scales, especially for large objects.
+Although Conv2 slightly improves small object detection (mAPs = 0.193), it performs worse overall.
+This confirms that applying depthwise convolution early, before semantic modeling via attention and FFN, provides better spatial enhancement for ARS images.
+Therefore, we adopt the Original configuration in our final design.
+5.3.5 Ablation of depthwise convolution kernel size
+TABLE VII: Object detection performance of RingMo-Aerial under different size depthwise separation convolution kernels on the DroneVehicle dataset (modality: RGB+IR).
+The results at 7 are from the original configuration.
+Kernel Size mAP mAP50 mAP75 mAPs mAPm mAPl
+3 0.485 0.722 0.564 0.192 0.489 0.661
+5 0.490 0.729 0.568 0.179 0.494 0.673
+7 0.492 0.735 0.570 0.228 0.497 0.684
+9 0.490 0.733 0.566 0.192 0.493 0.684
+TABLE VIII: Semantic segmentation performance of RingMo-Aerial under different size depthwise separation convolution kernels on the UAVid dataset. The result at 7 is from the original configuration.
+Kernel Size mIoU (%) mAcc (%)
+3 67.18 77.77
+5 67.57 78.21
+7 68.70 78.70
+9 67.03 78.15
+To investigate the impact of different sizes of depthwise separable convolution kernels during pre-training, we conducted additional experiments using kernels of sizes 3, 5, 7, and 9.
+The experimental results for object detection (DroneVehicle, RGB+IR) and semantic segmentation (UAVid) are presented in Table.
+VII and Table.
+VIII, respectively.
+The results show that a kernel size of 7 achieves the highest mAP and mIoU, demonstrating its effectiveness across different tasks.
+These findings suggest that kernels of sizes 3 and 5 provide insufficient receptive fields, limiting the ability to extract complete and informative features.
+Conversely, a kernel size of 9 results in an excessively large receptive field, introducing redundant information that interferes with the extraction of relevant features.
+Therefore, a kernel size of 7 strikes the optimal balance, enabling effective feature extraction and leading to superior performance.
+6 Conclusion
+RingMo-Aerial presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks.
+Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein.
+Our experiments demonstrate that RingMo-Aerial achieves SOTA performance across various datasets for ARS vision tasks.
+This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks.
+In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research.
+Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment.
+Additionally, we will explore the extension of RingMo-Aerial to other vision tasks and scenarios, aiming to expand its application scope.
+Acknowledgments
+This work was supported by the National Nature Science Foundation of China under Grant 62331027 and 62301538.
+Appendix A Frequency-Enhanced Multi-Head Self-Attention
+The following content provides a derivation analysis of the FE-MSA module from both the spatial domain and frequency domain perspectives.
+A.1 Spatial-Domain Analysis
+Self-attention performs “global” modeling by assigning correlation weights at the patch level.
+For small targets, fine-scale features remain under-integrated.
+Inserting a $7\times 7$ convolution before attention effectively enlarges the receptive field at each position from $1\times 1$ (a single point within a patch) to $7\times 7$, while a residual connection preserves the original features:
+$Y_{\mathrm{conv}}=X+\mathrm{DWConv}_{7\times 7}(X).$ (20)
+For each spatial location $i$,
+$y_{i}=x_{i}+\bigl{[}\mathrm{DWConv}_{7\times 7}(X)\bigr{]}_{i},$ (21)
+where $\bigl{[}\mathrm{DWConv}_{7\times 7}(X)\bigr{]}_{i}$ aggregates high-frequency (edge, texture) information from a $7\times 7$ neighborhood around $i$, thereby enhancing local edges and structural details of small targets.
+The convolutional branch thus accentuates small-target features in the local context, enabling subsequent attention to focus more precisely.
+Consider the standard self-attention formulation for head $h$:
+$A^{(h)}=\mathrm{Softmax}\Bigl{(}Q^{(h)}{K^{(h)}}^{\top}/\sqrt{d}\Bigr{)},$ (22)
+$Q^{(h)}=XW_{Q}^{(h)},$ (23)
+$K^{(h)}=XW_{K}^{(h)}.$ (24)
+After adding the convolutional branch,
+$Q^{\prime(h)}$ $=YW_{Q}^{(h)}=XW_{Q}^{(h)}+\mathrm{DWConv}(X)W_{Q}^{(h)},$ (25)
+$K^{\prime(h)}$ $=YW_{K}^{(h)}=XW_{K}^{(h)}+\mathrm{DWConv}(X)W_{K}^{(h)}.$ (26)
+For any query–key pair $(i,j)$, the modified dot-product logit expands to
+$\ell^{\prime}_{ij}$ $=Q^{\prime}_{i}\cdot K^{\prime}_{j}$
+$=\underbrace{x_{i}W_{Q}W_{K}^{\top}x_{j}}_{\ell_{ij}}$
+$\quad+\underbrace{x_{i}W_{Q}W_{K}^{\top}[\mathrm{Conv}(X)]_{j}+[\mathrm{Conv}(X)]_{i}W_{Q}W_{K}^{\top}x_{j}}_{\Delta_{ij}}$
+$\quad+\underbrace{[\mathrm{Conv}(X)]_{i}W_{Q}W_{K}^{\top}[\mathrm{Conv}(X)]_{j}}_{\delta_{ij}}.$ (27)
+Here, the original term $\ell_{ij}$ captures pure attention on the raw features, while the enhancement term $\Delta_{ij}$ and the double-enhancement term $\delta_{ij}$ arise from the convolutional branch.
+Since $\mathrm{Conv}(X)]$ is amplified at high-frequency regions corresponding to small-target edges and textures, these additional terms preferentially boost logits for positions inside or related to small targets.
+Consequently,
+$\ell^{\prime}_{ij}>\ell_{ij}$ (28)
+for small-target regions, and the softmax scaling further amplifies attention weights at these locations, resulting in more comprehensive aggregation of small-target information.
+A.2 Frequency-Domain Analysis
+Denote the 2D Fourier transform of $X\in\mathbb{R}^{H\times W\times C}$ as
+$\widehat{X}(\omega_{x},\omega_{y})=\sum_{m=0}^{H-1}\sum_{n=0}^{W-1}X[m,n]e^{-j(\omega_{x}m+\omega_{y}n)}.$ (29)
+Small targets concentrate energy at higher spatial frequencies; letting
+$E(\omega)$ $=\|\widehat{X}(\omega)\|^{2},$ (30)
+$E_{\mathrm{high}}$ $=\int_{\|\omega\|\geq\omega_{0}}E(\omega)\,d\omega$
+$>E_{\mathrm{low}}=\int_{\|\omega\|<\omega_{0}}E(\omega)\,d\omega.$ (31)
+Patch-level self-attention acts as a low-pass filter, responding strongly to low frequencies while attenuating high frequencies. A rough model is
+$H_{\mathrm{attn}}(\omega)=e^{-\alpha\|\omega\|},\quad\alpha>0,$ (32)
+or
+$H_{\mathrm{att}}(\omega)=\begin{cases}1,&\|\omega\|<\omega_{0},\\ \beta,&\|\omega\|\geq\omega_{0},\quad 0\leq\beta<1.\end{cases}$ (33)
+In contrast, a depth-wise separable $7\times 7$ convolution employs a spatial filter $h[m,n]$ whose frequency response
+$H_{\mathrm{conv}}(\omega_{x},\omega_{y})=\sum_{m=-3}^{3}\sum_{n=-3}^{3}h[m,n]e^{-j(\omega_{x}m+\omega_{y}n)}$ (34)
+acts as a band-pass filter with relatively high gain in the high-frequency range.
+Therefore, by preceding attention with this convolution, FE-MSA compensates for the attention filter’s high-frequency attenuation:
+$\bigl{|}H_{\mathrm{conv}}(\omega)\bigr{|}_{\|\omega\|\geq\omega_{0}}>\bigl{|}H_{\mathrm{attn}}(\omega)\bigr{|}_{\|\omega\|\geq\omega_{0}}.$ (35)
+This frequency-domain enhancement better preserves small-target detail during global aggregation.
+Appendix B Parameter analysis of ARS-Adapter
+Let $C$ be the number of channels (i.e., embedding dimension) of the encoder layer.
+The Transformer encoder layer includes a depthwise separable convolution layer with $C$ channels, and four linear layers with C channels.
+The Feed-Forward Network (FFN) layer includes a linear layer with input $C$ and output $4C$, and another linear layer with input $4C$ and output $C$.
+The total number of parameters in a single encoder layer is then:
+$\mathrm{P_{e}}=C\times 7\times 7+4\times C\times C+2\times C\times 4C=49C+12C^{2}$ (36)
+Besides, the total parameters added by the ARS-Adapter are calculated as follows:
+For each encoder layer, the added components include: Downsampling linear layers $2\times(C\times C/64)$, Upsampling linear layers $2\times(C\times C/64)$ and Depthwise separable convolution layer $\frac{C}{64}\times 3\times 3$.
+Then, the total additional parameters for each encoder layer are:
+$\mathrm{P_{a}}=4\times(C\times C/64)+\frac{C}{64}\times 3\times 3=\frac{9}{64}C+\frac{1}{16}C^{2}$ (37)
+Finally, if $C=512$, we can calculate the proportion of added parameters compared to the total parameters of the model:
+$\mathrm{Proportion}=\frac{\mathrm{P_{a}}}{\mathrm{P_{e}}+\mathrm{P_{a}}}\approx 0.5\%$ (38)
+Appendix C Semantic Segmentation
+TABLE IX: Semantic Segmentation: Comp. on UAVID
 Methods Building Tree Clutter Road Low vegetation Static car Moving car Human mAcc mIoU
@@
 RingMo(U)*[1] 90.9 85.4 82.6 74.3 73.1 56.8 81.6 73.7 71.9 58.2 69.9 58.3 75.4 67.5 27.4 21.0 71.6 61.9
+RingMo(M)*[1] 93.0 83.6 87.9 77.4 74.9 62.4 86.4 78.7 78.5 61.5 62.3 53.0 86.4 78.7 36.5 28.2 75.7 64.9
 RingMo-Aerial(U)* 92.6 83.8 88.5 77.2 78.4 65.5 87.5 80.2 73.3 60.7 72.5 65.3 84.4 71.6 34.8 26.9 76.5 66.4
@@
 Same below.
-In the methodological section of the paper, we expand upon the fine-tuning strategies employed for downstream tasks.
-Beyond the conventional approach of fine-tuning the entire model, this study introduces an adapter-based fine-tuning method tailored for an aviation foundation model.
-The proposed ARS-Adapter comprises a standard Adapter module coupled with an additional depthwise convolution module, shown as Fig.12.
-The latter is specifically designed to enhance the model’s performance in tasks such as detection and tracking.
-Like AdaptFormer [55], the introduced ARS-Adapter module operates in parallel with the model.
-The Adapter module featuring the depthwise convolution is paralleled with the MLP layer, while the Adapter module without this convolution module is paralleled with the Attention layer.
-Therefore, the forward propagation is:
-[…]
-[…]
-Specifically, the downsampling layer performs 16 times downsampling in the embedding dimension, and the corresponding upsampling layer performs 16 times upsampling.
-In parallel with FFN, the adapter introduces an additional convolution layer, which uses 7×7 depth-separated convolution to reduce fine-tuning parameters while introducing more spatial information.
-[…]
-5 Experiments
-TABLE II: Semantic Segmentation: Comp. on UDD6
+TABLE X: Semantic Segmentation: Comp. on UDD6
 Methods Other Facade Road Vegetation Vehicle Roof mAcc mIoU
@@
 RingMo(U) [1] 91 80.4 89.4 74.6 85.5 73.2 89.6 78.4 77.8 62.3 91.1 82.9 87.4 75.3
+RingMo(M)[1] 72.18 59.86 84.77 71.23 84.62 71.17 95.14 90.22 79.66 67.73 93.53 87.52 84.98 74.62
 RingMo-Aerial(U) 90.4 81.1 89.7 77.1 88.7 77.4 90.4 80.2 81.3 68.9 95.9 86.3 89.4 78.5
 RingMo-Aerial(M) 90.7 80.9 90.2 77.4 88.4 77.1 90.5 80.8 82.1 69.4 96.3 87.8 89.7 78.9
-TABLE III: Semantic Segmentation: Comp. on Floodnet
+TABLE XI: Semantic Segmentation: Comp. on Floodnet
 Methods Building Road Water Tree Vehicle Pool Grass OA mIoU
@@
 RingMo(U) [1] 71.95 66.80 44.22 74.24 61.97 77.97 23.11 33.01 85.22 88.46 59.83
+RingMo(M)*[1] 50.80 70.23 39.15 78.56 68.10 74.10 47.14 45.13 84.94 87.50 62.02
 RingMo-Aerial(U) 74.48 77.06 55.78 83.00 70.35 81.41 53.96 58.08 87.04 90.40 71.24
 RingMo-Aerial(M) 76.17 77.90 48.57 82.62 62.75 81.70 63.16 65.79 87.16 90.43 71.76
-5.1 Pre-training Implementation
-In the pre-training phase of our model, approximately 5.4 million image patches, as mentioned earlier, are utilized.
-Before pre-training, we initialize the model parameters using Kaiming initialization.
-Subsequently, we perform 100 epochs of masked image modeling (MIM) pre-training.
-During MIM pre-training, the input image size is 224, and random cropping and scaling are applied to the input images.
-The mask size is set to 32 pixels.
-The batch size is 2048, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 2 epochs.
-Following MIM pre-training, we continue with contrastive learning (CL) pre-training, jointly training for 36 epochs.
-The input image size remains 224, and similar random cropping and scaling are applied.
-The batch size is adjusted to 1024, with a base learning rate of 2e-4, a cosine annealing learning rate schedule, and a warm-up period of 1 epoch.
-5.2 Performance on Single/Multi-modal Downstream Tasks
-We validated the performance of RingMo-Aerial across six downstream tasks, covering both single-modal and multi-modal tasks, and involving a variety of data modalities and tasks commonly encountered in remote sensing scenarios.
-Detailed descriptions of the datasets, experimental setups, and additional visualizations for each task are provided in the Appendix.
-5.2.1 Semantic Segmantation
+Figure 15: Qualitative comparison of semantic segmentation results on the UAVID dataset using different methods.
+The figure illustrates differences in capturing details such as vegetation, moving vehicles, and buildings, with highlighted regions emphasizing variations in detail and performance.
+C.1 Datasets and Evaluation Metrics
+For semantic segmentation, the UAVID [66], UDD6 [67] and FloodNet[68] datasets are employed.
+  • UAVID.
+The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid, and 15 test), and each sequence contains ten pictures.
+In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
+A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
+  • UDD6.
+The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images.
+All of the images are 4K resolution.
+In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
+The training set is used for training, and the validation set is used for testing.
+  • FloodNet.
+The FloodNet dataset contains 500 high-resolution remote sensing images with a resolution of 4K.
+The dataset is divided into training, validation, and test sets, with the training set containing 300 images, the validation set containing 100 images, and the test set containing 100 images.
+Images are annotated as flood and non-flood areas, supporting multi-class flood impact assessment.
+All images are cropped into 1024×1024 pixel slices with a 256-pixel overlap in the experiment, facilitating the training and evaluation of machine learning models.
+In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc), and average accuracy (aAcc).
+C.2 Experiment Settings
+The experiment of the semantic segmentation is based on segmentation [69], UperNet [70], and Mask2Former [71] are employed as the segmentation framework.
+The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used).
+The AdamW [72] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
+In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
+C.3 Experiment Results
 Comparative results on the UAVID, UDD6, and FloodNet datasets are presented in Tabs.
-I, II, and III, respectively.
+IX,X, and XI, respectively.
 The combination of RingMo-Aerial with Mask2Former achieves segmentation outcomes that significantly outperform those of other methods.
 Furthermore, we present the qualitative results for a selection of methods in different scenarios, which primarily comprise roads, secondary roads, and car parks.
-RingMo-Aerial consistently produces superior segmentation maps.
+As illustrated in Fig.15, RingMo-Aerial consistently produces superior segmentation maps.
 Among these methods, the RingMo-Aerial approach demonstrates excellent segmentation accuracy and the ability to recognize small targets effectively when combining different architectures.
@@
 This illustrates the significant advantages of our method in complex urban scenes and diverse object classes.
-5.2.2 Object Detection
-TABLE IV: Object Detection: Comp. on VisDrone-DET
+Appendix D Object Detection
+D.1 Datasets and Evaluation Metrics
+To conduct a comprehensive comparison of object detection methods, we selected four unimodal benchmark datasets, including the VisDrone-DET[73], UAVDT[74] and ShipDataset[75], as well as one multimodal benchmark dataset, DroneVehicle[76] and our collection of SAR airborne remote sensing data, to demonstrate the performance of the representative methods discussed in this paper.
+  • VisDrone-DET.
+The Visdrone-DET dataset is the detection task part of the VisDrone dataset.
+The resolution of the images is about 2k.
+The RingMo-Aerial experiment crops the original image to 640×640 and uses multi-scale cropping and resizing for enhancement.
+The training set and validation set are used for training, and the labeled test-dev set is used for testing.
+  • UAVDT.
+The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with an average resolution of 1080 × 540 pixels.
+The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
+  • ShipDataset.
+The ShipDataset, captured at a resolution of $3840\times 2160$, encompasses five distinct scenarios in Shanghai, China, characterized by varying shooting angles and lighting conditions.
+This dataset has $96.97\%$ small ships.
+The images within this dataset predominantly feature congested ship traffic.
+The RingMo-Aerial experiment uses multi-scale resizing for enhancement.
+  • DroneVehicle.
+The DroneVehicle dataset is a large-scale dataset specifically designed for drone-based RGB-Infrared cross-modality vehicle detection tasks.
+It includes 28,439 pairs of RGB-Infrared images, totaling 56,878 images, covering various urban scenarios such as roads, residential areas, and parking lots under different lighting conditions, including day, night, and dark night.
+The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
+  • AerialSARData.
+To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside.
+This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection.
+The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels.
+The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
+TABLE XII: Object Detection: Comp. on VisDrone-DET
 Methods Backbone mAP mAP50 mAP75
@@
     RingMo-Aerial (C), (A), (D) represent the RingMo aerial backbone with Cascade RCNN detection algorithm, ARS-Adapter and Cascade RCNN algorithm, and DDQ DETR algorithm. Same below.
-TABLE V: Object Detection: Comp. on UAVDT
+TABLE XIII: Object Detection: Comp. on UAVDT
 Methods Backbone mAP mAP50 mAP75
@@
 RingMo-Aerial (D) RingMo-Aerial 25.2 39.9 27.1
-TABLE VI: Object Detection: Comp. on ShipDataset
+TABLE XIV: Object Detection: Comp. on ShipDataset
 Methods Backbone mAP mAP50 mAPs
@@
 YOLOX-Tiny[99] CSPDarkNet-53 71.1 96.4 30.1
-PP-PicoDet-L[87] ESNet 89.7 98.9 67.3
+$P$-PicoDet-L[101] ESNet 89.7 98.9 67.3
 Cascade R-CNN[81] Swin-Base 88.3 99.0 62.6
 RingMo-Aerial(C) RingMo-Aerial 83.7 99.1 72.3
-TABLE VII: Object Detection: Comp. on DroneVehicle
+TABLE XV: Object Detection: Comp. on DroneVehicle
 Methods Modality Backbone mAP50
@@
 YOLOv7[97] RGB EfficientNet-B0 68.5
-$S^{2}$et[90] RGB RetinaNet 57.3
+$S^{2}$ANet[104] RGB RetinaNet 57.3
 Oriented R-CNN[105] RGB ResNet101 62.3
@@
 YOLOv7[97] IR EfficientNet-B0 66.7
-$S^{2}$et[90] IR RetinaNet 64.8
+$S^{2}$ANet[104] IR RetinaNet 64.8
 Oriented R-CNN[105] IR ResNet101 65.5
@@
 Cascade R-CNN[81] RGB+IR Swin-Base 69.8
-RingMo-Aerial(C) RGB+IR RingMo-Aerial 71.6
-TABLE VIII: Object Detection: Comp. on AerialSARData
+RingMo-Aerial(C) RGB+IR RingMo-Aerial 73.5
+TABLE XVI: Object Detection: Comp. on AerialSARData
 Methods Backbone mAP mAP50
@@
     RingMo-Aerial (A), (C) represent the RingMo-Aerial backbone with ARS-Adapter and Cascade RCNN algorithm.
+In our object detection experiments, the evaluation metrics employed are mean average precision (mAP), mAP at an IoU threshold of 50 (mAP50), and mAP at an IoU threshold of 75 (mAP75).
+For datasets containing smaller objects, we also utilize mAPs, which represents the mean average precision for detecting small objects (with pixels smaller than 32×32).
+D.2 Experiment Settings
+The experiment of the object detection is based on mmdetection [109], Cascade RCNN [81] and DDQ DETR [92] are employed as the detection framework.
+The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used).
+For DDQ DETR, 36 epochs were spanned with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used).
+The AdamW [72] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
+In addition, the learning rate is decreased by 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
+Figure 16: RingMo-Aerial combined with Cascade RCNN achieves superior detection accuracy compared to the original RingMo model, particularly in complex environments and small object scenarios.
+It demonstrates reduced false positives and improved bounding box alignment with Ground Truth, highlighting its effectiveness for aerial imagery detection.
+Figure 17: Qualitative comparison of object detection results on infrared images from the DroneVehicle-IR dataset using different methods.
+The figure highlights differences in the detection quality of vehicles, road elements, and pedestrian areas, with emphasized regions showcasing variations in accuracy and detail among the compared methods.
+Figure 18: Qualitative comparison of object detection results on the AerialSARData dataset using different methods.
+The figure highlights differences in detection performance for SAR images, particularly focusing on identifying objects of interest with highlighted regions showing variations in accuracy and detection quality among the methods compared.
+D.3 Experiment Results
 An analysis of all tables reveals that the detection algorithms based on the proposed method as a pre-trained model have achieved state-of-the-art (SOTA) performance.
-On unimodal datasets such as VisDrone-DET (as shown in Tab.IV) and UAVDT (as shown in Tab.V), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods.
+On unimodal datasets such as VisDrone-DET (as shown in Tab.XII) and UAVDT (as shown in Tab.XIII), RingMo-Aerial combined with DDQ DETR (RingMo-Aerial(D))significantly outperforms other methods.
 The experimental results further demonstrate that RingMo-Aerial, compared to Swin-Base in Cascade RCNN and DDQ DETR, is a more suitable foundational model for UAV-based object detection tasks.
-What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Tab.VI), and PP-PicoDet-L follows.
+What’s more, RingMo-Aerial(C) exhibits the best performance in small object detection (mAPs) on the ShipDataset, which primarily targets small vessels (as shown in Tab.XIV), and PP-PicoDet-L follows.
 In contrast, YOLOX-L and YOLOX-Tiny show the worst results.
 Notably, RingMo-Aerial(C) consistently excels in small object detection across all tasks, further validating its superiority in complex scenarios.
-For the multimodal dataset, DroneVehicle (as shown in Tab.VII), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods.
+For the multimodal dataset, DroneVehicle (as shown in Tab.XV), RingMo-Aerial (C) achieves the best performance in the RGB+IR modality with an mAP50 of 71.6, significantly outperforming other methods.
 Although methods like Dual-YOLO and CIAN(OBB) also show strong performance in the RGB+IR modality, they still fall short compared to RingMo-Aerial.
@@
 Additionally, the outstanding performance of RingMo-Aerial in multimodal scenarios further validates its effectiveness as a general pre-trained model, particularly in providing higher accuracy and robustness in object detection tasks across diverse and complex environments.
-In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Tab.VIII), the RingMo-Aerial backbone network, integrated with various detection algorithms such as Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance.
+In the domain of Synthetic Aperture Radar (SAR) datasets (as shown in Tab.XVI), the RingMo-Aerial backbone network, integrated with various detection algorithms such as Cascade R-CNN, and the proposed ARS-Adapter, has demonstrated a notable enhancement in performance.
 Mainly, the RingMo-Aerial(C) configuration has achieved an mAP of 36.2% and a mAP50 of 84.4%, highlighting its superiority in SAR modality.
@@
 Additionally, we present the qualitative results of several methods across different scenarios, primarily focusing on roads and sports fields.
-For optical images, as shown in the appendix, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets.
+For optical images, as shown in Fig.16, the RingMo-Aerial detection model demonstrates superior performance, particularly in detecting small-scale targets.
 In the sports field scenario, RingMo-Aerial effectively identifies distant groups of people, even those not annotated in the test set, while RingMo fails to detect these groups.
@@
 RingMo-Aerial also demonstrates outstanding small target recognition capabilities for SAR and IR images.
-As shown in the appendix, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence.
+As shown in Fig.17 and 18, RingMo tends to misclassify densely overlapping vehicles at the end of the road, while RingMo-Aerial maintains a high level of class confidence.
 The visualized results confirm the model’s excellent performance on multimodal data, including optical, SAR, and IR images.
 RingMo-Aerial’s exceptional performance across multiple modalities proves its strong generalization ability and highlights its great potential as the first foundational model for ARS.
-5.2.3 Scene Classification
-TABLE IX: Scene Classification: Comp. on AID and RESISC-45
+Appendix E Change Detection
+E.1 Datasets and Evaluation Metrics
+For semantic segmentation, the LEVIR [110] and CDD [111] datasets are employed.
+  • LEVIR.
+The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset.
+It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5-meter resolution.
+Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels.
+Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases.
+  • CDD.
+The CDD dataset is a comprehensive collection of bi-temporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels.
+These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches.
+This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs.
+To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
+TABLE XVII: Change Detection: Comp. on LEVIR-CD and CDD
+Methods LEVIR-CD CDD
+F1 IoU F1 IoU
+FC-Siam-Di [112] 86.31 83.31 70.61 54.57
+FC-Siam-Conc [112] 83.69 76.77 75.11 60.14
+IFNet [113] 88.13 82.93 90.30 71.91
+STANet [110] 87.30 77.40 84.12 72.22
+BIT [114] 89.31 89.37 88.90 80.01
+SNUNet [115] 88.16 87.17 83.89 72.11
+ChangeFormer [116] 90.40 88.80 89.83 81.53
+RingMo [1] 89.53 82.38 90.90 84.09
+RingMo-Aerial 92.71 87.10 94.62 90.09
+E.2 Experiment Settings
+Our model adopts BIT with Swin Transformer as the backbone.
+The experiments use overall accuracy (OA) and F1 and IoU scores about change detection as the primary evaluation metrics.
+The model is trained with an AdamW optimizer with a weight decay of 0.01 and beta values equal to (0.9, 0.999).
+The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
+Figure 19: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset.
+That shows the results of aerial images, where RingMo-Aerial outperforms ResNet18 and RingMo, with fewer false positives and false negatives.
+RingMo-Aerial closely matches the Ground Truth, demonstrating the best accuracy in change Detection.
+The black region represents true negative, the white region represents true positive, the green region represents false negative, the red region represents false positive.
+E.3 Experiment Result
+The comparison results are shown in Tab.XVII.
+RingMo-Aerial combined with BIT can achieve change detection results that far exceed other methods.
+In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection.
+When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated into our architecture.
+The comparison of visual results on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks, as shown in Fig.19.
+The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images.
+In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
+RingMo-Aerial’s commendable performance not only makes it a key player in the field but also highlights the importance of the impact of foundation model innovation.
+These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
+Appendix F Scene Classification
+F.1 Datasets and Evaluation Metrics
+TABLE XVIII: Scene Classification: Comp. on AID and RESISC-45
 Model AID (OA) RESISC-45 (OA)
@@
 RingMo-Aerial 95.81/96.46 92.28/95.65
-The comparison results are shown in Tab.IX.
+For scene classification, the AID [129] and RESISC-45 [130] datasets are employed.
+  • AID. The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification. It contains images collected from Google Earth and covers 30 scenes.
+  • NWPU-RESISC45.
+The NWPU-RESISC45 dataset is a publicly available Remote Sensing Image Scene Classification (RESISC) benchmark.
+The dataset contains 31,500 images covering 45 scene classes, with 700 images in each class.
+In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively.
+In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively.
+The experiments use overall accuracy (OA) as a performance metric, the most commonly used performance metric in scene classification.
+F.2 Experiment Settings
+The experiment of the scene classification is based on mmpretrain [131], Cascade RCNN [81] is employed as the classification framework.
+The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs).
+The AdamW [72] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05.
+In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
+F.3 Experiment Results
+The comparison results are shown in Tab.XVIII.
 Although RingMo-Aerial does not perform as well as specialized satellite models like SkySense in nadir-view classification, it still demonstrates strong performance, reflecting its robust generalization capability.
-5.2.4 Change Detection
-TABLE X: Change Detection: Comp. on LEVIR-CD and CDD
-Methods LEVIR-CD CDD
-F1 IoU F1 IoU
-FC-Siam-Di [108] 86.31 83.31 70.61 54.57
-FC-Siam-Conc [108] 83.69 76.77 75.11 60.14
-IFNet [109] 88.13 82.93 90.30 71.91
-STANet [110] 87.30 77.40 84.12 72.22
-BIT [111] 89.31 89.37 88.90 80.01
-SNUNet [112] 88.16 87.17 83.89 72.11
-ChangeFormer [113] 90.40 88.80 89.83 81.53
-RingMo [1] 89.53 82.38 90.90 84.09
-RingMo-Aerial 92.71 87.10 94.62 90.09
-The comparison results are shown in Tab.X.
-RingMo-Aerial combined with BIT can achieve change detection results that far exceed other methods.
-In a rigorous comparative evaluation against current state-of-the-art technologies, RingMo-Aerial proved its capabilities in the specialized field of remote sensing change detection.
-When evaluated against LEVIR-CD and CDD datasets, the model demonstrated a clear competitive advantage, highlighting the robustness and effectiveness of the backbone network integrated into our architecture.
-The comparison of visual results on the LEVIR dataset reveals significant differences in the performance of models with different backbones in building detection and segmentation tasks, as shown in the appendix.
-The RingMo-Aerial model demonstrates superior accuracy, with fewer false positives (red areas) and false negatives (green areas), highlighting its advantages in processing remote sensing images.
-In contrast, the ResNet18 and Swin-Base models exhibit more false positives and false negatives, particularly in complex building layouts, where RingMo-Aerial shows greater robustness.
-RingMo-Aerial’s commendable performance not only makes it a key player in the field but also highlights the importance of the impact of foundation model innovation.
-These elements are critical to pushing the frontiers of remote sensing analysis and setting new benchmarks for future research and applications.
-5.2.5 3D-Reconstruction
-TABLE XI: 3D-Reconstruction: Comp. on LuoJia-MVS and WHU
+TABLE XIX: 3D-Reconstruction: Comp. on LuoJia-MVS and WHU
 Number of Views Model LuoJia-MVS WHU
-MAE $<$0.6m $<$3-interval MAE $<$0.6m $<$3-interval
+MAE <<0.6m <<3-interval MAE <<0.6m <<3-interval
 Three views * PatchmatchNet[132] 0.252 92.7 87.2 0.173 96.5 94.8
@@
     ‘Three Views’ and ‘Five Views’ represent the number of views for the network’s input is three and five, respectively.
-Tab.XI provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets.
+Appendix G 3D-Reconstruction
+G.1 Datasets and Evaluation Metrics
+For 3D-Reconstruction, the LuoJia-MVS[138] and WHU [136] datasets are employed.
+  • WHU.
+The WHU dataset is a synthetic aerial dataset designed for large-scale multi-view stereo reconstruction tasks.
+It was generated from a highly accurate 3D surface model and includes a complete set of aerial images as well as cropped sub-image sets specifically for deep learning applications.
+The images in the dataset cover an area of approximately 6.7 × 2.2 square kilometers in Meitan County, Guizhou Province, China, with a ground resolution of 0.1 meters.
+The WHU dataset provides crucial data support for the study of large-scale Earth surface and urban reconstruction and can be utilized for training and testing multi-view stereo matching algorithms.
+  • LuoJia-MVS.
+The LuoJia-MVS dataset comprises 7,972 five-view images, each with a spatial resolution of 10 cm, accompanied by pixel-wise depth information and precise camera parameters.
+These images were generated from a highly accurate digital surface model (DSM) constructed using thousands of stereo aerial images.
+During the experiment, we use three metrics, mean absolute error (MAE), <<0.6m, and <<3-interval, to evaluate the performance of the proposed method.
+G.2 Experiment Settings
+Our work is coded based on the PyTorch framework.
+The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1.
+During the training process, our method uses the Adam optimizer[139] for parameter optimization and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001.
+In addition, our method is consistent with the existing multi-stage networks[137].
+First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively.
+Then, deep optimization is performed through three stages of profound hypotheses, and the number of hypotheses is 48, 32, and 8, respectively.
+G.3 Experiment Results
+Tab.XIX provides a performance comparison of our method and other methods on the LuoJia-MVS and WHU datasets.
 Obviously, under the premise of taking five views as input, our method achieves the best estimation performance and the lowest prediction error.
-On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of $<$0.6m and $<$3-interval accuracy, which is better than the existing advanced method Cas-MVSNet.
-It improves MAE, $<$0.6m, $<$3-interval by 13.5% (0.122 vs.
+On the LuoJia-MVS dataset, our method achieves an MAE error of 0.122m, 98.1% and 96.5% of <<0.6m and <<3-interval accuracy, which is better than the existing advanced method Cas-MVSNet.
+It improves MAE, <<0.6m, <<3-interval by 13.5% (0.122 vs.
 0.141), 0.2% (98.1% vs.
 97.9%) and 1.1% (96.5% vs.95.4%), respectively, which is a significant improvement.
-In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a $<$0.6m accuracy of 98.0% and a $<$3-interval accuracy of 97.9%.
+In addition, on the WHU dataset, our method surpasses the advanced method Cas-MVSNet and achieves an MAE error of 0.091m, a <<0.6m accuracy of 98.0% and a <<3-interval accuracy of 97.9%.
 It can be seen that our proposed method also has superior generalization performance in multi-view stereo reconstruction methods.
-5.2.6 Object Tracking
-TABLE XII: Object Tracking: Comp. on Visdrone-MOT
+Appendix H Object Tracking
+H.1 Datasets and Evaluation Metrics
+TABLE $X$: Object Tracking: Comp. on Visdrone-MOT
 Methods MOTA IDF1
@@
 RingMo-Aerial 52.6 67.8
-TABLE XIII: Object Tracking: Comp. on AIR-MOT
+TABLE XXI: Object Tracking: Comp. on AIR-MOT
 Methods MOTA IDF1
@@
 RingMo-Aerial 68.6 81.3
-TABLE XIV: Object Tracking: Comp. on AIR-HSAO
+TABLE XXII: Object Tracking: Comp. on AIR-HSAO
 Methods MOTA IDF1
@@
 RingMo-Aerial 60.7 74.4
-The comparison results on the VisDrone-MOT dataset are shown in Tab.XII.
-RingMo-Aerial can achieve the best tracking results, far exceeding other methods.
-In the AIR-MOT dataset (as shown in Tab.XIII), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack.
-Meanwhile, in the AIR-HSAO dataset (as shown in Tab.XIV), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods.
-Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
-Furthermore, the visualization results of the AIR-MOT sequences demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios.
-The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions.
-This visualization provides qualitative solid evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in ARS videos.
-5.3 Ablation Experiments
-5.3.1 Ablation Experiments on pre-training
-Figure 13: The ablation experiment on the UAVID dataset and UperNet is used as the segmentation head.
-RingMo-Aerial pretrained with MIM and CL achieves the fastest convergence speed and optimal segmentation performance.
-TABLE XV: Ablation Experiments
-UAVID Visdrone-DET
-mAcc mIoU mAP mAP50
-RingMo 71.6 61.9 28.4 45.2
-RingMo-Aerial (MIM) 75.7 65.8 30.5 50.3
-RingMo-Aerial (MIM + CL) 76.5 66.4 31.1 51.8
-TABLE XVI: Ablation Experiments with the proposed ARS-Adapter
-UAVID Visdrone-DET
-mAcc mIoU mAP mAP50
-RingMo-Aerial (MIM) 71.6 62.9 25.3 42.7
-RingMo-Aerial (MIM + CL) 74.3 64.4 31.0 51.4
-We utilize UperNet model on UAVID and Cascade RCNN model on Visdrone-DET.
-The comparative analysis involves three foundation models: the Swin Transformer Base model pre-trained on ImageNet22K [22], the RingMo-Aerial model solely pre-trained with MIM, and the RingMo-Aerial model pre-trained with both MIM and CL.
-This comparative investigation aims to substantiate the efficacy and consistency of the proposed method.
-The comparison is shown in Fig.13 and Tab.XV.
-The experimental findings reveal a progressive convergence speed and performance enhancement across the three foundation models.
-Notably, the RingMo-Aerial model trained with MIM and CL exhibits the swiftest convergence and attains the optimal performance.
-This observation underscores the synergistic relationship between the FE-MSA module proposed in this study and the MIM + CL pre-training strategy, affirming their mutually reinforcing nature and consistent effectiveness.
-In addition, models pretrained with MIM combined with CL exhibit enhanced adaptability when used with ARS-Adapter.
-Consequently, ablation studies were conducted to compare ARS-Adapter between models solely pretrained with MIM and those pretrained with both MIM and CL.
-Comparative results on the UAVID and VisDrone-DET datasets are presented in Tab.XVI.
-The experimental findings indicate that models additionally trained with CL significantly improve performance in both detection and segmentation tasks when fine-tuned with ARS-Adapter.
-5.3.2 Performance Comparison Between ARS-Adapter and Other Fine-Tuning Methods
-TABLE XVII: Comp. between fine-tuning methods
-Methods Parameters UAVID VisDrone-DET
-mAcc mIoU mAP mAP50
-Froezn 0% 61.2 49.3 21.5 36.2
-Full 100% 76.5 66.4 31.1 51.8
-Adapter 0.39% 68.9 57.7 28.9 49.1
-AdaptFormer 0.39% 71.3 62.1 30.3 50.8
-ARS-Adapter * 0.41% 74.3 64.4 31.0 51.4
-    Although ARS-Adapter is less accurate than global fine-tuning, it significantly outperforms other fine-tuning methods when the proportion of fine-tuning parameters is almost constant.
-In the context of ARS, traditional model fine-tuning presents specific challenges.
-These include the propensity for overfitting due to the extensive pre-training on diverse tasks, which can lead to a model that is overly tailored to the nuances of the training data, thereby compromising its applicability to new, unseen data.
-Overfitting is particularly problematic in ARS, where models are not only expected to generalize across various tasks but also to adapt to the unique perspectives and resolutions inherent in aerial imagery.
-This contrasts with ApRS, which typically operates within a more constrained perspective and resolution, reducing the risk of overfitting to specific views.
-The ARS-Adapter, therefore, addresses these challenges by introducing a parameter-efficient fine-tuning approach that minimizes overfitting risks while maintaining high performance, as evidenced by the experimental results.
-This section compares ARS-Adapter against other fine-tuning methods, using UAVID and VisDrone-DET datasets with UperNet and Cascade RCNN as the segmentation and detection frameworks, respectively.
-The aim is to assess the effectiveness of ARS-Adapter compared to conventional fine-tuning techniques.
-The experimental results are presented in Tab.XVII.
-For clarity, we compare ARS-Adapter with full fine-tuning, zero fine-tuning, and two other fine-tuning methods (Adapter and AdaptFormer).
-Full fine-tuning is a performance upper bound, while zero fine-tuning provides a reference for minimal parameter adjustment.
-ARS-Adapter requires a similar proportion of parameters(0.41%) to be adjusted as Adapter(0.39%) and AdaptFormer(0.39%), but achieves superior fine-tuning efficiency.
-Additionally, the performance loss compared to full fine-tuning(100%) is limited to approximately 2%, demonstrating that ARS-Adapter can achieve high performance with minimal parameter adjustments.
-These results indicate that ARS-Adapter maintains high performance while significantly reducing the number of parameters requiring adjustment, demonstrating its efficiency in fine-tuning.
-6 Conclusion
-The RingMo-Aerial model presented in this paper represents the first large foundation model specifically tailored for Aerial Remote Sensing (ARS) vision tasks.
-Following an in-depth analysis of the unique aspects of ARS vision tasks, this study identified the necessity for foundation model research in this domain and proposed a series of innovative solutions based on the RingMo framework to address the specific challenges therein.
-Our experiments demonstrate that the RingMo-Aerial model achieves SOTA performance across various datasets for ARS vision tasks.
-This demonstrates the utility and effectiveness of our approach in improving ARS vision tasks.
-In the future, RingMo-Aerial and its associated technologies are poised to provide a robust foundation model for ARS vision research.
-Future work will focus on improving the model’s generalization capabilities, computational efficiency, and broader real-world deployment.
-Additionally, we will explore the extension of the RingMo-Aerial model to other vision tasks and scenarios, aiming to expand its application scope.
-Appendix A Semantic Segmentation
-A.1 Datasets and Evaluation Metrics
-For semantic segmentation, the UAVID [140], UDD6 [141] and FloodNet[142] datasets are employed.
-  • UAVID.
-The UAVID dataset has 42 sequences of 4K resolution in total (20 train, 7 valid, and 15 test), and each sequence contains ten pictures.
-In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
-A total of 27 sequences in the training and validation sets are used for training, and 15 sequences in the test set are used for testing.
-  • UDD6.
-The training set of the UDD6 dataset includes 106 images, and the validation set includes 35 images.
-All of the images are 4K resolution.
-In the experiment, all images were cropped into 1024×1024 pixel slices with 256 pixels of overlap.
-The training set is used for training, and the validation set is used for testing.
-  • FloodNet.
-The FloodNet dataset contains 500 high-resolution remote sensing images with a resolution of 4K.
-The dataset is divided into training, validation, and test sets, with the training set containing 300 images, the validation set containing 100 images, and the test set containing 100 images.
-Images are annotated as flood and non-flood areas, supporting multi-class flood impact assessment.
-All images are cropped into 1024×1024 pixel slices with a 256-pixel overlap in the experiment, facilitating the training and evaluation of machine learning models.
-In our experiment of semantic segmentation, the evaluation metrics are mean intersection over union (mIoU), mean accuracy (mAcc), and average accuracy (aAcc).
-A.2 Experiment Settings
-The experiment of the semantic segmentation is based on segmentation [143], UperNet [144], and Mask2Former [145] are employed as the segmentation framework.
-The training process spanned a total of 40,000 iterations with a batch size of 8 total (2 per GPU and 4 NVIDIA A40 GPU are used).
-The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
-In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
-Figure 14: Qualitative comparison of semantic segmentation results on the UAVID dataset using different methods.
-The figure illustrates differences in capturing details such as vegetation, moving vehicles, and buildings, with highlighted regions emphasizing variations in detail and performance.
-Appendix B Object Detection
-B.1 Datasets and Evaluation Metrics
-To conduct a comprehensive comparison of object detection methods, we selected four unimodal benchmark datasets, including the VisDrone-DET[147], UAVDT[148] and ShipDataset[149], as well as one multimodal benchmark dataset, DroneVehicle[92] and our collection of SAR airborne remote sensing data, to demonstrate the performance of the representative methods discussed in this paper.
-  • VisDrone-DET.
-The Visdrone-DET dataset is the detection task part of the VisDrone dataset.
-The resolution of the images is about 2k.
-The RingMo-Aerial experiment crops the original image to 640×640 and uses multi-scale cropping and resizing for enhancement.
-The training set and validation set are used for training, and the labeled test-dev set is used for testing.
-  • UAVDT.
-The UAVDT dataset is another popular ARS detection dataset, containing 23,258 images for training and 15,069 images for testing with an average resolution of 1080 × 540 pixels.
-The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
-  • ShipDataset.
-The ShipDataset, captured at a resolution of $3840\times 2160$, encompasses five distinct scenarios in Shanghai, China, characterized by varying shooting angles and lighting conditions.
-This dataset has $96.97\%$ small ships.
-The images within this dataset predominantly feature congested ship traffic.
-The RingMo-Aerial experiment uses multi-scale resizing for enhancement.
-  • DroneVehicle.
-The DroneVehicle dataset is a large-scale dataset specifically designed for drone-based RGB-Infrared cross-modality vehicle detection tasks.
-It includes 28,439 pairs of RGB-Infrared images, totaling 56,878 images, covering various urban scenarios such as roads, residential areas, and parking lots under different lighting conditions, including day, night, and dark night.
-The RingMo-Aerial experiment uses multi-scale cropping and resizing for enhancement.
-  • AerialSARData.
-To evaluate the performance of our foundational model across various modalities, we employed a hexacopter drone (1800 Pro) equipped with a miniature Synthetic Aperture Radar (MS-102) mounted on its underside.
-This setup was utilized to collect a dataset of SAR-based aerial remote sensing for vehicle detection.
-The dataset comprises a total of 711 images, with the training set consisting of 536 images and the test set comprising 175 images, all at a resolution of 800x800 pixels.
-The images encompass a diverse range of scenarios, including airports, roads, and residential areas.
-In our object detection experiments, the evaluation metrics employed are mean average precision (mAP), mAP at an IoU threshold of 50 (mAP50), and mAP at an IoU threshold of 75 (mAP75).
-For datasets containing smaller objects, we also utilize mAPs, which represents the mean average precision for detecting small objects (with pixels smaller than 32×32).
-B.2 Experiment Settings
-The experiment of the object detection is based on mmdetection [150], Cascade RCNN [67] and DDQ DETR [78] are employed as the detection framework.
-The training process spanned a total of 12 epochs for Cascade RCNN, with a batch size of 32 total (8 per GPU and 4 NVIDIA A40 GPU are used).
-For DDQ DETR, 36 epochs were spanned with a batch size of 16 total (4 per GPU and 4 NVIDIA A40 GPU are used).
-The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0001, accompanied by a weight decay of 0.05.
-In addition, the learning rate is decreased by 10 after the 8th and 11th epochs for Cascade RCNN and after the 27th and 33rd epochs for DDQ DETR.
-Figure 15: RingMo-Aerial combined with Cascade RCNN achieves superior detection accuracy compared to the original RingMo model, particularly in complex environments and small object scenarios.
-It demonstrates reduced false positives and improved bounding box alignment with Ground Truth, highlighting its effectiveness for aerial imagery detection.
-Figure 16: Qualitative comparison of object detection results on infrared images from the DroneVehicle-IR dataset using different methods.
-The figure highlights differences in the detection quality of vehicles, road elements, and pedestrian areas, with emphasized regions showcasing variations in accuracy and detail among the compared methods.
-Figure 17: Qualitative comparison of object detection results on the AerialSARData dataset using different methods.
-The figure highlights differences in detection performance for SAR images, particularly focusing on identifying objects of interest with highlighted regions showing variations in accuracy and detection quality among the methods compared.
-Appendix C Change Detection
-C.1 Datasets and Evaluation Metrics
-For semantic segmentation, the LEVIR [110] and CDD [151] datasets are employed.
-  • LEVIR.
-The LEVIR-CD is a publicly available, extensive remote sensing change detection dataset.
-It comprises 637 pairs of high-resolution images, each with a dimension of 1024 × 1024 pixels at a 0.5-meter resolution.
-Adhering to the standard dataset partitioning, we have processed the original images by extracting non-overlapping patches, each measuring 256 × 256 pixels.
-Consequently, this results in a distribution of 7120, 1024, and 2048 image patches allocated for training, validation, and testing phases.
-  • CDD.
-The CDD dataset is a comprehensive collection of bi-temporal remote sensing imagery, comprising seven pairs of images with a resolution of 4725×2700 pixels and an additional four pairs with dimensions of 1900×1000 pixels.
-These images have been meticulously synchronized and cropped to yield 16,000 pairs of 256×256 pixel patches.
-This curated subset is strategically divided into a training set of 10,000 pairs, a validation set of 3,000 pairs, and a test set of the remaining 3,000 pairs.
-To enhance the realism and applicability of the trained models, the dataset incorporates seasonal variations, thereby providing a more challenging and convincing foundation for network training.
-C.2 Experiment Settings
-Our model adopts BIT with Swin Transformer as the backbone.
-The experiments use overall accuracy (OA) and F1 and IoU scores about change detection as the primary evaluation metrics.
-The model is trained with an AdamW optimizer with a weight decay of 0.01 and beta values equal to (0.9, 0.999).
-The LR is 2e-3 for the LEVIR-CD dataset and linearly decays to 0 until trained for 200 epochs, with a total batch size 16 for datasets (8 per GPU, using 2 NVIDIA A40 GPUs).
-Figure 18: Comparison of Detection Visualization Results Using Different Backbone Networks with the BIT Algorithm on the LEVIR Dataset.
-That shows the results of aerial images, where RingMo-Aerial outperforms ResNet18 and RingMo, with fewer false positives and false negatives.
-RingMo-Aerial closely matches the Ground Truth, demonstrating the best accuracy in change Detection.
-The black region represents true negative, the white region represents true positive, the green region represents false negative, the red region represents false positive.
-Appendix D Scene Classification
-D.1 Datasets and Evaluation Metrics
-For scene classification, the AID [152] and RESISC-45 [153] datasets are employed.
-  • AID. The AID dataset is a large-scale data set used to evaluate the performance of aviation scene classification. It contains images collected from Google Earth and covers 30 scenes.
-  • NWPU-RESISC45.
-The NWPU-RESISC45 dataset is a publicly available Remote Sensing Image Scene Classification (RESISC) benchmark.
-The dataset contains 31,500 images covering 45 scene classes, with 700 images in each class.
-In our experiment of scene classification, the training proportions in the AID dataset were set to TR = 20% and TR = 50%, respectively.
-In the NWPU-RESISC dataset, the proportion of training samples was set to TR = 10% and TR = 20%, respectively.
-The experiments use overall accuracy (OA) as a performance metric, the most commonly used performance metric in scene classification.
-D.2 Experiment Settings
-The experiment of the scene classification is based on mmpretrain [154], Cascade RCNN [67] is employed as the classification framework.
-The training process spanned a total of 300 calendar elements, with a total batch size of 16 for dataset AID (8 per GPU, using 2 NVIDIA A40 GPUs) and 128 for dataset NWPU-RESISC45 (64 per GPU, using 2 NVIDIA A40 GPUs).
-The AdamW [146] optimizer is utilized for the optimization, and the base learning rate is set to 0.0005, accompanied by a weight decay of 0.05.
-In addition, the Poly learning rate strategy is used for learning rate scheduling, and the final learning rate drops to 0.
-Appendix E 3D-Reconstruction
-E.1 Datasets and Evaluation Metrics
-For 3D-Reconstruction, the and WHU [118] datasets are employed.
-  • WHU.
-The WHU dataset is a synthetic aerial dataset designed for large-scale multi-view stereo reconstruction tasks.
-It was generated from a highly accurate 3D surface model and includes a complete set of aerial images as well as cropped sub-image sets specifically for deep learning applications.
-The images in the dataset cover an area of approximately 6.7 × 2.2 square kilometers in Meitan County, Guizhou Province, China, with a ground resolution of 0.1 meters.
-The WHU dataset provides crucial data support for the study of large-scale Earth surface and urban reconstruction and can be utilized for training and testing multi-view stereo matching algorithms.
-  • LuoJia-MVS.
-The LuoJia-MVS dataset comprises 7,972 five-view images, each with a spatial resolution of 10 cm, accompanied by pixel-wise depth information and precise camera parameters.
-These images were generated from a highly accurate digital surface model (DSM) constructed using thousands of stereo aerial images.
-During the experiment, we used the WHU dataset[118] and the LuoJia-MVS dataset[155] to verify the effectiveness of our proposed method.
-WHU dataset[118] is the first large-scale aerial multi-view stereo dataset for remote sensing.
-It was collected from Meitan, Guizhou, China, and was obtained by taking large-scale photos with a tilted five-eye camera on a drone platform.
-The area of the entire dataset is approximately $6.7\times 2.2km^{2}$, and the image resolution is about 0.1m.
-In addition, the tilt angle of each camera to the central camera is 40°.
-The WHU dataset gathers high-rise buildings, mountains, forests, and factory scenes, providing a good benchmark for remote sensing multi-view stereo.
-LuoJia-MVS[155] dataset is a large-scale multi-view stereo dataset collected from a variety of landform types.
-It was collected in a part of Baiyun, Guiyang, Guizhou, China.
-The image size of the entire dataset is $768\times 384$, and the image resolution is about 0.1m.
-The data consists of five views and pixel-by-pixel depth labels.
-Land types cover cultivated land, forest land, industrial areas, residential areas, etc.
 Figure 20: Qualitative results on the target tracking dataset AIR-MOT using RingMo-Aerial.
@@
 The category and target ID are displayed above the bounding boxes, demonstrating RingMo-Aerial’s ability to effectively distinguish and track individual targets.
-E.2 Experiment Settings
-Our work is coded based on the PyTorch framework.
-The proposed method is trained and tested on NVIDIA A40, and the batch size is set to 1.
-During the training process, our method uses the Adam optimizer[156] for parameter optimization and trains for 24 epochs, where the parameter $\beta_{1}$ is set to 0.9, the parameter $\beta_{2}$ is set to 0.999, and the initial learning rate is 0.001.
-In addition, our method is consistent with the existing multi-stage networks[119].
-First, multi-scale features are learned through a multi-scale feature extractor, and the scales are 1/16, 1/4, and 1 of the input size, respectively.
-Then, deep optimization is performed through three stages of profound hypotheses, and the number of hypotheses is 48, 32, and 8, respectively.
-During the experiment, we use three metrics, mean absolute error (MAE), $<$0.6m, and $<$3-interval, to evaluate the performance of the proposed method.
-Appendix F Object Tracking
-F.1 Datasets and Evaluation Metrics
 For Object Tracking, the following aerial dataset is employed.
@@
 In our experiment of object tracking, Multiple Object Tracking Accuracy (MOTA) and Identification Flatness (IDF1) are used as evaluation metrics.
-F.2 Experiment Settings
+H.2 Experiment Settings
 In the proposed experiment, object tracking is predicated on the detection model of RingMo-Aerial on VisDrone-DET.
@@
 Subsequently, the model’s multi-object tracking performance will be quantified and analyzed.
-Acknowledgments
-This work was supported by the National Nature Science Foundation of China under Grant 62331027 and 62301538.
+H.3 Experiment Results
+The comparison results on the VisDrone-MOT dataset are shown in Tab.XX.
+RingMo-Aerial can achieve the best tracking results, far exceeding other methods.
+In the AIR-MOT dataset (as shown in Tab.XXI), which contains more small aircraft and boat targets, RingMo-Aerial achieved a MOTA of 68.6 and an IDF1 of 81.3, though slightly trailing behind CFTrack.
+Meanwhile, in the AIR-HSAO dataset (as shown in Tab.XXII), RingMo-Aerial maintained its robust performance with a MOTA of 60.7 and an IDF1 of 74.4, approaching the best-performing methods.
+Overall, the RingMo-Aerial algorithm exhibited excellent multi-object tracking capabilities in these complex scenarios, coming close to or surpassing the state-of-the-art methods in several key metrics.
+Furthermore, the visualization results of the AIR-MOT sequences (as shown in Fig.20) demonstrate the tracking performance of the RingMo-Aerial method in four selected video sequences, particularly in complex and dynamic scenarios.
+The clarity and precision of the bounding boxes further highlight the method’s effectiveness in distinguishing and tracking various objects under different environmental conditions.
+This visualization provides qualitative solid evidence of the proposed algorithm’s performance in handling multi-object tracking tasks in ARS videos.